### v1

# organization=Université de Rennes, IRISA – UMR 6074,addressline=263 Av. Général Leclerc, city=Rennes, postcode=35000, country=France

Abstract
In this article, we give an extended space formulation for the induced tree and path polytopes of chordal graphs with variables associated with the edge and vertex sets. Whereas the formulation for the induced tree polytope is easily seen to have a compact size, the system we provide for the induced path polytope has an exponential number of inequalities. We show which of these inequalities define facets and exhibit a superset of the facet-defining ones that can be enumerated in polynomial time. We show that for some graphs, the latter superset contains redundant inequalities. As corollaries, we obtain that the problems of finding an induced tree or path maximizing a linear function over the edges and vertices are solvable in polynomial time for the class of chordal graphs .
Keywords— induced path , induced trees , extended formulations , polytope , chordal graphs
Definitions
All the graphs in this article are simple and connected unless specified. Given a graph $G=(V,E)$, we denote its complement by $\overline{G}=(V,\overline{E})$, where $\overline{E}=\{e\in\binom{V}{2}:e\notin E\}$. We respectively denote by $V(G)$ and $E(G)$ the vertex set and the edge set of $G$. Two vertices $u$ and $v$ are adjacent if $uv\in E(G)$. Given a subset of vertices $W\subseteq V$, let $E(W)$ denote the set of edges of $G$ having both endpoints in $W$ and $\delta(W)$ the set of all edges having exactly one endpoint in $W$. When $W$ is a singleton $\{w\}$, we will simply write $\delta(w)$. We say that the edges in $\delta(w)$ are incident to $w$, and two edges sharing an extremity are said to be adjacent. For $F\subseteq E$, let $V(F)$ denote the set of vertices incident to any edge of $F$. Given $W\subseteq V$, the graph $G[W]=(W,E(W))$ is the subgraph of $G$ induced by $W$. When $H$ is an induced subgraph of $G$, we say that $G$ contains $H$. Given a vertex $u\in V(G)$, we denote by $N_{G}(u)=\{w\in V(G):uw\in E(G)\}$ its neighborhood, and by $N_{G}[u]=N_{G}(u)\cup\{u\}$ its closed neighborhood, when clear from context we simply write $N(u)$ and $N[u]$. Two vertices $u$ and $v$ are true twins if $N[u]=N[v]$ and false twins if $N(u)=N(v)$. A vertex $w$ is complete to a subset of vertices $W$ if it is adjacent to each vertex of $W$ and does not belong to $W$. A vertex $u$ complete to $V\setminus\{u\}$ is said to be universal. A vertex is simplicial if its neighborhood induces a clique.
An induced path (resp. hole) is a graph induced by a sequence of vertices $(v_{1},\dots,v_{p})$ whose edge set is $\{v_{i}v_{i+1}:\ i=1,\dots,p-1\}$ (resp. $\{v_{i}v_{i+1}:\ i=1,\dots,p-1\}\cup\{v_{1}v_{p}\}$ ). A subset of vertices induces a path (resp. hole) if its elements can be ordered into a sequence inducing a path (resp. hole). The length of a path or hole is its number of edges. A path or hole is called long if its length is at least 4. The parity of a path or hole is the parity of its length. Given a path induced by the sequence $(w_{1},\dots,w_{k})$, we call $w_{1}$, $w_{k}$ its extremities and the set of vertices $\{w_{2},\dots,w_{k-1}\}$ its interior. A tree is a graph with no hole.
The contraction of an edge $uv$ in $G$ yields a new graph $G/uv$ built from $G$ by deleting $u$, $v$, adding a new vertex $w$, and adding the edges $wz$ for all $z\in N_{G}(u)\cup N_{G}(v)$. For $F\subseteq E$, we denote by $G/F$ the graph obtained from $G$ by contracting all the edges in $F$.
A graph is $k$-chordal if its holes have maximum size $k$, 3-chordal graphs are called chordal. Equivalently, chordal graphs are those for which an order $(v_{1},\dots,v_{k})$ of its vertices exists such that $v_{i}$ is simplicial in $G_{i}=G[\{v_{i},\dots,v_{k}\}]$, such an order is called a perfect elimination order. Moreover, the maximal cliques of $G$ are among $N_{G_{i}}[v_{i}]$, which shows that they come in a linear number. A tree is a connected graph containing no holes.
Let $\mathcal{P}=\{x:Ax\leq b\}\subseteq\mathds{R}^{n}$ be a polyhedron, that is, the intersection of finitely many half-spaces. The dimension of $\mathcal{P}\subseteq\mathds{R}^{n}$, denoted by dim $\mathcal{P}$, is the maximum number of affinely independent points in $\mathcal{P}$ minus one. If $a\in\mathds{R}^{n}\setminus\{0\}$, $\alpha\in\mathds{R}$, then the inequality $a^{\top}x\leq\alpha$ is said to be valid for $\mathcal{P}$ if $\mathcal{P}\subseteq\{x\in\mathds{R}^{n}:a^{\top}x\leq\alpha\}.$ We say that a valid inequality $a^{\top}x\leq\alpha$ defines a facet of $\mathcal{P}$ if $\textrm{dim}(\mathcal{P}\cap\{x\in\mathds{R}^{n}:a^{\top}x=\alpha\})$ is equal to $\textrm{dim}~\mathcal{P}$ minus one. The polyhedron $\mathcal{P}\subseteq\mathds{R}^{n}$ is full dimensional when dim $\mathcal{P}$ is equal to $n$. In that case, a linear system $Ax\leq b$ that defines $\mathcal{P}$ is minimal if each inequality of the system defines a facet of $\mathcal{P}$. Moreover, these facet-defining inequalities are unique up to scalar multiplication. A face of $\mathcal{P}$ is a polyhedron contained in $\mathcal{P}$ obtained by setting to equality some inequalities in $Ax\leq b$. Equivalently, a face is the intersection of the supporting hyperplane of a valid inequality and $\mathcal{P}$. Extreme points of a polyhedron are its faces of dimension zero. Each extreme point satisfies $\textrm{dim}~\mathcal{P}$ linearly independent valid inequalities for $\mathcal{P}$ with equality. A polyhedron $\mathcal{P}\subseteq\mathds{R}^{n}$ is said to be integer if all its non-empty faces contain an integer point, it is binary if it is included in the $n$-dimensional hypercube and all its extreme points are binary. A polytope is a bounded polyhedron, equivalently, it can be defined as the convex hull of a finite number of points $P$ and is denoted by $\textrm{conv}(P)$.
In this work, the polyhedra we consider are convex hulls of points encoding structured induced subgraphs of an input graph $G=(V,E)$. Such an induced subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ (note that $E^{\prime}=E(V^{\prime})$) can be encoded by its natural incidence vector that is $\chi^{V^{\prime}}\in\{0,1\}^{V}$ whose components associated with elements of $V^{\prime}$ are equal to 1 and the others equal to 0, or by the extended incidence vector that is $\xi^{V^{\prime}}=(\chi^{V^{\prime}},\zeta^{E^{\prime}})\in\{0,1\}^{V}\times\{0,1\}^{E}$ where the components of $\zeta^{E^{\prime}}$ associated with elements of $E$ are equal to 1 and the others equal to 0. Given a point = $(x,y)\in\mathds{R}^{V\cup E}$ we denote by $(x,y)_{G^{\prime}}$ its restriction to elements of $G^{\prime}$. A cone is a polyhedron described by a zero right-handed linear system; it is pointed if it contains no affine space. The elements of a cone can be described by conic combinations of finite sets of elements called generating. The minimal generating families of pointed cones are unique up to positive scalar multiplication; we denote the element of such a family by its generators.
Motivations
In this paper, we study the problems of finding maximum weight induced trees and paths where both the edges and the vertices have weight, and in particular, filling a gap in polyhedral results associated with these problems.
Induced trees
The problem of finding the size of the largest induced tree in a graph is a problem that has been approached through extremal combinatorics [ERDOS198661, MATOUSEK2007307, FOX2009494]. In particular, an important question is that of bounding this number according to several graph parameters, such as the number of vertices, edges, radius, independence number, maximum clique, and connectivity. In [ERDOS198661], a Ramsey-type bound is proposed on the minimum integer $n(k,t)$ such that every graph having at least $n(k,t)$ vertices contains either a clique of size $k$ or an induced tree of size $t$. Finally, they show that every connected graph with radius $r$ has an induced tree of size at least 2r12r-1 by exhibiting a large induced path. A related problem is the one of finding large induced forests [inducedforest] in which bounds for the largest induced forest are given with respect to maximum degree or stability number. Besides these studies, no polyhedral study of the induced tree polytope on particular classes has been proposed. The problem of computing the maximum size of an induced tree is known to be NP-hard [completePathTree] in general, but it is also the case in bipartite and triangle-free graphs with maximal degree 3 [MATOUSEK2007307]. Several integer linear programs have been proposed to find maximum weighted such subsets [ILPinducedtree], and extensive computational experiments have been made on general graphs.
Induced paths
The problem of finding a longest induced path has applications in assessing the robustness of a network. It can be viewed as a variant of the graph diameter. In a graph where vertices represent communication devices that may fail, and edges represent possible transmission links, the length of the longest induced path indicates the worst-case number of intermediate devices required for a one-to-one communication between any two vertices $u$ and $v$ after the failure of a set of devices that do not disconnect $u$ from $v$. Another application is about error correcting codes [error]. The induced paths of a graph are induced trees, and the type of questions raised in the previous paragraph can be raised considering this structure. The authors of [FOX2009494] show that every connected triangle-free graph on $n$ vertices has an induced path of size at least $\sqrt{n}$, and their approach can be generalized to obtain lower bounds for graphs with no clique of given size $k$. In [esperet], the authors show that every 3-connected planar graph contains an induced path on $\text{log}(|V|)$ vertices. In [jfraymond], it is observed that a graph may contain a large path but no large induced path; for this reason, they characterize the sets of edges to be forbidden along a path of size $n$ to force the existence of an induced path of size $log(n)$. This result allows them to show that if a graph has a path of length $n$ and no $K_{t}$ as a topological minor, then it contains an induced path of order $(log n)^{\Omega(1/tlog^{2}t)}$.
Finding a maximum vertex weighted induced path is also NP-hard [completePathTree], but polynomially solvable in $k$-chordal graphs, interval filament, graphs decomposable by clique cutset or by splits whose prime subgraphs are polynomially solvable [GAVRIL2002203], improved in [HANSEN2009135] and on graphs of bounded clique width, asteroidal triple free graphs [inducedpathcliquewifth]. Besides these combinatorial algorithms, no polyhedral study of the corresponding polytope has been made. A related polyhedra has, however, been characterized in [goel2025halfspacerepresentationspathpolytopes]: the polytope whose extreme points are incidence vectors of paths between pairs of leaves in a tree: in trees, induced path and path are the same subgraphs, so the latter paper describes the induced path (between leaves) polytope of trees. Finally, extended space ILPs for this problem have been proposed [ILPinducedPath, ILPinducedPath2], leading to computational experiments.
Contributions
In this paper, we first exhibit an extended formulation of compact size characterizing the induced tree polytope of chordal graphs, showing that the problem of computing a maximum vertex and/or edge weighted induced tree is solvable in polynomial time in chordal graphs. For a given graph, the induced tree polytope contains the induced path polytope, and both characterizations follow similar schemes of proof.
Even if the maximum vertex weighted induced path is already known to be solvable in polynomial time in chordal graphs [GAVRIL2002203, HANSEN2009135] it is unknown whether it is the case for vertex and edge weighted graphs. For this reason, we provide an extended formulation for the induced path polytope. The set of inequalities of this formulation is exponential, and we characterize which of them define facets. Finally, we show that a polynomial superset of the facet-defining ones can be enumerated efficiently. This exhibits a compact formulation for the induced path polytope of chordal graphs. Note that both latter results are necessary as our characterization of the facet-defining ones does not show that they come in a polynomial number.
1 Induced tree polytope of chordal graphs
In this section, we give an extended formulation for the induced tree polytope of a chordal graph. Given a graph $G$, and a vertex $v$, we denote by $\mathcal{K}^{v}$ the set of maximal cliques containing $v$ and $\mathcal{T}(G)$ the following system of inequalities:
$\sum_{u\in K}y_{uv}$ $\leq x_{v}$ $\forall v\in V,\ K\in\mathcal{K}^{v}$ (1)
$-x_{v}$ $\leq 0$ $\forall v\in V$ (2)
$-y_{e}$ $\leq 0$ $\forall e\in E$ (3)
$x(V)-y(E)$ $=1$ (4)
We now prove that this system fully characterizes induced tree polytopes in chordal graphs.
Theorem 1.1.
A graph $G$ is chordal if and only if $\mathcal{T}(G)$ is binary.
Proof. $\Rightarrow"$ Note that the system of inequalities (1)–(3) defines a pointed cone $\mathcal{C}(G)$. The extreme points of $\mathcal{T}(G)$ are exactly the scalings of $\mathcal{C}(G)$’s generators that satisfy (4). Observe that induced trees’ extended incidence vectors belong to $\mathcal{C}(G)$ and satisfy (4). We show that they are the generators of $\mathcal{C}(G)$ by induction on a simplicial decomposition of $G$.
To initialize the induction, consider the graph $G$ with a single vertex, $\mathcal{C}(G)$ is the positive half line trivially generated by the unit vector.
Induction goes as follows: let $v$ be a simplicial vertex of $G$ and suppose that $\mathcal{C}(H)$ is generated by induced trees’ extended incidence vectors for $H=G\setminus\{v\}$. Let $(x^{*},y^{*})$ be a point of $\mathcal{C}(G)$, $(x^{*},y^{*})_{H}$ is a point of $\mathcal{C}(H)$ by construction. By the induction hypothesis, let $(x^{*},y^{*})_{H}$ be generated by the induced trees $\Gamma$ with coefficients $\lambda$. The proof consists of iteratively modifying $\Gamma$ and $\lambda$ so that they yield a conic combination equal to $(x^{*},y^{*})$. At each step, we denote by $(x^{\prime},y^{\prime})$ the point given by the current conic combination associated with $\Gamma$ and $\lambda$. Each operation preserves the value of $(x^{\prime},y^{\prime})_{H}$ but increments either $y_{uv}^{\prime}$ or $x_{v}^{\prime}$. We first modify the conic combination so that $y_{uv}^{*}=y_{uv}^{\prime}\ \forall uv\in E$. We consider the following notation $\lambda(\Gamma^{\prime})=\sum_{T\in\Gamma^{\prime}}\lambda_{T}$ for every $\Gamma^{\prime}\subseteq\Gamma$. Let us consider the following subsets of $\Gamma$:
  • $\Gamma^{u}\ \forall u\in N(v)$ containing its elements whose intersection with $N[v]$ equals $\{u\}$.
  • $\Gamma^{u}_{2}\ \forall u\in N(v)$ containing its elements whose intersection with $N(v)$ is $\{u,w\}$ for some $w\in N(v)$.
  • $\Gamma^{u}_{v}\ \forall u\in N(v)$ containing the elements having $v$ as an extremity and containing $u$
We refer to Figure 1 for an illustration of the sets of induced trees considered. Note that $u$ is either a leaf or an internal node in the elements of $\Gamma^{u}_{2}$, we provide two different corresponding induced trees in different shades of green, which are the two bottom ones.
To each element $T$ of $\Gamma^{u}$, $T\cup\{v\}$ is an induced tree of $G$. This makes it possible to apply the Algorithm (1) over each element of $\bigcup_{u\in N(v)}\Gamma^{u}$, incrementing $y_{uv}^{\prime}$ and $x_{u}^{\prime}$.
Let $T$ be an element of $\Gamma^{u}$
$\alpha=\min(\lambda_{T},y_{uv}^{*}-y_{uv}^{\prime})$
$T^{\prime}=T\cup\{v\}$
$\lambda_{T}=\lambda_{T}-\alpha$
$\lambda_{T^{\prime}}=\lambda_{T^{\prime}}+\alpha$
$T$ is removed from $\Gamma$ if $\alpha=\lambda_{T}$
$T^{\prime}$ is added to $\Gamma$
Algorithm 1 Reduction of $\Gamma^{u}$
Algorithm 1 increases $y_{uv}^{\prime}$ unless $y_{uv}^{\prime}=y_{uv}^{*}$ or $\Gamma^{u}$ is empty.
Suppose that $\Gamma^{u}$ is empty then, $y_{uv}^{*}\leq y_{uv}^{\prime}\ \forall uv\in\delta(v)$. As $(x^{*},y^{*})$ belongs to $\mathcal{C}(G)$ we have the following:
$y_{uv}^{*}+\lambda(\Gamma_{2}^{u})=y^{*}_{uv}+y^{*}(\delta(u,N(v)\setminus\{u\}))\leq x_{u}^{*}=x_{u}^{\prime}=\lambda(\Gamma^{u}_{v})+\lambda(\Gamma_{2}^{u})=y_{uv}^{\prime}+\lambda(\Gamma_{2}^{u})$
The first equality is given by construction, the inequality is given by inequality (1) associated with $u$, and the remaining equalities are obtained by construction. This show how $y_{uv}^{*}=y_{uv}^{\prime}$ can be achieved on all edges $uv\in\delta(v)$. Finally, the singleton $\{v\}$ can be added to $\Gamma$ and $\lambda_{v}$ set to $x_{v}^{*}-y(\delta(v))^{*}$ to achieve $(x^{\prime},y^{\prime})=(x^{*},y^{*})$ and exhibit a conic comination of $(x^{*},y^{*})$.
$\Leftarrow"$ By contradiction, suppose that $G$ is not chordal, and let $C$ induce a hole of size at least 4 in $G$, and $v$ belongs to $C$. By setting $x_{w}=1\ \forall w\in V\setminus v$, $y_{e}=1\ \forall e\in E(C)$ and $x_{v}=2$ and 0 on all other components, we obtain a point which belongs to $\mathcal{T}(G)$ and is not included in the hypercube implying that $\mathcal{T}(G)$ is itself not contained in the hypercube and hence not binary. ∎
[…]
Figure 1: Sets of induced trees considered in the proof of Theorem 1.1
Corollary 1.2.
The induced tree polytope of a graph $G$ in the extended variable space is equal to $\mathcal{T}(G)$ if and only if $G$ is chordal.
The set of inclusion-wise maximal cliques of chordal graphs is polynomial (1)–(3), implying that Optimizing a linear function on the vertex and edge set over $\mathcal{T}(G)$ is doable in polynomial time.
Corollary 1.3.
The edge and vertex weighted maximum induced tree is solvable in polynomial time on chordal graphs.
2 Induced path polytope of chordal graphs
We first present a technical lemma concerning edge contraction in chordal graphs exhibited in [dupontbouillard2024contractions]. After that, we give an exponential description of the induced path polytope of chordal graphs and characterize the inequalities that define facets. Finally, we show that the non-dominated such constraints come in a polynomial number.
Lemma 2.1 ([dupontbouillard2024contractions]).
If $G=(V,E)$ is chordal, then $G/F$ also is chordal for any set $F\subseteq E$.
Given a set of vertices $W\subseteq V$, we denote by $C_{G}(W)$ the vertices complete to $W$ in $G$, we simply write $C(W)$ when clear from context. We denote by $\mathcal{P}(G)$ the following system:
$y(\delta(w,C(K\cup\{w\})))+2y(\delta(w,K))$ $\leq 2x_{w}$ $\ \forall w\in V,\ \text{ clique }K\subseteq N(w)$ (5)
$-x_{w}$ $\leq 0$ $\forall w\in V$ (6)
$-y_{e}$ $\leq 0$ $\forall e\in E$ (7)
$x(V)-y(E)$ $=1$ (8)
The next theorem shows that $\mathcal{P}(G)$ describes the induced path polytope of chordal graphs. Inequalities (5) express the fact that an induced path contains at most two edges in $\delta(w)$ (for $K=\emptyset$) and the extremitie of two such edges cannot induce a triangle (for $K\neq\emptyset$). We refer to Figure 2 for an illustration where the blue edges are the ones involved with coefficient 2 and the green ones are involved with coefficient one. At most two green edges belong to a same induced path every blue edge is not contained in any induced path containing some other colored edge.
Theorem 2.2.
A graph $G=(V,E)$ is chordal if and only if $\mathcal{P}(G)$ is binary:
Proof. $\Rightarrow$ We proceed by showing that the pointed cone $\mathcal{C}(G):=$(5)–(7) is generated by the extended incidence vectors of induced paths. The extreme points of (5)–(8) are scalings of the generators of $\mathcal{C}(G)$ satisfying (8).
We now show that the set of induced paths’ extended induced path generates $\mathcal{C}(G)$. To do so, we proceed by induction on a perfect elimination order of $G$. Initialization of the induction is given by the single vertex graph $G$ for which $\mathcal{C}(G)=\mathds{R}_{+}$.
Induction goes as follows: let $v$ be a simplicial vertex of $G$ and suppose $\mathcal{C}(H)$ to be generated by $H$’s induced path extended incidence vectors for $H=G\setminus v$. Let $(x^{*},y^{*})$ be a point of $\mathcal{C}(G)$, $(x^{*},y^{*})_{H}$ is a point of $C(H)$.
By the induction hypothesis, let $(x^{*},y^{*})_{H}$ be generated by the induced path $\Psi$ with coefficients $\lambda$. The proof consists of iteratively modifying $\Psi$ and $\lambda$ so that they yield a conic combination equal to $(x^{*},y^{*})$. At each step, we denote by $(x^{\prime},y^{\prime})$ the point given by the current conic combination associated with $\Psi$ and $\lambda$; an induced path $P$ is removed from $\Psi$ if $\lambda_{P}$ is set to 0. Each operation preserves the value of $(x^{\prime},y^{\prime})_{H}=(x^{*},y^{*})_{H}$. Let us fix an edge $uv\in\delta(v)$, we exhibit a conic combination whose $y_{uv}$ component is equal to $y_{uv}^{*}$, as applying the procedure will preserve the value of $y_{e}$ for the other edge $e\in\delta(v)\setminus\{uv\}$, it can be applied until the conic combination is equal to $(x^{*},y^{*})$ on all components but $x_{v}$.
Let us consider the following subsets of $\Psi$:
  • $\Psi^{u}$ is the subset of $\Psi$ whose elements have $u$ as an extremity and do not contain any other element of $N[v]$
  • $\Psi_{2}^{u}$ is the subset of $\Psi$ having $u$ as an extremity and containing another element of $N(v)$
  • $\Psi^{u}_{I}$ is the subset of $\Psi$ whose interior contains $u$ and does not contain $v$
  • $\Psi_{u}^{v}$ is the subset of $\Psi$ containing both $u$ and $v$, note that $v$ is necessarily an extremity of $\Psi_{u}^{v}$’s elements
  • $\Psi_{\bot}$ contains the remaining elements.
We refer to Figure 3 for an illustration of the different types of path considered. To each element of $\Psi^{u}$, the addition of $v$ yields another induced path which contains $\{u,v\}$. This implies that unless $y_{uv}^{\prime}$ is already equal to $y_{uv}^{*}$, a conic combination preserving $(x^{\prime},y^{\prime})_{H}$ and strictly increasing over $y_{uv}$ can be computed. Algorithm 2 describes such an operation.
let $P$ be an element of $\Psi^{u}$
$\alpha=\min\{y_{uv}^{*}-y_{uv}^{\prime},\lambda_{P}\}$
$P^{\prime}=P\cup\{v\}$
$(x^{\prime},y^{\prime})=(x^{\prime},y^{\prime})-\alpha\chi^{P}+\alpha\chi^{P^{\prime}}$
$\lambda_{P}=\lambda_{P}-\alpha$
$\lambda_{P^{\prime}}=\lambda_{P^{\prime}}+\alpha$
Algorithm 2 Reduction of $\Psi^{u}\ \forall u\in N(v)$
We can now suppose that $\Psi^{u}$ is empty (otherwise equality of the conic combination’s $y_{uv}$ component and $y_{uv}^{*}$ is already achieved) which implies the following:
$x^{*}_{u}$ $=\lambda(\Psi_{2}^{u})+\lambda(\Psi_{I}^{u})+\lambda(\Psi_{u}^{v})$ (9)
$y_{uv}^{\prime}$ $=\lambda(\Psi_{u}^{v})$ (10)
Claim 2.3.
If $\Psi_{I}^{u}$ is empty, then $y_{uv}^{\prime}\geq y_{uv}^{*}$.
Proof. Suppose that $\Psi^{u}_{I}$ is empty. Inequality (5) associated with $u$ and $K=N[v]\setminus\{u\}$ applied to $(x^{*},y^{*})$ yields: $x_{u}^{*}\geq y^{*}(\delta(u,K))$ and the right hand side is larger than or equal to $\lambda(\Psi_{2}^{u})+y_{uv}^{*}$. Using equality (9) and (10) we get : $y_{uv}^{\prime}\geq y_{uv}^{*}.$
Suppose now that $\Psi_{I}^{u}$ is non-empty.
let $P$ be an element of $\Psi_{I}^{u}$
let $P_{1}$ and $P_{2}$ be the two different maximal subpath of $P$ ending in $u$.
$\alpha=\min(\lambda_{P},\lambda_{u,v}$)
$\lambda_{P}=\lambda_{P}-\alpha$
Let $i\in\{1,2\}$ be such that $P_{i}\cup\{v\}$ induces a path
$j=i+1\mod 2$
$\lambda_{u,v}=\lambda_{u,v}-\alpha$
$\lambda_{P_{j}}=\lambda_{P_{j}}+\alpha$
$\lambda_{P_{i}\cup\{v\}}=\lambda_{P_{i}\cup\{v\}}+\alpha$
Algorithm 3 Removal of $\{u,v\}$ from $\Psi_{u}^{v}$
Algorithm (3) is used to ensure that each element of $\Psi_{u}^{v}$ intersects $\delta(u)$ twice, such a path always contains $uv$ and another edge in $\delta(u)$. When Algorithm (3) cannot be applied, either $\Psi_{I}^{u}$ is empty and Claim (2.3) concludes or $\lambda_{u,v}=0$.
Claim 2.4.
If $\Psi_{2}^{u}$ is empty, then $y_{uv}^{\prime}\geq y_{uv}^{*}$.
Proof. Suppose that $\Psi_{2}^{u}$ is empty. Inequality (5) associated with $u$ and $K=\emptyset$ applied to $(x^{*},y^{*})$ yields: $2x_{u}\geq y^{*}(\delta(u))$. Note that the right hand side is larger than or equal to $2\lambda(\Psi_{I}^{u})+\lambda(\Psi_{u}^{v})+y_{uv}^{*}$ as each $\lambda_{P}$ for $P\in\Psi_{I}^{u}\cup\Psi^{v}_{u}$ contributes two times in $y^{*}(\delta(u))$ and $\lambda(\Psi^{v}_{u})=y_{uv}^{*}$. Equalities (9) and (10) yield $y_{uv}^{\prime}\geq y_{uv}^{*}$ and conclude the proof. ∎
Suppose now that $\Psi_{2}^{u}$ is nonempty, let us consider the clique $K\subseteq N(u)\cap N(v)$ of vertices contained in at least one element of $\Psi_{2}^{u}$. We now consider the partition of $\Psi_{I}^{u}$ into $\Psi_{I_{1}}^{u}$ containing the elements intersecting $K$ and $\Psi_{I_{0}}^{u}$ containing the others. Algorithm 4 permits to remove elements from $\Psi_{I_{1}}^{u}$ and adds elements in $\Psi_{u}^{v}$ or $\Psi_{2}^{u}$.
Let $P$ be in $\Psi_{I_{1}}^{u}$
Decompose $P$ into $P_{1}$ and $P_{2}$ the two maximal subpath of $P$ ending in $u$
$\alpha=\min(y_{uv}^{*}-y_{uv}^{\prime},\lambda_{P})$
$\lambda_{P}=\lambda_{P}-\alpha$
Let $i\in\{1,2\}$ be such that $P_{i}\cup\{v\}$ induces a path
$j=i+1\mod 2$
$\lambda_{P_{i}\cup\{v\}}=\lambda_{P_{i}\cup\{v\}}+\alpha$
$\lambda_{P_{j}}=\lambda_{P_{j}}+\alpha$
Algorithm 4 Reduction of $\Psi_{I_{1}}^{u}$
Using Algorithm 4 permits obtaining a conic combination of induced path such that $y_{uv}^{\prime}=y_{uv}^{*}$, or no element of $\Psi_{I}^{u}$ intersects $K$ that is $\Psi_{I}^{u}=\Psi_{I_{0}}^{u}$.
Claim 2.5.
Given $P_{1},P_{2}$ be two induced paths of a chordal graph having a similar extremity $u$, such that the neighbor $u_{1}$ of $u$ in $P_{1}$ is different from the neighbor $u_{2}$ of $u$ in $P_{2}$. Either $u_{1}$ is adjacent to $u_{2}$ or $P_{2}\cup P_{1}$ induces a path.
Proof. By contradiction, suppose that $P_{2}\cup P_{1}$ does not induce a path. Let $w_{1}$ be the other element of $P_{1}$ adjacent to $u_{1}$, by hypothesis, $w_{1}\neq u_{2}$.
As $P_{1}$ and $P_{2}$ have an extremity in common but their union does not induce a path, either they intersect in some other vertex than $u$ or there exists an edge going from $P_{1}\setminus u$ to $P_{2}\setminus u$. In both situations, either $u_{2}$ is adjacent to $u_{1}$ or there exist a cycle $C\subseteq P_{1}\cup P_{2}\textasciidieresis$ containing $\{u,u_{1},w_{1},u_{2}\}$. Let us contract all the vertices of $P_{1}\cup P_{2}\setminus\{u_{1},u_{2},u\}$ to a single vertex $w$. The subgraph induced by $\{u_{1},u_{2},u,w\}$ contains a cycle of size 4, and $w$ is non-adjacent to $u$ by construction. By chordality of $G$ and Lemma (2.1), $u_{2}$ is adjacent to $u_{1}$. ∎
[…]
Figure 2: Structure of inequalities (5) for a given vertex $w$ and clique $K$ of its neighborhood.
Claim (2.5) now permits to decompose the elements of $\Psi_{I_{0}}^{u}$ for which a neighbor of $u$ is not complete to $K$ into their two maximal subpaths ending in $v$. We then combine one of these subpaths with an element of $\Psi_{2}^{u}$ to obtain other induced paths.
Algorithm 5 gives the procedure. Note that Algorithm 5 may add elements to $\Psi^{u}$ and $\Psi_{I_{1}}^{u}$ which can be empty again using Algorithm 2 and 4. Note that the fact that Algorithm 2 and 4 strictly increase the value of $y_{uv}^{\prime}$ ensures that the sequence of updates on $\lambda$ and $\Psi$ converges.
Let $P_{0}\in\Psi_{I_{0}}^{u}$ and $P_{2}\in\Psi_{2}^{u}$ be such that the $P_{0}^{\prime}\cup P_{2}$ induces a path where $P_{0}^{\prime}$ is a maximal subpath of $P_{0}$ ending in $u$.
$\alpha=\min(\lambda_{P_{2}},\lambda_{P_{0}})$
$\lambda_{P_{2}}=\lambda_{P_{2}}-\alpha$
$\lambda_{P_{0}}=\lambda_{P_{0}}-\alpha$
$\lambda_{P_{2}\cup P_{0}^{\prime}}=\lambda_{P_{2}\cup P_{0}^{\prime}}+\alpha$
$\lambda_{(P_{0}\setminus P_{0}^{\prime})\cup\{v\}}=\lambda_{(P_{0}\setminus P_{0}^{\prime})\cup\{v\}}+\alpha$
Algorithm 5 Replacing the elements of $\Psi_{I_{0}}^{u}$ so that in the remaining ones the neighbors of $u$ are complete to $K$
Now we suppose that the elements of $\Psi^{u}_{I_{0}}$ are complete to $K$ (otherwise, $\Psi^{u}_{2}$ is empty and Claim 2.4 concludes).
Using another time Claim 2.5, we get rid of elements in $\Psi_{u}^{v}$ for which the neighbor of $u$ distinct from $v$ is not complete to $K$ with Algorithm 6. Note that Algorithm 6 increments $\lambda_{uv}$ which is then decreased back using Algorithm 3. It also increments $\Psi_{I_{1}}^{u}$, which is then decreased back using Algorithm 4 if a path in which a neighbor of $u$ is not complete to $K$ is added. Such a sequence of modifications strictly increases $y_{uv}^{\prime}$, ensuring convergence. Note that the neighbor of $u$ in $P_{i}$ as defined in Algorithm 3 is now complete to $K$ as we already used Algorithm 5.
Let $P_{4}$ be in $\Psi_{u}^{v}\setminus\{u,v\}$ and $P_{2}\in\Psi_{2}^{u}$ such that $P_{4}\cup P_{2}\setminus\{v\}$ induces a path
$\alpha=\min(\lambda_{P_{4}},\lambda_{P_{2}})$
$\lambda_{P_{4}}=\lambda_{P_{4}}-\alpha$
$\lambda_{P_{2}}=\lambda_{P_{2}}-\alpha$
$\lambda_{P_{4}\cup P_{2}\setminus\{v\}}=\lambda_{P_{4}\cup P_{2}\setminus\{v\}}+\alpha$
$\lambda_{u,v}=\lambda_{uv}+\alpha$.
Algorithm 6 The neighbors of $u$ in $G[P]$ for $P\in\Psi^{u}$ are complete to $K$
As $(x^{*},y^{*})$ is a point of $\mathcal{C}(G)$, it satisfies the constraint (5) associated with $K$ and $u$ that is:
$2x_{u}^{\prime}=2x_{u}^{*}\geq 2y^{*}(\delta(u,K))+y^{*}(\delta(u,C(K\cup\{u\})))$ (11)
By construction, each element $P_{4}$ of $\Psi_{u}^{v}$ contains both $v$ and another neighbor of $u$ and both are complete to $K$ implying that $\lambda_{P_{4}}$ contributes one time in $y^{*}[\delta(u,C(K\cup\{u\})\setminus\{uv\})]$. By construction, $y^{*}(\delta(u,C(K\cup\{u\})))=y^{*}_{uv}+y^{*}[\delta(u,C(K\cup\{u\})\setminus\{uv\})]=y^{*}_{uv}+\lambda(\Psi_{u}^{v})+2\lambda(\Psi_{I}^{u})$ . Using equalities (9) and (10) we get:
$2x_{uv}^{*}-2x_{uv}^{*}\geq 2\lambda(\Psi_{2}^{u})+y^{*}_{uv}+2\lambda(\Psi_{I}^{u})+\lambda(\Psi_{u}^{v})-2\lambda(\Psi_{2}^{u})-2\lambda(\Psi_{I}^{u})-2\lambda(\Psi_{u}^{v})$
Which finally shows that $y_{uv}^{\prime}\geq y_{uv}^{*}$. One can repeat this procedure until $y_{uv}^{*}=y_{uv}^{\prime}\ \forall uv\in\delta(v)$. Setting $\lambda_{\{v\}}=(x^{*}_{u}-y^{*}(\delta(v)))$ yields the desired conic combination of $(x^{*},y^{*})$ using induced paths’ extended incidence vectors.
$(\Leftarrow)$ The proof is by contradiction. The same non binary point exhibited in the proof of Theorem 1.1 also belongs to $\mathcal{P}(G)$. ∎
[…]
Figure 3: Induced paths considered in the proof of Theorem 2.2
Note that inequalities (6) are mandatory only when the graph has isolated vertices. If the input graph has no such vertex, then inequalities (5) and (7) ensure that $x$ is positive. These constraints have been added to give easy initializations of the induction. The proof of Theorem 2.2 shows that the extreme points of $\mathcal{P}(G)$ are induced paths’ extended incidence vectors, which implies the following.
Corollary 2.6.
A graph $G$ is chordal if and only if its extended space induced path polytope is equal to $\mathcal{P}(G)$.
It turns out that not all inequalities (5) define facets, we characterize which do.
Theorem 2.7.
Given a graph $G=(V,E)$, a vertex $w$, and a clique $K\in N(w)$, the associated inequality (5) defines a facet if and only if no connected component of $\overline{G}[C_{G}(K)]$ is bipartite.
Proof. $(\Rightarrow)$ Let us consider the connected components $\overline{G}_{1},\dots,\overline{G}_{k}$ of $\overline{G}[C_{G}(K)]$ and suppose by contradiction that for $i\leq\ell$, $\overline{G}_{i}$ is bipartite composed of the stable sets induced by $S_{i}^{1},S_{i}^{2}$. By construction, $K_{j}=\bigcup_{i=1}^{\ell}S_{i}^{j}\cup K,\ j=1,2$ is a clique of $G$ such that $C_{G}(K_{j})=\bigcup_{i=\ell+1}^{k}V(\overline{G}_{i})$. By summing $2x_{w}\geq 2y(\delta(w,K_{j}))+y(\delta(w,C(K_{j})))$ for $j=1,2$ we obtain the inequality associated with $K$ and $w$ showing that it is dominated and hence not facet-defining.
$(\Leftarrow)$ We exhibit |V|+|E|1|V|+|E|-1 linearly independent points satisfying (5) with equality, the fact that the equality constraints submatrix’s rank is at least 1 because of equality 8 concludes:
  1. 1.
    $\xi^{v}\ \forall v\in V\setminus\{w\}$
  2. 2.
    $\xi^{u,v}\ \forall uv\in E\setminus\delta(w)$
  3. 3.
    $\xi^{u,w}\ \forall uw\in\delta(w,K)$
  4. 4.
    $\xi^{u,w,v}\ \forall uw\in\delta(w)\setminus\delta(K)\setminus\delta(C(K))$ for some vertex $v\in K$ non adjacent to $u$
The remaining $|\delta(w,C(K))|$ points are built the following way: for each connected components $\overline{G}_{c}$ of $\overline{G}[C(K)]$, let us consider $T_{c}\subseteq E(\overline{G}_{c})$ inducing a subgraph containing an odd cycle and hitting every vertex of $G_{c}$, $T_{c}$ exists by hypothesis and may choosen such that $|T_{c}|=|V(\overline{G}_{c})|$. This can be done by starting from an odd cycle and then adding the missing edges for this set to be spanning the vertices. Now, match each edge of $\delta(w,C(K))$ to a distinct edge of $T=\bigcup T_{c}$ sharing an extremity, note that $|T|=|C(K)|=|\delta(w,C(K))|$. Let $uw$ be in $\delta(w,C(K))$ and $uv$ be its corresponding edge of $T$. In $G$, $u$ and $v$ are non adjacent hence $\xi^{u,v,w}$ is the incidence vector of an induced path for which inequality (5) for $w$ and $K$ is tight. This construction is necessary as such path will contain two edges of $\delta(w,C(K))$ making it hard to build an indentity of wanted size over these components.
We refer to Figure 4 for an illustration of the matrix whose rows are the points previously exhibited, and ”?” stands for the submatrices associated with incidence vectors of 3-sized paths and whose value does not need to be explicit for the validity of the proof. The rank of $H$ is unkown, however, we know that the restriction of the last $|\delta(w,C(K))|$ points to elements of $G\setminus\delta(w,C(K))$ contains a incidence matrix of a non bipartite graph $T$, by Theorem 2.1 of [GROSSMAN1995213] it has full row rank $|C(K)|$.
Linear independence is obtained by the fact that the matrix is block triangular with a non-zero diagonal over the points (1)–(4), and the fact that the latter set of points forms a submatrix of rank $|C(K)|$ which are the only ones involving elements of $\delta(w,C(K))$. ∎
[…]
Figure 4: Matrix of points built in the proof of Theorem 2.7 where $G^{\prime}$ is the graph with vertex set $V$ and edge set $(E\setminus\delta(w))\cup(\delta(w)\cap\delta(K))$
The characterization of Theorem 2.7 gives no information about the number of such facet-defining inequalities. We show that the non-dominated inequalities (5) come in a polynomial number and can be enumerated efficiently. Given a fixed vertex $w$, for any two cliques $K\subseteq N(w)$ and $K^{\prime}\subseteq N(w)$, if $K\subseteq K^{\prime}$ then we have $C(K^{\prime})\subseteq C(K)$, for this reason, we say that a clique defines an orbit if there exists no $K^{\prime}$ strictly containing $K$ such that $K\cup C(K)\subseteq K^{\prime}\cup C(K^{\prime})$. Suppose that a clique $K$ is not orbit-defining, let $K^{\prime}$ be the orbit-defining clique containing $K$, inequality (5) associated with $(w,K)$ is dominated with respect to inequality (5) associated with $(w,K^{\prime})$ and non-negativity constraints.
Theorem 2.8.
Given a chordal graph $G$, the set of non-dominated inequalities inequalities (5) has polynomial size.
Proof. We build a chordal auxiliary graph $H$ to $G[N(w)]$, whose maximal cliques are in bijection with orbit-defining cliques of $G[N(w)]$. Polynomial size of $H$ and linear number of maximal cliques of chordal graphs will conclude.
The auxiliary graph $H$ is built from $G[N(w)]$ by first adding a true twin to each vertex of $N(w)$, let $W$ be the set of added twins. Secondly, add all the edges between elements of $W$. A perfect elimination ordering of $H$ is given by any perfect elimination of $G[N(w)]$ and followed by any permutation of the vertices of $W$. This proves that $H$ is chordal.
Let $K_{H}$ be a maximal clique of $H$. It is partitioned into $K=K_{H}\cap N(w)$ and $K_{W}=K_{H}\cap W$. We associate with $K_{H}$ a unique orbit-defining clique of $G[N(w)]$. By construction, $K$ is a clique of $G[N(w)]$, and we show that it defines an orbit. By contradiction, suppose that a clique $K^{\prime}$ of $G[N(w)]$ strictly contains $K$ and is such that $K\cup C(K)\subseteq K^{\prime}\cup C(K^{\prime})$. Let $K_{W^{\prime}}\subseteq W$ be the set of twins of vertices in $K^{\prime}\cup C(K^{\prime})$. By construction, $K^{\prime}\cup K_{W^{\prime}}$ is a clique of $H$ strictly containing $K_{H}$, a contradiction. Now, let $K$ define an orbit of $G[N(w)]$, we associate to $K$ a unique maximal clique of $H$ as follows. Let $W_{K}\subseteq W$ be the set of twins of $C(K)\cup K$, $K_{H}=W_{K}\cup K$ is a clique of $H$ by construction. We show that $K_{H}$ is maximal, let $v$ be a vertex of $W\setminus W_{K}$, $v$ is complete to $W_{K}$ by construction, and suppose it to also be complete to $K$. Then, its twin $u$ also is complete to $K$, implying that $u\in C(K)$ and $v\in W_{k}$, a contradiction. Let now $v\in N_{G}(w)$ be a vertex complete to $K_{H}$, then, in particular, it is complete to $W_{K}$, implying that it is complete to $C(K)$. By construction we have $K\subsetneq K\cup\{v\}$ and $K\cup C(K)\subseteq K\cup\{v\}\cup C(K\cup\{v\})$, a contradiction to $K$ defining an orbit. ∎
Corollary 2.9.
The maximum vertex and weighted induced path problem is solvable in polynomial time in chordal graphs.
It turns out that the target of Theorems 2.7 and 2.8 differ, as there exist orbit-defining cliques in chordal graphs whose associated inequalities are not facet-defining as given by Figure 5. Vertices 1, 2, and 3 form an orbit defining clique such that $C(\{1,2,3\})=\{4,5,6,7\}$ however, the complement of the subgraph induced by $\{4,5,6,7\}$ is a hole of size 4 which is bipartite therefore, Theorem 2.7 implies that its associated inequalities are not facet defining.
11223344556677
Figure 5: An orbit defining clique whose associated inequalities (5) are not facet defining
Conclusion
In this paper, we propose extended formulations for the induced tree and path polytopes of chordal graphs of compact size, implying the polynomial solvability of finding a maximum vertex and edge weighted such induced subgraph in chordal graphs. Our proofs rely heavily on the structure of the induced trees and paths cones; it may be interesting to investigate the structures of the linear programs whose constraint sets correspond to incidence vectors of induced paths and trees.
Acknowledgment
The author thanks the ANR project ANR-24-CE23-1621 EVARISTE for financial support.

### v2

@@
-organization=Université de Rennes, IRISA – UMR 6074,addressline=263 Av. Général Leclerc, city=Rennes, postcode=35000, country=France
+Extended formulations for induced tree and path polytopes of chordal graphs
 Abstract
-In this article, we give an extended space formulation for the induced tree and path polytopes of chordal graphs with variables associated with the edge and vertex sets.
+In this article, we give two extended space formulations, respectively, for the induced tree and path polytopes of chordal graphs with vertex and edge variables.
+These formulations are obtained by proving that the induced tree and path extended incidence vectors of chordal graphs form Hilbert basis.
+This also shows that both polytopes have the integer decomposition property in chordal graphs.
 Whereas the formulation for the induced tree polytope is easily seen to have a compact size, the system we provide for the induced path polytope has an exponential number of inequalities.
@@
 As corollaries, we obtain that the problems of finding an induced tree or path maximizing a linear function over the edges and vertices are solvable in polynomial time for the class of chordal graphs.
-Keywords— induced path , induced trees , extended formulations , polytope , chordal graphs
+keywords:
+[…]
 Definitions
@@
 When $W$ is a singleton $\{w\}$, we will simply write $\delta(w)$.
-We say that the edges in $\delta(w)$ are incident to $w$, and two edges sharing an extremity are said to be adjacent.
+We say that the edges in $\delta(w)$ are incident to $w$, and two edges sharing an endpoint are said to be adjacent.
+Given two subsets of vertices $W_{1}$, $W_{2}$, we denote by $\delta(W_{1},W_{2})$ the set of edges having one endpoint in both $W_{1}$ and $W_{2}$.
 For $F\subseteq E$, let $V(F)$ denote the set of vertices incident to any edge of $F$.
@@
 A vertex $u$ complete to $V\setminus\{u\}$ is said to be universal.
+Given a set of vertices $W\subseteq V$, we denote by $C_{G}(W)$ the vertices complete to $W$ in $G$; we simply write $C(W)$ when clear from context.
 A vertex is simplicial if its neighborhood induces a clique.
@@
 $\{v_{i}v_{i+1}:\ i=1,\dots,p-1\}\cup\{v_{1}v_{p}\}$ ).
+Note that sets of vertices inducing cliques with three vertices are considered holes.
+An induced tree is a set of vertices inducing an acyclic graph.
 A subset of vertices induces a path (resp.
@@
 A tree is a graph with no hole.
-The contraction of an edge $uv$ in $G$ yields a new graph $G/uv$ built from $G$ by deleting $u$, $v$, adding a new vertex $w$, and adding the edges $wz$ for all $z\in N_{G}(u)\cup N_{G}(v)$.
+The contraction of an edge $uv$ in $G$ yields a new graph $G/uv$ built from $G$ by deleting $u$, $v$, adding a new vertex $w$, and adding the edges $wz$ for all $z\in N_{G}(u)\cup N_{G}(v)$; parallel edges and loops are removed.
 For $F\subseteq E$, we denote by $G/F$ the graph obtained from $G$ by contracting all the edges in $F$.
-A graph is $k$-chordal if its holes have maximum size $k$, 3-chordal graphs are called chordal.
+A graph is $k$-chordal if its holes have maximum size $k$; 3-chordal graphs are called chordal.
 Equivalently, chordal graphs are those for which an order $(v_{1},\dots,v_{k})$ of its vertices exists such that $v_{i}$ is simplicial in $G_{i}=G[\{v_{i},\dots,v_{k}\}]$, such an order is called a perfect elimination order.
 Moreover, the maximal cliques of $G$ are among $N_{G_{i}}[v_{i}]$, which shows that they come in a linear number.
-A tree is a connected graph containing no holes.
+The size of a linear system is its number of rows.
 Let $\mathcal{P}=\{x:Ax\leq b\}\subseteq\mathds{R}^{n}$ be a polyhedron, that is, the intersection of finitely many half-spaces.
-The dimension of $\mathcal{P}\subseteq\mathds{R}^{n}$, denoted by dim $\mathcal{P}$, is the maximum number of affinely independent points in $\mathcal{P}$ minus one.
+The dimension of $\mathcal{P}\subseteq\mathds{R}^{n}$, denoted by $\textrm{dim}~\mathcal{P}$, is the maximum number of affinely independent points in $\mathcal{P}$ minus one.
 If $a\in\mathds{R}^{n}\setminus\{0\}$, $\alpha\in\mathds{R}$, then the inequality $a^{\top}x\leq\alpha$ is said to be valid for $\mathcal{P}$ if $\mathcal{P}\subseteq\{x\in\mathds{R}^{n}:a^{\top}x\leq\alpha\}.$ We say that a valid inequality $a^{\top}x\leq\alpha$ defines a facet of $\mathcal{P}$ if $\textrm{dim}(\mathcal{P}\cap\{x\in\mathds{R}^{n}:a^{\top}x=\alpha\})$ is equal to $\textrm{dim}~\mathcal{P}$ minus one.
@@
 Each extreme point satisfies $\textrm{dim}~\mathcal{P}$ linearly independent valid inequalities for $\mathcal{P}$ with equality.
-A polyhedron $\mathcal{P}\subseteq\mathds{R}^{n}$ is said to be integer if all its non-empty faces contain an integer point, it is binary if it is included in the $n$-dimensional hypercube and all its extreme points are binary.
+A polyhedron $\mathcal{P}\subseteq\mathds{R}^{n}$ is said to be integer if all its non-empty faces contain an integer point, it is binary if it is bounded and all its extreme points are binary.
 A polytope is a bounded polyhedron, equivalently, it can be defined as the convex hull of a finite number of points $P$ and is denoted by $\textrm{conv}(P)$.
+A cone is a polyhedron described by a linear system with a right hand side equal to 0; it is pointed if it contains no affine space.
+The elements of a cone can be described by conic combinations of finite sets of elements called generators.
+The minimal generating sets of pointed cones are unique up to positive scalar multiplication; we denote the element of such a family by its generators they are in one to one correspondence with its face of dimension 1.
+A polyhedron $\mathcal{P}$ is said to have the integer decomposition property (IDP) if every integer point in the dilation $k\mathcal{P}$ for $k\in\mathds{Z}_{+}$ is obtained as the sum of $k$ elements of $\mathcal{P}$.
+A set of points $X=\{\chi_{1},\dots,\chi_{n}\}$ forms a Hilbert basis if every integer point in the conic hull of $X$ is expressible as an integer nonnegative combination of elements of $X$.
+The Hilbert basis and integer decomposition property are closely related.
+If the extreme points of a polytope $\mathcal{P}$ form a Hilbert basis, then $\mathcal{P}$ has the IDP.
 In this work, the polyhedra we consider are convex hulls of points encoding structured induced subgraphs of an input graph $G=(V,E)$.
-Such an induced subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ (note that $E^{\prime}=E(V^{\prime})$) can be encoded by its natural incidence vector that is $\chi^{V^{\prime}}\in\{0,1\}^{V}$ whose components associated with elements of $V^{\prime}$ are equal to 1 and the others equal to 0, or by the extended incidence vector that is $\xi^{V^{\prime}}=(\chi^{V^{\prime}},\zeta^{E^{\prime}})\in\{0,1\}^{V}\times\{0,1\}^{E}$ where the components of $\zeta^{E^{\prime}}$ associated with elements of $E$ are equal to 1 and the others equal to 0.
-Given a point = $(x,y)\in\mathds{R}^{V\cup E}$ we denote by $(x,y)_{G^{\prime}}$ its restriction to elements of $G^{\prime}$.
-A cone is a polyhedron described by a zero right-handed linear system; it is pointed if it contains no affine space.
-The elements of a cone can be described by conic combinations of finite sets of elements called generating.
-The minimal generating families of pointed cones are unique up to positive scalar multiplication; we denote the element of such a family by its generators.
+Such an induced subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ (note that $E^{\prime}=E(V^{\prime})$) can be encoded by its natural incidence vector that is $\chi^{V^{\prime}}\in\{0,1\}^{V}$ whose components associated with elements of $V^{\prime}$ are equal to 1 and the others equal to 0, or by its extended incidence vector that is $\xi^{V^{\prime}}=(\chi^{V^{\prime}},\zeta^{E^{\prime}})\in\{0,1\}^{V\cup E}$ where the components of $\zeta^{E^{\prime}}$ associated with elements of $E$ are equal to 1 and the others equal to 0.
+Given a point $(x,y)\in\mathds{R}^{V\cup E}$ we denote by $(x,y)|_{G^{\prime}}$ its restriction to elements of $G^{\prime}$.
+In this paper, given a graph $G$, we consider the following polyhedra:
+  • its induced tree cone $\text{cone}\{\xi^{W}:\text{ $W$ induces a tree of $G$}\}$
+  • its induced tree polytope $\text{conv}\{\xi^{W}:\text{ $W$ induces a tree of $G$}\}$
+  • its induced path cone $\text{cone}\{\xi^{W}:\text{ $W$ induces a path of $G$}\}$
+  • its induced path polytope $\text{conv}\{\xi^{W}:\text{ $W$ induces a path of $G$}\}$
 Motivations
-In this paper, we study the problems of finding maximum weight induced trees and paths where both the edges and the vertices have weight, and in particular, filling a gap in polyhedral results associated with these problems.
+In this paper, we study the problems of finding maximum weight induced trees and paths where both the edges and the vertices have weight, and in particular, filling a gap in polyhedral literature associated with these problems.
 Induced trees
@@
 A related polyhedra has, however, been characterized in [goel2025halfspacerepresentationspathpolytopes]: the polytope whose extreme points are incidence vectors of paths between pairs of leaves in a tree: in trees, induced path and path are the same subgraphs, so the latter paper describes the induced path (between leaves) polytope of trees.
-Finally, extended space ILPs for this problem have been proposed [ILPinducedPath, ILPinducedPath2], leading to computational experiments.
+Finally, extended space integer linear programs for this problem have been proposed [ILPinducedPath, ILPinducedPath2], leading to computational experiments.
 Contributions
-In this paper, we first exhibit an extended formulation of compact size characterizing the induced tree polytope of chordal graphs, showing that the problem of computing a maximum vertex and/or edge weighted induced tree is solvable in polynomial time in chordal graphs.
-For a given graph, the induced tree polytope contains the induced path polytope, and both characterizations follow similar schemes of proof.
-Even if the maximum vertex weighted induced path is already known to be solvable in polynomial time in chordal graphs [GAVRIL2002203, HANSEN2009135] it is unknown whether it is the case for vertex and edge weighted graphs.
-For this reason, we provide an extended formulation for the induced path polytope.
-The set of inequalities of this formulation is exponential, and we characterize which of them define facets.
-Finally, we show that a polynomial superset of the facet-defining ones can be enumerated efficiently.
-This exhibits a compact formulation for the induced path polytope of chordal graphs.
-Note that both latter results are necessary as our characterization of the facet-defining ones does not show that they come in a polynomial number.
+In this paper, we study both the induced tree polytope and the induced path polytope of chordal graphs in two separate sections.
+It turns out that beside being included one in the other, these two polytopes have similar properties.
+Each section starts by respectively showing that the induced path and tree extended incidence vectors of chordal graphs form a Hilbert basis of some cones, which we explicitly describe by linear systems of inequalities.
+We show that intersecting these cones with the correct hyperplane yields the induced tree and path polytope of chordal graphs in extended variable space.
+While the system we provide for the induced tree polytope is easily seen to be compact in chordal graphs, the one we provide for the induced path polytope has an exponential number of inequalities.
+Therefore, we characterize the ones that define facets and show that a superset can be enumerated in polynomial time.
+This implies that there exists a compact system describing the induced path polytope of chordal graphs.
+Finally, these results imply that the problems of finding a maximum vertex and edge weight induced tree or path are polynomially solvable in chordal graphs.
 1 Induced tree polytope of chordal graphs
-In this section, we give an extended formulation for the induced tree polytope of a chordal graph.
-Given a graph $G$, and a vertex $v$, we denote by $\mathcal{K}^{v}$ the set of maximal cliques containing $v$ and $\mathcal{T}(G)$ the following system of inequalities:
-$\sum_{u\in K}y_{uv}$ $\leq x_{v}$ $\forall v\in V,\ K\in\mathcal{K}^{v}$ (1)
+In this section, we show that the induced tree’s extended incidence vectors of chordal graphs form a Hilbert basis.
+Then we use this fact to exhibit a linear system describing the induced tree polytope of chordal graphs.
+Given a graph $G$, and a vertex $v$, we denote by $\mathcal{K}^{v}$ the set of maximal cliques containing $v$ and $\mathcal{C}_{T}(G)$ the cone defined by the following system of inequalities:
+$\sum_{u\in K\setminus\{v\}}y_{uv}$ $\leq x_{v}$ $\forall v\in V,\ K\in\mathcal{K}^{v}$ (1)
 $-x_{v}$ $\leq 0$ $\forall v\in V$ (2)
 $-y_{e}$ $\leq 0$ $\forall e\in E$ (3)
-$x(V)-y(E)$ $=1$ (4)
-We now prove that this system fully characterizes induced tree polytopes in chordal graphs.
+Observe that induced trees’ extended incidence vectors belong to $\mathcal{C}_{T}(G)$.
 Theorem 1.1.
-A graph $G$ is chordal if and only if $\mathcal{T}(G)$ is binary.
+The induced tree extended incidence vectors of a chordal graph form a Hilbert basis.
 Proof.
-$\Rightarrow"$ Note that the system of inequalities (1)–(3) defines a pointed cone $\mathcal{C}(G)$.
-The extreme points of $\mathcal{T}(G)$ are exactly the scalings of $\mathcal{C}(G)$’s generators that satisfy (4).
-Observe that induced trees’ extended incidence vectors belong to $\mathcal{C}(G)$ and satisfy (4).
-We show that they are the generators of $\mathcal{C}(G)$ by induction on a simplicial decomposition of $G$.
-To initialize the induction, consider the graph $G$ with a single vertex, $\mathcal{C}(G)$ is the positive half line trivially generated by the unit vector.
-Induction goes as follows: let $v$ be a simplicial vertex of $G$ and suppose that $\mathcal{C}(H)$ is generated by induced trees’ extended incidence vectors for $H=G\setminus\{v\}$.
-Let $(x^{*},y^{*})$ be a point of $\mathcal{C}(G)$, $(x^{*},y^{*})_{H}$ is a point of $\mathcal{C}(H)$ by construction.
-By the induction hypothesis, let $(x^{*},y^{*})_{H}$ be generated by the induced trees $\Gamma$ with coefficients $\lambda$.
+We show that induced trees extended incidence vectors of a chordal graph $G=(V,E)$ form a Hilbert basis of $\mathcal{C}_{T}(G)$.
+As $\mathcal{C}_{T}(G)$ is defined with rational coefficients, each one-dimensional face contains integer points.
+This implies that a Hilbert basis of $\mathcal{C}_{T}(G)$ generates $\mathcal{C}_{T}(G)$.
+We proceed by induction. For initialization, consider the graph $G$ with a single vertex. $\mathcal{C}_{T}(G)$ is the positive half-line trivially generated by the unit vector.
+Induction goes as follows: let $v$ be a simplicial vertex of $G$ and suppose that the integer points of $\mathcal{C}_{T}(G\setminus v)$ are obtained as positive integer combinations of induced trees’ extended incidence vectors.
+Let $(x^{*},y^{*})$ be an integer point of $\mathcal{C}_{T}(G)$, $(x^{*},y^{*})|_{G\setminus v}$ is an integer point of $\mathcal{C}_{T}(G\setminus v)$ by construction.
+By the induction hypothesis, let $(x^{*},y^{*})|_{G\setminus v}$ be generated by the induced trees $\Gamma$ with integer coefficients $\lambda$.
 The proof consists of iteratively modifying $\Gamma$ and $\lambda$ so that they yield a conic combination equal to $(x^{*},y^{*})$.
 At each step, we denote by $(x^{\prime},y^{\prime})$ the point given by the current conic combination associated with $\Gamma$ and $\lambda$.
-Each operation preserves the value of $(x^{\prime},y^{\prime})_{H}$ but increments either $y_{uv}^{\prime}$ or $x_{v}^{\prime}$.
-We first modify the conic combination so that $y_{uv}^{*}=y_{uv}^{\prime}\ \forall uv\in E$.
+For fixed $uv\in\delta(v)$, each operation preserves the value of $(x^{\prime},y^{\prime})|_{H}$ for $H=G\setminus uv$, and preserves nonnegativity and integrality of $\lambda$.
+We define the slackness $\Delta=\sum_{uv\in\delta(v)}\Delta(u,v)$ for $\Delta(u,v)=y_{uv}^{*}-y_{uv}^{\prime}$.
 We consider the following notation $\lambda(\Gamma^{\prime})=\sum_{T\in\Gamma^{\prime}}\lambda_{T}$ for every $\Gamma^{\prime}\subseteq\Gamma$.
-Let us consider the following subsets of $\Gamma$:
+Let us consider the following subsets of $\Gamma$, illustrated in Figure 1 :
   • $\Gamma^{u}\ \forall u\in N(v)$ containing its elements whose intersection with $N[v]$ equals $\{u\}$.
@@
   • $\Gamma^{u}_{v}\ \forall u\in N(v)$ containing the elements having $v$ as an extremity and containing $u$
-We refer to Figure 1 for an illustration of the sets of induced trees considered.
 Note that $u$ is either a leaf or an internal node in the elements of $\Gamma^{u}_{2}$, we provide two different corresponding induced trees in different shades of green, which are the two bottom ones.
-To each element $T$ of $\Gamma^{u}$, $T\cup\{v\}$ is an induced tree of $G$.
-This makes it possible to apply the Algorithm (1) over each element of $\bigcup_{u\in N(v)}\Gamma^{u}$, incrementing $y_{uv}^{\prime}$ and $x_{u}^{\prime}$.
-Let $T$ be an element of $\Gamma^{u}$
-$\alpha=\min(\lambda_{T},y_{uv}^{*}-y_{uv}^{\prime})$
-$T^{\prime}=T\cup\{v\}$
-$\lambda_{T}=\lambda_{T}-\alpha$
-$\lambda_{T^{\prime}}=\lambda_{T^{\prime}}+\alpha$
-$T$ is removed from $\Gamma$ if $\alpha=\lambda_{T}$
-$T^{\prime}$ is added to $\Gamma$
-Algorithm 1 Reduction of $\Gamma^{u}$
-Algorithm 1 increases $y_{uv}^{\prime}$ unless $y_{uv}^{\prime}=y_{uv}^{*}$ or $\Gamma^{u}$ is empty.
-Suppose that $\Gamma^{u}$ is empty then, $y_{uv}^{*}\leq y_{uv}^{\prime}\ \forall uv\in\delta(v)$. As $(x^{*},y^{*})$ belongs to $\mathcal{C}(G)$ we have the following:
+We define the following operation to be applied on $(\Gamma,\lambda)$, which decreases $\Delta$ by an integer value.
+$\mathbf{\Gamma^{u}}$-Reduction: let $T$ be an element of $\Gamma^{u}$.
+Set $\alpha=\min(\lambda_{T},y_{uv}^{*}-y_{uv}^{\prime})$ increase $\lambda_{T\cup\{v\}}$ and decrease $\lambda_{T}$ by $\alpha$.
+By construction, for any $T\in\Gamma^{u}$, $T\cup\{v\}$ induces a tree and that $\mathbf{\Gamma^{u}}$-Reduction preserves integrality of $\lambda$.
+Moreover, $\xi^{T}|_{H}=\xi^{T\cup\{v\}}|_{H}$ implying that $\lambda(\Psi)|_{H}$ is preserved.
+$\mathbf{\Gamma^{u}}$-Reduction applies until $\Gamma^{u}$ is empty or $\Delta(u,v)=0$.
+Suppose that $\Gamma^{u}$ is empty then, $y_{uv}^{*}\leq y_{uv}^{\prime}\ \forall uv\in\delta(v)$. As $(x^{*},y^{*})$ belongs to $\mathcal{C}_{T}(G)$ we have the following:
 $y_{uv}^{*}+\lambda(\Gamma_{2}^{u})=y^{*}_{uv}+y^{*}(\delta(u,N(v)\setminus\{u\}))\leq x_{u}^{*}=x_{u}^{\prime}=\lambda(\Gamma^{u}_{v})+\lambda(\Gamma_{2}^{u})=y_{uv}^{\prime}+\lambda(\Gamma_{2}^{u})$
 The first equality is given by construction, the inequality is given by inequality (1) associated with $u$, and the remaining equalities are obtained by construction.
-This show how $y_{uv}^{*}=y_{uv}^{\prime}$ can be achieved on all edges $uv\in\delta(v)$.
-Finally, the singleton $\{v\}$ can be added to $\Gamma$ and $\lambda_{v}$ set to $x_{v}^{*}-y(\delta(v))^{*}$ to achieve $(x^{\prime},y^{\prime})=(x^{*},y^{*})$ and exhibit a conic comination of $(x^{*},y^{*})$.
+This show how $\Delta(u,v)=0$ can be achieved on all edges $uv\in\delta(v)$.
+Finally, the singleton $\{v\}$ can be added to $\Gamma$ and $\lambda_{v}$ set to $x_{v}^{*}-y(\delta(v))^{*}$ to achieve $(x^{\prime},y^{\prime})=(x^{*},y^{*})$ with integer $\lambda$.
+Note that Theorem 1.1 is not a characterization of the graphs for which the induced tree’s extended incidence vectors form a Hilbert basis.
+In fact, this statement is also true for $C_{4}$.
+However, it is straightforward to see that this property is closed by taking induced subgraphs.
+A natural question would be to characterize the graphs for which this property holds.
+Theorem 1.1 implies the following.
+Corollary 1.2.
+The induced tree polytope and cone of a chordal graphs have IDP.
+[…]
+Figure 1: Sets of induced trees considered in the proof of Theorem 1.3
+Let us denote by $\mathcal{T}(G)$ the polytope defined as the intersection of $\mathcal{C}_{T}(G)$ and the following hyperplane:
+$\{(x,y):x(V)-y(E)=1\}$ (4)
+We now prove that $\mathcal{T}(G)$ characterizes induced tree polytopes in chordal graphs.
+Theorem 1.3.
+A graph $G$ is chordal if and only if $\mathcal{T}(G)$ is binary.
+Proof. $\Rightarrow"$ The extreme points of $\mathcal{T}(G)$ are exactly the scalings of $\mathcal{C}_{T}(G)$’s generators that belong to (4). Theorem 1.1 concludes.
 $\Leftarrow"$ By contradiction, suppose that $G$ is not chordal, and let $C$ induce a hole of size at least 4 in $G$, and $v$ belongs to $C$.
@@
 ∎
-[…]
-Figure 1: Sets of induced trees considered in the proof of Theorem 1.1
+The set of inclusion-wise maximal cliques of chordal graphs is polynomial, implying that (1)–(3) has compact size.
+Optimizing a linear function on the vertex and edge set over $\mathcal{T}(G)$ is doable in polynomial time.
 Corollary 1.4.
-The induced tree polytope of a graph $G$ in the extended variable space is equal to $\mathcal{T}(G)$ if and only if $G$ is chordal.
-The set of inclusion-wise maximal cliques of chordal graphs is polynomial (1)–(3), implying that Optimizing a linear function on the vertex and edge set over $\mathcal{T}(G)$ is doable in polynomial time.
+The induced tree polytope of a graph $G$ is equal to $\mathcal{T}(G)$ if and only if $G$ is chordal. Moreover, $\mathcal{T}(G)$ is described by a linear system of size $O(|V||\mathcal{K}|)$.
 Corollary 1.5.
@@
 2 Induced path polytope of chordal graphs
-We first present a technical lemma concerning edge contraction in chordal graphs exhibited in [dupontbouillard2024contractions].
-After that, we give an exponential description of the induced path polytope of chordal graphs and characterize the inequalities that define facets.
-Finally, we show that the non-dominated such constraints come in a polynomial number.
+We first recall that chordal graphs are closed by edge contraction as exhibited in [dupontbouillard2024contractions]. We use this result to prove a technical graph theory lemma.
+After that, we show that the extended incidence vectors of induced paths of a chordal graph form a Hilbert basis.
+This allows us to give an exponential description of the induced path polytope of chordal graphs and characterize the inequalities that define facets.
+Finally, we show that the non-dominated constraints come in a polynomial number.
 Lemma 2.1 ([dupontbouillard2024contractions]).
-If $G=(V,E)$ is chordal, then $G/F$ also is chordal for any set $F\subseteq E$.
-Given a set of vertices $W\subseteq V$, we denote by $C_{G}(W)$ the vertices complete to $W$ in $G$, we simply write $C(W)$ when clear from context. We denote by $\mathcal{P}(G)$ the following system:
-$y(\delta(w,C(K\cup\{w\})))+2y(\delta(w,K))$ $\leq 2x_{w}$ $\ \forall w\in V,\ \text{ clique }K\subseteq N(w)$ (5)
-$-x_{w}$ $\leq 0$ $\forall w\in V$ (6)
-$-y_{e}$ $\leq 0$ $\forall e\in E$ (7)
-$x(V)-y(E)$ $=1$ (8)
-The next theorem shows that $\mathcal{P}(G)$ describes the induced path polytope of chordal graphs.
-Inequalities (5) express the fact that an induced path contains at most two edges in $\delta(w)$ (for $K=\emptyset$) and the extremitie of two such edges cannot induce a triangle (for $K\neq\emptyset$).
-We refer to Figure 2 for an illustration where the blue edges are the ones involved with coefficient 2 and the green ones are involved with coefficient one.
-At most two green edges belong to a same induced path every blue edge is not contained in any induced path containing some other colored edge.
-Theorem 2.2.
-A graph $G=(V,E)$ is chordal if and only if $\mathcal{P}(G)$ is binary:
-Proof.
-$\Rightarrow$ We proceed by showing that the pointed cone $\mathcal{C}(G):=$(5)–(7) is generated by the extended incidence vectors of induced paths.
-The extreme points of (5)–(8) are scalings of the generators of $\mathcal{C}(G)$ satisfying (8).
-We now show that the set of induced paths’ extended induced path generates $\mathcal{C}(G)$.
-To do so, we proceed by induction on a perfect elimination order of $G$.
-Initialization of the induction is given by the single vertex graph $G$ for which $\mathcal{C}(G)=\mathds{R}_{+}$.
-Induction goes as follows: let $v$ be a simplicial vertex of $G$ and suppose $\mathcal{C}(H)$ to be generated by $H$’s induced path extended incidence vectors for $H=G\setminus v$.
-Let $(x^{*},y^{*})$ be a point of $\mathcal{C}(G)$, $(x^{*},y^{*})_{H}$ is a point of $C(H)$.
-By the induction hypothesis, let $(x^{*},y^{*})_{H}$ be generated by the induced path $\Psi$ with coefficients $\lambda$.
-The proof consists of iteratively modifying $\Psi$ and $\lambda$ so that they yield a conic combination equal to $(x^{*},y^{*})$.
-At each step, we denote by $(x^{\prime},y^{\prime})$ the point given by the current conic combination associated with $\Psi$ and $\lambda$; an induced path $P$ is removed from $\Psi$ if $\lambda_{P}$ is set to 0.
-Each operation preserves the value of $(x^{\prime},y^{\prime})_{H}=(x^{*},y^{*})_{H}$.
-Let us fix an edge $uv\in\delta(v)$, we exhibit a conic combination whose $y_{uv}$ component is equal to $y_{uv}^{*}$, as applying the procedure will preserve the value of $y_{e}$ for the other edge $e\in\delta(v)\setminus\{uv\}$, it can be applied until the conic combination is equal to $(x^{*},y^{*})$ on all components but $x_{v}$.
-Let us consider the following subsets of $\Psi$:
-  • $\Psi^{u}$ is the subset of $\Psi$ whose elements have $u$ as an extremity and do not contain any other element of $N[v]$
-  • $\Psi_{2}^{u}$ is the subset of $\Psi$ having $u$ as an extremity and containing another element of $N(v)$
-  • $\Psi^{u}_{I}$ is the subset of $\Psi$ whose interior contains $u$ and does not contain $v$
-  • $\Psi_{u}^{v}$ is the subset of $\Psi$ containing both $u$ and $v$, note that $v$ is necessarily an extremity of $\Psi_{u}^{v}$’s elements
-  • $\Psi_{\bot}$ contains the remaining elements.
-We refer to Figure 3 for an illustration of the different types of path considered.
-To each element of $\Psi^{u}$, the addition of $v$ yields another induced path which contains $\{u,v\}$.
-This implies that unless $y_{uv}^{\prime}$ is already equal to $y_{uv}^{*}$, a conic combination preserving $(x^{\prime},y^{\prime})_{H}$ and strictly increasing over $y_{uv}$ can be computed.
-Algorithm 2 describes such an operation.
-let $P$ be an element of $\Psi^{u}$
-$\alpha=\min\{y_{uv}^{*}-y_{uv}^{\prime},\lambda_{P}\}$
-$P^{\prime}=P\cup\{v\}$
-$(x^{\prime},y^{\prime})=(x^{\prime},y^{\prime})-\alpha\chi^{P}+\alpha\chi^{P^{\prime}}$
-$\lambda_{P}=\lambda_{P}-\alpha$
-$\lambda_{P^{\prime}}=\lambda_{P^{\prime}}+\alpha$
-Algorithm 2 Reduction of $\Psi^{u}\ \forall u\in N(v)$
-We can now suppose that $\Psi^{u}$ is empty (otherwise equality of the conic combination’s $y_{uv}$ component and $y_{uv}^{*}$ is already achieved) which implies the following:
-$x^{*}_{u}$ $=\lambda(\Psi_{2}^{u})+\lambda(\Psi_{I}^{u})+\lambda(\Psi_{u}^{v})$ (9)
-$y_{uv}^{\prime}$ $=\lambda(\Psi_{u}^{v})$ (10)
-Claim 2.3.
-If $\Psi_{I}^{u}$ is empty, then $y_{uv}^{\prime}\geq y_{uv}^{*}$.
-Proof.
-Suppose that $\Psi^{u}_{I}$ is empty.
-Inequality (5) associated with $u$ and $K=N[v]\setminus\{u\}$ applied to $(x^{*},y^{*})$ yields: $x_{u}^{*}\geq y^{*}(\delta(u,K))$ and the right hand side is larger than or equal to $\lambda(\Psi_{2}^{u})+y_{uv}^{*}$.
-Using equality (9) and (10) we get : $y_{uv}^{\prime}\geq y_{uv}^{*}.$
-Suppose now that $\Psi_{I}^{u}$ is non-empty.
-let $P$ be an element of $\Psi_{I}^{u}$
-let $P_{1}$ and $P_{2}$ be the two different maximal subpath of $P$ ending in $u$.
-$\alpha=\min(\lambda_{P},\lambda_{u,v}$)
-$\lambda_{P}=\lambda_{P}-\alpha$
-Let $i\in\{1,2\}$ be such that $P_{i}\cup\{v\}$ induces a path
-$j=i+1\mod 2$
-$\lambda_{u,v}=\lambda_{u,v}-\alpha$
-$\lambda_{P_{j}}=\lambda_{P_{j}}+\alpha$
-$\lambda_{P_{i}\cup\{v\}}=\lambda_{P_{i}\cup\{v\}}+\alpha$
-Algorithm 3 Removal of $\{u,v\}$ from $\Psi_{u}^{v}$
-Algorithm (3) is used to ensure that each element of $\Psi_{u}^{v}$ intersects $\delta(u)$ twice, such a path always contains $uv$ and another edge in $\delta(u)$.
-When Algorithm (3) cannot be applied, either $\Psi_{I}^{u}$ is empty and Claim (2.3) concludes or $\lambda_{u,v}=0$.
-Claim 2.4.
-If $\Psi_{2}^{u}$ is empty, then $y_{uv}^{\prime}\geq y_{uv}^{*}$.
-Proof.
-Suppose that $\Psi_{2}^{u}$ is empty.
-Inequality (5) associated with $u$ and $K=\emptyset$ applied to $(x^{*},y^{*})$ yields: $2x_{u}\geq y^{*}(\delta(u))$.
-Note that the right hand side is larger than or equal to $2\lambda(\Psi_{I}^{u})+\lambda(\Psi_{u}^{v})+y_{uv}^{*}$ as each $\lambda_{P}$ for $P\in\Psi_{I}^{u}\cup\Psi^{v}_{u}$ contributes two times in $y^{*}(\delta(u))$ and $\lambda(\Psi^{v}_{u})=y_{uv}^{*}$.
-Equalities (9) and (10) yield $y_{uv}^{\prime}\geq y_{uv}^{*}$ and conclude the proof.
-∎
-Suppose now that $\Psi_{2}^{u}$ is nonempty, let us consider the clique $K\subseteq N(u)\cap N(v)$ of vertices contained in at least one element of $\Psi_{2}^{u}$.
-We now consider the partition of $\Psi_{I}^{u}$ into $\Psi_{I_{1}}^{u}$ containing the elements intersecting $K$ and $\Psi_{I_{0}}^{u}$ containing the others.
-Algorithm 4 permits to remove elements from $\Psi_{I_{1}}^{u}$ and adds elements in $\Psi_{u}^{v}$ or $\Psi_{2}^{u}$.
-Let $P$ be in $\Psi_{I_{1}}^{u}$
-Decompose $P$ into $P_{1}$ and $P_{2}$ the two maximal subpath of $P$ ending in $u$
-$\alpha=\min(y_{uv}^{*}-y_{uv}^{\prime},\lambda_{P})$
-$\lambda_{P}=\lambda_{P}-\alpha$
-Let $i\in\{1,2\}$ be such that $P_{i}\cup\{v\}$ induces a path
-$j=i+1\mod 2$
-$\lambda_{P_{i}\cup\{v\}}=\lambda_{P_{i}\cup\{v\}}+\alpha$
-$\lambda_{P_{j}}=\lambda_{P_{j}}+\alpha$
-Algorithm 4 Reduction of $\Psi_{I_{1}}^{u}$
-Using Algorithm 4 permits obtaining a conic combination of induced path such that $y_{uv}^{\prime}=y_{uv}^{*}$, or no element of $\Psi_{I}^{u}$ intersects $K$ that is $\Psi_{I}^{u}=\Psi_{I_{0}}^{u}$.
+If $G=(V,E)$ is chordal, then $G/F$ is also chordal for any set $F\subseteq E$.
+The proof that chordal graphs’ induced path incidence vectors form a Hilbert basis is constructive; therefore, it is necessary to build induced paths with "good properties," and this is what the following lemma stands for.
 Claim 2.2.
@@
 Let us contract all the vertices of $P_{1}\cup P_{2}\setminus\{u_{1},u_{2},u\}$ to a single vertex $w$.
-The subgraph induced by $\{u_{1},u_{2},u,w\}$ contains a cycle of size 4, and $w$ is non-adjacent to $u$ by construction.
+The subgraph induced by $\{u_{1},u_{2},u,w\}$ contains a cycle of size 4, and $w$ is non-adjacent to $u$ as otherwise $P_{1}$ and $P_{2}$ would not induce paths.
 By chordality of $G$ and Lemma (2.1), $u_{2}$ is adjacent to $u_{1}$.
-∎
+To show that a set of vectors forms a Hilbert basis, it can be easier to know the polyhedral description of the cone they generate.
+Therefore, we introduce $\mathcal{C}_{P}(G)$ as the cone defined by the following linear system and show in the next theorem that it is generated by the Hilbert basis formed by the induced paths’ extended incidence vectors.
 […]
 Figure 2: Structure of inequalities (5) for a given vertex $w$ and clique $K$ of its neighborhood.
-Claim (2.5) now permits to decompose the elements of $\Psi_{I_{0}}^{u}$ for which a neighbor of $u$ is not complete to $K$ into their two maximal subpaths ending in $v$.
-We then combine one of these subpaths with an element of $\Psi_{2}^{u}$ to obtain other induced paths.
-Algorithm 5 gives the procedure.
-Note that Algorithm 5 may add elements to $\Psi^{u}$ and $\Psi_{I_{1}}^{u}$ which can be empty again using Algorithm 2 and 4.
-Note that the fact that Algorithm 2 and 4 strictly increase the value of $y_{uv}^{\prime}$ ensures that the sequence of updates on $\lambda$ and $\Psi$ converges.
-Let $P_{0}\in\Psi_{I_{0}}^{u}$ and $P_{2}\in\Psi_{2}^{u}$ be such that the $P_{0}^{\prime}\cup P_{2}$ induces a path where $P_{0}^{\prime}$ is a maximal subpath of $P_{0}$ ending in $u$.
-$\alpha=\min(\lambda_{P_{2}},\lambda_{P_{0}})$
-$\lambda_{P_{2}}=\lambda_{P_{2}}-\alpha$
-$\lambda_{P_{0}}=\lambda_{P_{0}}-\alpha$
-$\lambda_{P_{2}\cup P_{0}^{\prime}}=\lambda_{P_{2}\cup P_{0}^{\prime}}+\alpha$
-$\lambda_{(P_{0}\setminus P_{0}^{\prime})\cup\{v\}}=\lambda_{(P_{0}\setminus P_{0}^{\prime})\cup\{v\}}+\alpha$
-Algorithm 5 Replacing the elements of $\Psi_{I_{0}}^{u}$ so that in the remaining ones the neighbors of $u$ are complete to $K$
-Now we suppose that the elements of $\Psi^{u}_{I_{0}}$ are complete to $K$ (otherwise, $\Psi^{u}_{2}$ is empty and Claim 2.4 concludes).
-Using another time Claim 2.5, we get rid of elements in $\Psi_{u}^{v}$ for which the neighbor of $u$ distinct from $v$ is not complete to $K$ with Algorithm 6.
-Note that Algorithm 6 increments $\lambda_{uv}$ which is then decreased back using Algorithm 3.
-It also increments $\Psi_{I_{1}}^{u}$, which is then decreased back using Algorithm 4 if a path in which a neighbor of $u$ is not complete to $K$ is added.
-Such a sequence of modifications strictly increases $y_{uv}^{\prime}$, ensuring convergence.
-Note that the neighbor of $u$ in $P_{i}$ as defined in Algorithm 3 is now complete to $K$ as we already used Algorithm 5.
-Let $P_{4}$ be in $\Psi_{u}^{v}\setminus\{u,v\}$ and $P_{2}\in\Psi_{2}^{u}$ such that $P_{4}\cup P_{2}\setminus\{v\}$ induces a path
-$\alpha=\min(\lambda_{P_{4}},\lambda_{P_{2}})$
-$\lambda_{P_{4}}=\lambda_{P_{4}}-\alpha$
-$\lambda_{P_{2}}=\lambda_{P_{2}}-\alpha$
-$\lambda_{P_{4}\cup P_{2}\setminus\{v\}}=\lambda_{P_{4}\cup P_{2}\setminus\{v\}}+\alpha$
-$\lambda_{u,v}=\lambda_{uv}+\alpha$.
-Algorithm 6 The neighbors of $u$ in $G[P]$ for $P\in\Psi^{u}$ are complete to $K$
-As $(x^{*},y^{*})$ is a point of $\mathcal{C}(G)$, it satisfies the constraint (5) associated with $K$ and $u$ that is:
-$2x_{u}^{\prime}=2x_{u}^{*}\geq 2y^{*}(\delta(u,K))+y^{*}(\delta(u,C(K\cup\{u\})))$ (11)
-By construction, each element $P_{4}$ of $\Psi_{u}^{v}$ contains both $v$ and another neighbor of $u$ and both are complete to $K$ implying that $\lambda_{P_{4}}$ contributes one time in $y^{*}[\delta(u,C(K\cup\{u\})\setminus\{uv\})]$.
-By construction, $y^{*}(\delta(u,C(K\cup\{u\})))=y^{*}_{uv}+y^{*}[\delta(u,C(K\cup\{u\})\setminus\{uv\})]=y^{*}_{uv}+\lambda(\Psi_{u}^{v})+2\lambda(\Psi_{I}^{u})$ .
+$y(\delta(w,C(K\cup\{w\})))+2y(\delta(w,K))$ $\leq 2x_{w}$ $\ \forall w\in V,\ \text{ clique }K\subseteq N(w)$ (5)
+$-x_{w}$ $\leq 0$ $\forall w\in V$ (6)
+$-y_{e}$ $\leq 0$ $\forall e\in E$ (7)
+Inequalities (5) express the fact that an induced path contains at most two edges in $\delta(w)$ (for $K=\emptyset$), and the extremities of two such edges cannot induce a triangle (for $K\neq\emptyset$).
+We refer to Figure 2 for an illustration where the blue edges are the ones involved with coefficient 2 and the green ones are involved with coefficient 1.
+At most two green edges belong to a same induced path; every blue edge is not contained in any induced path containing some other colored edge.
+This shows that extended incidence vectors of induced paths belong to $\mathcal{C}_{P}(G)$.
+2.1 Induced paths of chordal graphs and Hilbert basis
+In this section, $G$ denotes a chordal graph, $v$ a simplicial vertex, $u$ a neighbor of $v$ and $H$ the graph obtained by removing $uv$ from $G$.
+We denote by $\Psi$ a family of vertex subsets of $G$ inducing paths, and consider $\lambda$ to be an integer vector indexed by the induced paths of $G$.
+The pair $(\Psi,\lambda)$ acts as the pair $(\Gamma,\lambda)$ used in the previous section.
+We prove that the extended incidence vectors of induced paths of chordal graphs form a Hilbert basis.
+Although the proof follows the same inductive scheme as in the case of induced trees, it is substantially more technical.
+In particular, it relies on a refined decomposition of the family $\Psi$.
+We first introduce this decomposition of $\Psi$, then define five operations acting on $(\Psi,\lambda)$.
+For each operation, we establish a corresponding lemma showing that it preserves the property that all elements of $\Psi$ induce paths of $G$, the nonnegativity of $\lambda$ and the value of $\lambda(\Psi)|_{H}:=\sum_{P\in\Psi}\lambda_{P}\xi^{P}|_{H}$.
+Moreover, we precise its impact on the coefficients $\lambda$.
+Finally, these ingredients are combined to complete the proof.
+Let us consider the following partition of $\Psi$:
+  • $\Psi^{u}$ is the subset of $\Psi$ whose elements have $u$ as an extremity and do not contain any other element of $N[v]$
+  • $\Psi_{2}^{u}$ is the subset of $\Psi$ having $u$ as an extremity and containing another element of $N(v)$
+  • $\Psi^{u}_{I}$ is the subset of $\Psi$ whose interior contains $u$ and does not contain $v$
+  • $\Psi_{u}^{v}$ is the subset of $\Psi$ containing both $u$ and $v$, note that $v$ is necessarily an extremity of $\Psi_{u}^{v}$’s elements
+  • $\Psi_{\bot}$ contains the remaining elements.
+[…]
+Figure 3: Types of induced paths considered in Section 2.1
+We refer to Figure 3 for an illustration of the different types of path considered, it may be used as a visual aid to keep track of the definitions .
+We denote by $K_{\Psi}$ the elements of $N(v)\setminus\{u\}$ contained in at least one path in $\Psi_{2}^{u}$:
+$K_{\Psi}:=\{w\in N(v)\setminus\{u\}:\exists P\in\Psi_{2}^{u}\text{ s.t. }w\in P\}$
+$K_{\Psi}$ is considered to be changing accordingly to $\Psi$. We further partition $\Psi_{I}^{u}$ into:
+  • $\Psi_{K_{\Psi}}^{u}$ containing the elements of $\Psi$ intersecting $K_{\Psi}$
+  • $\Psi_{\overline{K}_{\Psi}}^{u}$ containing the others.
+We now define some operations to be applied on ($\Psi$$\lambda)$.
+We say that such an operation is valid if it preserves the integrality and nonnegativity of $\lambda$, if the vertex subsets added to $\Psi$ induce paths, and if it preserves the value of $\lambda(\Psi)|_{H}$.
+In each reduction, $\alpha$ is chosen as an integer satisfying $0<\alpha\leq\min\{\lambda_{P}:P\in\Psi,\ \lambda_{P}\text{ is decreased}\}$.
+This ensures that $\lambda$ stays nonnegative.
+$\mathbf{\Psi^{u}}$-Reduction : let $P$ be an element of $\Psi^{u}$. Increase $\lambda_{P\cup\{v\}}$ and decrease $\lambda_{P}$ by $\alpha$.
+$\mathbf{\Psi_{K_{\Psi}}^{u}}$-Reduction : Let $P$ be an element of $\Psi_{K_{\Psi}}^{u}$.
+$P$ contains the two inclusion-wise maximal distinct subsets $P_{1}$ and $P_{2}$ inducing paths ending in $u$.
+Let $P_{2}$ be the one intersecting $N(v)\setminus u$.
+Increase $\lambda_{P_{2}}$, $\lambda_{P_{1}\cup\{v\}}$ and decrease $\lambda_{P}$ by $\alpha$.
+$\mathbf{\Psi_{\overline{K}_{\Psi}}^{u}}$-Reduction : Let $P$ be an element of $\Psi_{\overline{K}_{\Psi}}^{u}$ such that a neighbor $u_{1}$ of $u$ in $P$ is not complete to $K_{\Psi}$.
+Let $P_{1}$ be its inclusion-wise maximal subset inducing a path ending in $u$ and containing $u_{1}$, and let $P^{\prime}$ be the other inclusion-wise maximal subset inducing a path ending in $u$.
+Let $P_{2}$ be an element of $\Psi_{2}^{u}$ whose element $u_{2}$ in $K_{\Psi}$ is non-adjacent to $u_{1}$.
+Decrease $\lambda_{P}$, $\lambda_{P_{2}}$ and increase $\lambda_{P_{1}\cup P_{2}}$, $\lambda_{P^{\prime}\cup\{v\}}$ by $\alpha$.
+$\mathbf{\Psi_{u}^{v}}$-Reduction : Let $P$ be a path of $\Psi_{u}^{v}$ such that the neighbor $u_{1}$ of $u$ distinct from $v$ is not complete to a vertex $u_{2}$ of $K_{\Psi}$ contained in the path $P_{2}\in\Psi_{2}^{u}$.
+Let $P_{1}$ be the inclusion-wise maximal subset of $P$ inducing a path ending in $u$ and containing $u_{1}$.
+Decrease $\lambda_{P}$, $\lambda_{P_{2}}$ and increase $\lambda_{P_{1}\cup P_{2}}$, $\lambda_{\{u,v\}}$ by $\alpha$.
+$\mathbf{\lambda_{\{u,v\}}}$-Reduction : Let $P$ be an element of $\Psi_{\overline{K}_{\Psi}}^{u}$.
+$P$ contains the two distinct inclusion-wise maximal subset $P_{1}$ and $P_{2}$ both inducing paths ending in $u$.
+Note that $P_{1}\cup\{v\}$ induces a path.
+Decrease $\lambda_{u,v}$ and increase $\lambda_{P_{2}}$, $\lambda_{P_{1}\cup\{v\}}$ by $\alpha$.
+Note that in each of the above operations, $\alpha$ is integer, implying that they preserve integrality of $\lambda$.
+We show that under some conditions, the above reductions are valid.
+It remains to verify that the sets of vertices built induce paths and that the value of $\lambda(\Psi)|_{H}$ is preserved.
+Lemma 2.3.
+Suppose that $\Psi^{u}$ is nonempty. $\mathbf{\Psi^{u}}$-Reduction is valid.
+Proof. Given $\in\Psi^{u}, PP$ ends in $u$ and does not intersect $N(v)\setminus\{u\}$, therefore $P\cup\{v\}$ induces a path. Moreover, $\xi^{P}|_{H}=\xi^{P\cup\{v\}}|_{H}$. ∎
+Lemma 2.4.
+Suppose that $\Psi_{K_{\Psi}}^{u}$ is non-empty. $\mathbf{\Psi_{K_{\Psi}}^{u}}$-Reduction is valid and strictly increments $\lambda(\Psi^{u}_{2})$, $\lambda(\Psi^{v}_{u})$.
+Proof.
+By construction, $P\in\Psi_{K_{\Psi}}^{u}$ is decomposed into its two inclusion-wise maximal subsets $P_{1}$ and $P_{2}$ that induce paths ending in $u$.
+At most one of them intersects $N(v)\setminus\{u\}$, which implies that $P_{1}\cup\{v\}$ induces a path.
+Moreover, $\xi^{P}|_{H}=\xi^{P_{1}\cup\{v\}}|_{H}+\xi^{P_{2}}|_{H}.$ Therefore, $\mathbf{\Psi_{K_{\Psi}}^{u}}$-Reduction is valid and increments both $\lambda_{P_{1}\cup\{v\}}$ and $\lambda_{P_{2}}$.
+∎
+Lemma 2.5.
+Suppose that $\Psi_{\overline{K}_{\Psi}}^{u}$ contains an element $P$ such that a neighbor $u_{1}\in P$ of $u$ is not complete to $K_{\Psi}$, in particular $\Psi_{2}^{u}$ is non empty.
+Then $\mathbf{\Psi_{\overline{K}_{\Psi}}^{u}}$-Reduction is valid.
+Proof.
+Let $P_{1}$ be the inclusion-wise maximal subset of $P$ inducing a path ending in $u$ and containing $u_{1}$, let $P^{\prime}$ be its other subset inducing a path ending in $u$.
+Let $P_{2}$ be an element of $\Psi^{u}_{2}$ such that the neighbor $u_{2}$ of $u$ in $P_{2}$ is non adjacent to $u_{1}$.
+Claim 2.2 implies that $P_{1}\cup P_{2}$ induces a path.
+Moreover, $\xi^{P}|_{H}+\xi^{P_{2}}|_{H}=\xi^{P_{1}\cup P_{2}}|_{H}+\xi^{P^{\prime}\cup\{v\}}|_{H}.$ Therefore $\mathbf{\Psi_{\overline{K}_{\Psi}}^{u}}$-Reduction is valid.
+∎
+Lemma 2.6.
+Suppose that $\Psi_{u}^{v}$ contains an element $P$ such that the neighbor $u_{1}$ of $u$, distinct from $v$, is not complete to a vertex $u_{2}$ of $K_{\Psi}$ contained in the path $P_{2}\in\Psi_{2}^{u}$.
+$\mathbf{\Psi_{u}^{v}}$-Reduction is valid and strictly increases $\lambda_{\{u,v\}}$.
+Proof. Let $P_{1}=P\setminus\{v\}$, by Claim 2.2 $P_{1}\cup P_{2}$ induces a path and $\{u,v\}$ induces a path. Moreover, $\xi^{P}|_{H}+\xi^{P_{2}}|_{H}=\xi^{P_{1}}|_{H}+\xi^{\{u,v\}}|_{H}$. ∎
+Lemma 2.7.
+Suppose that $\Psi_{\overline{K}_{\Psi}}^{u}$ is non empty. $\mathbf{\lambda_{\{u,v\}}}$-Reduction is valid and strictly decreases $\lambda_{\{u,v\}}$.
+Proof.
+Given $P\in\Psi_{\overline{K}_{\Psi}}^{u}$, we consider its two inclusion-wise maximal subsets $P_{1}$ and $P_{2}$ inducing a path ending in $u$.
+$P_{1}\cup\{v\}$ induces a path.
+Moreover, $\xi^{P}|_{H}+\xi^{\{u,v\}}|_{H}=\xi^{P_{1}\cup\{v\}}|_{H}+\xi^{P_{2}}|_{H}$.
+∎
+The proof of the following theorem proceeds by induction over the number of vertices and makes use of the previous definitions and operations.
+Theorem 2.8.
+The extended incidence vectors of induced paths of a chordal graph form a Hilbert basis of $\mathcal{C}_{P}(G)$.
+Proof.
+The statement is trivial for $|V|=1$.
+Induction goes as follows: As $G$ is chordal, it contains a simplicial vertex $v$.
+The graph $G\setminus v$ is chordal and by hypothesis, the induced paths’ extended incidence vectors of $G\setminus v$ form a Hilbert basis of $\mathcal{C}_{P}(G\setminus v)$.
+Let $(x^{*},y^{*})$ belong to $\mathcal{C}_{P}(G)\cap\mathds{Z}^{V\cup E}$, its restriction $(x^{*},y^{*})|_{G\setminus v}$ belongs to $\mathcal{C}_{P}(G\setminus v)\cap\mathds{Z}^{V\cup E}$.
+From now on, $(\Psi,\lambda)$ denotes an arbitrary family of induced paths of $G$ with integer coefficients satisfying $(x’,y’)|_{G\setminus v}=(x^{*},y^{*})|_{G\setminus v}$, where (x,y)(x’,y’) is the point induced by $(\Psi,\lambda)$.
+All operations defined above transform $(\Psi,\lambda)$ into another such family and preserve: integrality of $\lambda$, the fact that the elements of $\Psi$ induce paths of $G$ and $(x’,y’)|_{H}=(x^{*},y^{*})|_{H}$.
+Note that the latter requirement implies that the operations not only preserve the value of $\lambda(\Psi)$ over elements in $G\setminus v$ but also on edges in $\delta(v)\setminus\{uv\}$.
+The goal of the upcoming sequence of reductions is to achieve $(x^{\prime},y^{\prime})=(x^{*},y^{*})$. For this we consider the slackness;
+$\Delta:=\sum_{uv\in\delta(v)}\Delta(u,v)\text{ where }\Delta(u,v)=y_{uv}^{*}-y_{uv}^{\prime}.$
+The proof proceeds by fixing a neighbor $u$ of $v$ and modifying $(\Psi,\lambda)$ until the equality $\Delta(u,v)=0$ holds.
+Since reductions for a fixed $u$ do not affect $\Psi_{w}^{v}$ for $w\neq u$, and that elements of $\Psi^{u}_{w}$ are the only ones contributing to the value of $y_{vw}^{\prime}$, the procedure can be applied independently for each neighbor of $v$.
+An induced path $P$ is removed from $\Psi$ if $\lambda_{P}$ is set to 0.
+Lemmas 2.32.5 give conditions for some valid reductions strictly decreasing $\Delta(u,v)$ to be applied for respectively $\alpha=\min(\Delta(u,v),\lambda_{P})$, $\alpha=\min(\Delta(u,v),\lambda_{P})$, and $\alpha=\min(\lambda_{P},\lambda_{P_{2}},\Delta(u,v))$.
+Since the corresponding reductions strictly decrease either $\Delta(u,v)$, at some point, none of Lemma 2.32.6 apply for the fixed neighbor $u$.
+This implies that, either $\alpha=0$ and $\Delta(u,v)=0$ which ends the proof or the following holds:
+  • $\Psi^{u}$, $\Psi^{u}_{I_{1}}$ both are empty
+  • the neighbors of $u$ in the elements of $\Psi^{u}_{I_{0}}$ are complete to $K_{\Psi}$.
+By construction, the only elements of $\Psi$ which contribute to the value of $y_{uv}^{\prime}$ belong to $\Psi_{u}^{v}$, therefore the following holds :
+$y_{uv}^{\prime}=\lambda(\Psi_{u}^{v}).$ (8)
+As $\Psi^{u}$ is empty, the following holds:
+$x^{*}_{u}=x_{u}^{\prime}=\lambda(\Psi_{2}^{u})+\lambda(\Psi_{\overline{K}_{\Psi}}^{u})+\lambda(\Psi_{u}^{v})\\$ (9)
+Suppose that $\Psi^{u}_{I_{0}}$ is empty.
+As $(x^{*},y^{*})$ belongs to $\mathcal{C}_{P}(G)$, by inequality (5) associated with $u$ and $K=N[v]\setminus\{u\}$ we get: $x_{u}^{*}\geq y^{*}(\delta(u,K))$ and the right hand side is larger than or equal to $\lambda(\Psi_{2}^{u})+y_{uv}^{*}$.
+Using equalities (8) and (9) we get : $y_{uv}^{\prime}\geq y_{uv}^{*}$ and hence $\Delta(u,v)=0$.
+Assume henceforth that $\Psi_{\overline{K}_{\Psi}}^{u}$ is non-empty.
+Suppose $\Psi_{2}^{u}$ to be empty, Inequality (5) associated with $u$ and $K=\emptyset$ applied to $(x^{*},y^{*})$ yields: $2x_{u}\geq y^{*}(\delta(u))$.
+Note that the right hand side is larger than or equal to $2\lambda(\Psi_{I}^{u})+\lambda(\Psi_{u}^{v})+y_{uv}^{*}$ as each $\lambda_{P}$ for $P\in\Psi_{I}^{u}\cup\Psi^{v}_{u}$ contributes two times in $y^{*}(\delta(u))$ and $\lambda(\Psi^{v}_{u})=y_{uv}^{*}$.
+Equalities (9) and (8) yield $y_{uv}^{\prime}\geq y_{uv}^{*}$ and $\Delta(u,v)=0$.
+We may now assume that $\Psi_{2}^{u}$ is nonempty.
+Therefore, $K$ is non-empty and the neighbors of $u$ in the elements of $\Psi_{\overline{K}_{\Psi}}^{u}$ are complete to $K_{\Psi}$.
+Suppose that the conditions of Lemma 2.6 hold, then applying $\Psi_{u}^{v}$-Reduction with $\alpha=\min(\lambda_{P},\lambda_{P_{2}})$ yields an element of $\Psi_{K_{\Psi}}^{u}$, therefore, the conditions of Lemma 2.4 hold and $\Psi_{K_{\Psi}}^{u}$-Reduction applies.
+This sequence of operations strictly decreases $\Delta(u,v)$.
+Suppose that the conditions of Lemma 2.7 hold, applying $\lambda_{u}^{v}$-Reduction with $\alpha=\min(\lambda_{P},\lambda_{u,v}$) yields an element of $\Psi_{u}^{v}$, and one of $\Psi^{u}$.
+Therefore, the conditions of Lemma 2.3 hold, and $\Psi^{u}$-Reduction applies.
+This sequence of operations also decreases $\Delta(u,v)$.
+As $\Psi^{u}$-Reduction, $\Psi_{K_{\Psi}}^{u}$-Reduction, $\Psi_{\overline{K}_{\Psi}}^{u}$-Reduction, and the two latter sequences of reductions strictly decrease $\Delta(u,v)$, we may assume that none of Lemmas 2.32.7 apply or that $\Delta(u,v)=0$.
+Therefore, the neighbors of $u$ in the elements of $\Psi^{v}_{u}$ are complete to $K_{\Psi}$ and $\lambda_{\{u,v\}}=0$.
+As $(x^{*},y^{*})$ is a point of $\mathcal{C}_{P}(G)$, it satisfies the constraint (5) associated with $K_{\Psi}$ and $u$ that is (where $C(S)$ denotes the set of vertices complete to $S$):
+$2x_{u}^{\prime}=2x_{u}^{*}\geq 2y^{*}(\delta(u,K_{\Psi}))+y^{*}(\delta(u,C(K_{\Psi}\cup\{u\})))$ (10)
+By construction, each element $P$ of $\Psi_{u}^{v}$ contains both $v$ and another neighbor of $u$ and both are complete to $K_{\Psi}$.
+This implies that $\lambda_{P}$ contributes two times in $y^{*}[\delta(u,C(K_{\Psi}\cup\{u\}))]$.
+By construction, $y^{*}(\delta(u,C(K_{\Psi}\cup\{u\})))=y^{*}_{uv}+y^{*}[\delta(u,C(K_{\Psi}\cup\{u\})\setminus\{uv\})]=y^{*}_{uv}+\lambda(\Psi_{u}^{v})+2\lambda(\Psi_{\overline{K}_{\Psi}}^{u})$.
 Using equalities (8) and (9) we get:
-$2x_{uv}^{*}-2x_{uv}^{*}\geq 2\lambda(\Psi_{2}^{u})+y^{*}_{uv}+2\lambda(\Psi_{I}^{u})+\lambda(\Psi_{u}^{v})-2\lambda(\Psi_{2}^{u})-2\lambda(\Psi_{I}^{u})-2\lambda(\Psi_{u}^{v})$
+$0\geq y^{*}_{uv}-\lambda(\Psi_{u}^{v})$
 which finally shows that $y_{uv}^{\prime}\geq y_{uv}^{*}$.
-One can repeat this procedure until $y_{uv}^{*}=y_{uv}^{\prime}\ \forall uv\in\delta(v)$.
-Setting $\lambda_{\{v\}}=(x^{*}_{u}-y^{*}(\delta(v)))$ yields the desired conic combination of $(x^{*},y^{*})$ using induced paths’ extended incidence vectors.
+Repeat this procedure until $y_{uv}^{*}=y_{uv}^{\prime}\ \forall uv\in\delta(v)$.
+Setting $\lambda_{\{v\}}=(x^{*}_{v}-y^{*}(\delta(v)))$ for the trivial induced path $\{v\}$ yields the desired integer combination of $(x^{*},y^{*})$ using induced paths’ extended incidence vectors.
+As induced trees of $C_{4}$ coincide with its induced path, Theorem 2.8 is not a characterization of the graph for which the induced path extended incidence vectors have IDP.
+2.2 Extended formulation for the induced path polytope
+In this section, we discuss the consequences of Theorem 2.8 and use it to give a compact description of the induced path polytope of chordal graphs.
+Corollary 2.9.
+The induced path cone and polytope of chordal graphs both have IDP.
+Similarly to what we do for the induced tree polytope, we exhibit a description of the induced path polytope of chordal graphs.
+We denote by $\mathcal{P}(G)$ the polytope defined as the intersection of $\mathcal{C}_{P}(G)$ and the hyperplane:
+$\{(x,y)~|~x(V)-y(E)=1\}.$ (11)
+Theorem 2.10.
+A graph $G=(V,E)$ is chordal if and only if $\mathcal{P}(G)$ is binary.
+Proof.
+$(\Rightarrow)$ The extreme points of $\mathcal{P}(G)$ are scalings of $\mathcal{C}_{P}(G)$’s generators which belong to (11).
+Theorem 2.8 tells that the set of induced paths’ extended incidence vectors generate $\mathcal{C}_{P}(G)$ moreover, they belong to (11), which concludes the proof.
 $(\Leftarrow)$ The proof is by contradiction. The same non binary point exhibited in the proof of Theorem 1.3 also belongs to $\mathcal{P}(G)$. ∎
-[…]
-Figure 3: Induced paths considered in the proof of Theorem 2.2
 Note that inequalities (6) are mandatory only when the graph has isolated vertices.
 If the input graph has no such vertex, then inequalities (5) and (7) ensure that $x$ is positive.
-These constraints have been added to give easy initializations of the induction.
 The proof of Theorem 2.10 shows that the extreme points of $\mathcal{P}(G)$ are induced paths’ extended incidence vectors, which implies the following.
 Corollary 2.11.
-A graph $G$ is chordal if and only if its extended space induced path polytope is equal to $\mathcal{P}(G)$.
-It turns out that not all inequalities (5) define facets, we characterize which do.
+A graph $G$ is chordal if and only if its induced path polytope is equal to $\mathcal{P}(G)$.
+It turns out that not all inequalities (5) define facets; we characterize which ones do.
 Theorem 2.12.
-Given a graph $G=(V,E)$, a vertex $w$, and a clique $K\in N(w)$, the associated inequality (5) defines a facet if and only if no connected component of $\overline{G}[C_{G}(K)]$ is bipartite.
+Given a graph $G=(V,E)$, a vertex $w$, and a clique $K\subseteq N(w)$, the associated inequality (5) defines a facet if and only if no connected component of $\overline{G}[C_{G}(K\cup\{w\})]$ is bipartite.
 Proof.
-$(\Rightarrow)$ Let us consider the connected components $\overline{G}_{1},\dots,\overline{G}_{k}$ of $\overline{G}[C_{G}(K)]$ and suppose by contradiction that for $i\leq\ell$, $\overline{G}_{i}$ is bipartite composed of the stable sets induced by $S_{i}^{1},S_{i}^{2}$.
-By construction, $K_{j}=\bigcup_{i=1}^{\ell}S_{i}^{j}\cup K,\ j=1,2$ is a clique of $G$ such that $C_{G}(K_{j})=\bigcup_{i=\ell+1}^{k}V(\overline{G}_{i})$.
+$(\Rightarrow)$ Let us consider the connected components $\overline{G}_{1},\dots,\overline{G}_{k}$ of $\overline{G}[C_{G}(K\cup\{w\})]$ and suppose by contradiction that for $i\leq\ell$, $\overline{G}_{i}$ is bipartite composed of the stable sets induced by $S_{i}^{1},S_{i}^{2}$.
+By construction, $K_{j}=\bigcup_{i=1}^{\ell}S_{i}^{j}\cup K\cup\{w\},\ j=1,2$ is a clique of $G$ such that $C_{G}(K_{j})=\bigcup_{i=\ell+1}^{k}V(\overline{G}_{i})$.
 By summing $2x_{w}\geq 2y(\delta(w,K_{j}))+y(\delta(w,C(K_{j})))$ for $j=1,2$ we obtain the inequality associated with $K$ and $w$ showing that it is dominated and hence not facet-defining.
-$(\Leftarrow)$ We exhibit |V|+|E|1|V|+|E|-1 linearly independent points satisfying (5) with equality, the fact that the equality constraints submatrix’s rank is at least 1 because of equality 8 concludes:
+$(\Leftarrow)$ We exhibit |V|+|E|1|V|+|E|-1 linearly independent points satisfying (5) with equality, the fact that the equality constraints submatrix’s rank is at least 1 because of equality 11 concludes.
+We consider $K_{w}=K\cup\{w\}$.
+The points are as follows:
   1. 1.
-    $\xi^{v}\ \forall v\in V\setminus\{w\}$
+    $\xi^{\{v\}}\ \forall v\in V\setminus\{w\}$
   2. 2.
-    $\xi^{u,v}\ \forall uv\in E\setminus\delta(w)$
+    $\xi^{\{u,v\}}\ \forall uv\in E\setminus\delta(w)$
   3. 3.
-    $\xi^{u,w}\ \forall uw\in\delta(w,K)$
+    $\xi^{\{u,w\}}\ \forall uw\in\delta(w,K)$
   4. 4.
-    $\xi^{u,w,v}\ \forall uw\in\delta(w)\setminus\delta(K)\setminus\delta(C(K))$ for some vertex $v\in K$ non adjacent to $u$
-The remaining $|\delta(w,C(K))|$ points are built the following way: for each connected components $\overline{G}_{c}$ of $\overline{G}[C(K)]$, let us consider $T_{c}\subseteq E(\overline{G}_{c})$ inducing a subgraph containing an odd cycle and hitting every vertex of $G_{c}$, $T_{c}$ exists by hypothesis and may choosen such that $|T_{c}|=|V(\overline{G}_{c})|$.
-This can be done by starting from an odd cycle and then adding the missing edges for this set to be spanning the vertices.
-Now, match each edge of $\delta(w,C(K))$ to a distinct edge of $T=\bigcup T_{c}$ sharing an extremity, note that $|T|=|C(K)|=|\delta(w,C(K))|$.
-Let $uw$ be in $\delta(w,C(K))$ and $uv$ be its corresponding edge of $T$.
-In $G$, $u$ and $v$ are non adjacent hence $\xi^{u,v,w}$ is the incidence vector of an induced path for which inequality (5) for $w$ and $K$ is tight.
-This construction is necessary as such path will contain two edges of $\delta(w,C(K))$ making it hard to build an indentity of wanted size over these components.
-We refer to Figure 4 for an illustration of the matrix whose rows are the points previously exhibited, and ”?” stands for the submatrices associated with incidence vectors of 3-sized paths and whose value does not need to be explicit for the validity of the proof.
-The rank of $H$ is unkown, however, we know that the restriction of the last $|\delta(w,C(K))|$ points to elements of $G\setminus\delta(w,C(K))$ contains a incidence matrix of a non bipartite graph $T$, by Theorem 2.1 of [GROSSMAN1995213] it has full row rank $|C(K)|$.
-Linear independence is obtained by the fact that the matrix is block triangular with a non-zero diagonal over the points (1)–(4), and the fact that the latter set of points forms a submatrix of rank $|C(K)|$ which are the only ones involving elements of $\delta(w,C(K))$.
+    $\xi^{\{u,w,v\}}\ \forall uw\in\delta(w)\setminus(\delta(K)\cup\delta(C(K_{w})))$ for some vertex $v\in K$ non adjacent to $u$
+The remaining $|\delta(w,C(K_{w}))|$ points are built the following way: for each connected components $\overline{G}_{c}$ of $\overline{G}[C(K_{w})]$, let us consider $T_{c}\subseteq E(\overline{G}_{c})$ inducing a subgraph containing an odd cycle and hitting every vertex of $G_{c}$, $T_{c}$ exists by hypothesis and may choosen such that $|T_{c}|=|V(\overline{G}_{c})|$.
+This can be done by starting from an odd cycle and then adding the missing edges for this set to span the vertices.
+Now, match each edge of $\delta(w,C(K_{w}))$ to a distinct edge of $T=\bigcup T_{c}$ sharing an extremity, note that $|T|=|C(K_{w})|=|\delta(w,C(K_{w}))|$.
+Let $uw$ be in $\delta(w,C(K_{w}))$ and $uv$ be its corresponding edge of $T$.
+In $G$, $u$ and $v$ are non-adjacent, hence $\xi^{u,v,w}$ is the incidence vector of an induced path for which inequality (5) for $w$ and $K$ is tight.
+This construction is necessary as such a path will contain two edges of $\delta(w,C(K_{w}))$, making it impossible to continue a diagonal equal to 1.
+We refer to Figure 4 for an illustration of the matrix whose rows are the points previously exhibited, and "?" stands for the submatrices associated with incidence vectors of 3-sized paths and whose value does not need to be explicit for the validity of the proof.
+$H$ is similar to the incidence matrix of the non bipartite graph induced by $T$, by Theorem 2.1 of [GROSSMAN1995213] it has full row rank $|C(K_{w})|$.
+Full row rank is obtained by the fact that the matrix is block triangular with a non-zero diagonal over the points (1)–(4), and the fact that the last set of points forms a submatrix of rank $|C(K_{w})|$ which are the only ones involving elements of $\delta(w,C(K_{w}))$.
+This implies that the matrix has rank $|V|+|E|$, therefore the inequality is facet defining.
 ∎
@@
 We show that the non-dominated inequalities (5) come in a polynomial number and can be enumerated efficiently.
-Given a fixed vertex $w$, for any two cliques $K\subseteq N(w)$ and $K^{\prime}\subseteq N(w)$, if $K\subseteq K^{\prime}$ then we have $C(K^{\prime})\subseteq C(K)$, for this reason, we say that a clique defines an orbit if there exists no $K^{\prime}$ strictly containing $K$ such that $K\cup C(K)\subseteq K^{\prime}\cup C(K^{\prime})$.
+Given a fixed vertex $w$, for any two cliques $K\subseteq N(w)$ and $K^{\prime}\subseteq N(w)$, let $K_{w}=K\cup\{w\}$, $K_{w}^{\prime}=K^{\prime}\cup\{w\}$.
+If $K\subseteq K^{\prime}$, then we have $C(K_{w}^{\prime})\subseteq C(K_{w})$, for this reason, we say that a clique defines an orbit of $w$ if there exists no $K^{\prime}$ strictly containing $K$ such that $K_{w}\cup C(K_{w})\subseteq K_{w}^{\prime}\cup C(K_{w}^{\prime})$.
 Suppose that a clique $K$ is not orbit-defining, let $K^{\prime}$ be the orbit-defining clique containing $K$, inequality (5) associated with $(w,K)$ is dominated with respect to inequality (5) associated with $(w,K^{\prime})$ and non-negativity constraints.
 Theorem 2.13.
-Given a chordal graph $G$, the set of non-dominated inequalities inequalities (5) has polynomial size.
+Given a chordal graph $G$, the set of non-dominated inequalities (5) has polynomial size.
 Proof.
@@
 Let $K_{H}$ be a maximal clique of $H$.
-It is partitioned into $K=K_{H}\cap N(w)$ and $K_{W}=K_{H}\cap W$.
+It is decomposed into $K=K_{H}\cap N(w)$ and $K_{W}=K_{H}\cap W$.
 We associate with $K_{H}$ a unique orbit-defining clique of $G[N(w)]$.
@@
 ∎
+Since for each vertex $w$, orbit-defining cliques of $G[N(w)]$ are in one-to-one correspondence with the maximal cliques of a chordal graph with at most $2|V|$ vertices, their number is $O(V)$ for fixed $w$.
+Summing over all $w\in V$, the total number of non-dominated inequalities (5) is $O(|V|^{2})$.
 Corollary 2.14.
+There exists an extended formulation for the induced path polytope of chordal graphs with size $O(|V|^{2})$.
 The maximum vertex and weighted induced path problem is solvable in polynomial time in chordal graphs.
-It turns out that the target of Theorems 2.7 and 2.8 differ, as there exist orbit-defining cliques in chordal graphs whose associated inequalities are not facet-defining as given by Figure 5.
+It turns out that the target of Theorems 2.12 and 2.13 differ, as there exist orbit-defining cliques in chordal graphs whose associated inequalities are not facet-defining as given by Figure 5.
 Vertices 1, 2, and 3 form an orbit defining clique such that $C(\{1,2,3\})=\{4,5,6,7\}$ however, the complement of the subgraph induced by $\{4,5,6,7\}$ is a hole of size 4 which is bipartite therefore, Theorem 2.12 implies that its associated inequalities are not facet defining.
@@
 Conclusion
-In this paper, we propose extended formulations for the induced tree and path polytopes of chordal graphs of compact size, implying the polynomial solvability of finding a maximum vertex and edge weighted such induced subgraph in chordal graphs.
-Our proofs rely heavily on the structure of the induced trees and paths cones; it may be interesting to investigate the structures of the linear programs whose constraint sets correspond to incidence vectors of induced paths and trees.
-Acknowledgment
-The author thanks the ANR project ANR-24-CE23-1621 EVARISTE for financial support.
+In this paper, we show that the induced tree and path extended incidence vectors of chordal graphs form a Hilbert basis.
+We use this fact to give a compact linear system of inequalities whose binaricity characterizes chordality.
+It turns out that there exist non-chordal graphs for which induced tree and path extended incidence vectors form a Hilbert basis, which raises the question of characterizing such graphs.
+In general, showing that a set of vectors $\mathcal{V}$ forms a Hilbert basis is used to show that a linear system whose rows are elements of $\mathcal{V}$ is TDI.
+In this paper, we are not considering the programs whose rows are extended induced path or induced tree of chordal graphs.
+Finding some problems that are naturally formulated as integer linear programs whose linear constraint set has such a structure would lead to TDI-ness results.
+Statements and Declarations
+The author thanks the ANR project ANR-24-CE23-1621 EVARISTE for financial support and declares that he has no conflicts of interest.

### v3

@@
 Abstract
-In this article, we give two extended space formulations, respectively, for the induced tree and path polytopes of chordal graphs with vertex and edge variables.
-These formulations are obtained by proving that the induced tree and path extended incidence vectors of chordal graphs form Hilbert basis.
-This also shows that both polytopes have the integer decomposition property in chordal graphs.
-Whereas the formulation for the induced tree polytope is easily seen to have a compact size, the system we provide for the induced path polytope has an exponential number of inequalities.
-We show which of these inequalities define facets and exhibit a superset of the facet-defining ones that can be enumerated in polynomial time.
-We show that for some graphs, the latter superset contains redundant inequalities.
+In this article, we give an extended space formulation for the induced tree polytope and another one for the induced path polytope of chordal graphs, both with vertex and edge variables.
+These formulations are obtained by proving that, in chordal graphs, the extreme points of those polytopes respectively form two Hilbert bases.
+Since these polytopes are binary, it also shows that they have integer decomposition property.
+While the formulation for the induced tree polytope has a polynomial size, the formulation we propose for the induced path polytope involves an exponential number of inequalities.
+We identify which of these inequalities are facet‑defining and exhibit a polynomially enumerable superset containing all of them, although this superset may include redundant inequalities for some graphs.
 As corollaries, we obtain that the problems of finding an induced tree or path maximizing a linear function over the edges and vertices are solvable in polynomial time for the class of chordal graphs.
@@
 Definitions
-All the graphs in this article are simple and connected unless specified.
+All the graphs in this article are simple.
 Given a graph $G=(V,E)$, we denote its complement by $\overline{G}=(V,\overline{E})$, where $\overline{E}=\{e\in\binom{V}{2}:e\notin E\}$.
-We respectively denote by $V(G)$ and $E(G)$ the vertex set and the edge set of $G$.
+We denote by $V(G)$ and $E(G)$ the vertex set and the edge set of $G$, respectively.
 Two vertices $u$ and $v$ are adjacent if $uv\in E(G)$.
-Given a subset of vertices $W\subseteq V$, let $E(W)$ denote the set of edges of $G$ having both endpoints in $W$ and $\delta(W)$ the set of all edges having exactly one endpoint in $W$.
-When $W$ is a singleton $\{w\}$, we will simply write $\delta(w)$.
-We say that the edges in $\delta(w)$ are incident to $w$, and two edges sharing an endpoint are said to be adjacent.
-Given two subsets of vertices $W_{1}$, $W_{2}$, we denote by $\delta(W_{1},W_{2})$ the set of edges having one endpoint in both $W_{1}$ and $W_{2}$.
-For $F\subseteq E$, let $V(F)$ denote the set of vertices incident to any edge of $F$.
+Given a subset of vertices $W\subseteq V$, we denote by $E(W)$ the set of edges of $G$ having both endpoints in $W$, and by $\delta(W)$ the set of all edges having exactly one endpoint in $W$.
+When $W$ is a singleton $\{w\}$, we simply write $\delta(w)$.
+We say that the edges in $\delta(w)$ are incident to $w$.
+Given two subsets of vertices $W_{1}$, $W_{2}$, we denote by $\delta(W_{1},W_{2})$ the set of edges having one endpoint in $W_{1}$ and the other in $W_{2}$.
+For $F\subseteq E$, we denote by $V(F)$ the set of vertices incident to any edge of $F$.
 Given $W\subseteq V$, the graph $G[W]=(W,E(W))$ is the subgraph of $G$ induced by $W$.
 When $H$ is an induced subgraph of $G$, we say that $G$ contains $H$.
+We denote by $G\setminus v$ (resp.
+$G\setminus uv$¨) the obtained from $G$ by removing the vertex $v$ (resp.
+edge $uv$).
 Given a vertex $u\in V(G)$, we denote by $N_{G}(u)=\{w\in V(G):uw\in E(G)\}$ its neighborhood, and by $N_{G}[u]=N_{G}(u)\cup\{u\}$ its closed neighborhood, when clear from context we simply write $N(u)$ and $N[u]$.
@@
 A vertex $u$ complete to $V\setminus\{u\}$ is said to be universal.
-Given a set of vertices $W\subseteq V$, we denote by $C_{G}(W)$ the vertices complete to $W$ in $G$; we simply write $C(W)$ when clear from context.
+Given a set of vertices $W\subseteq V$, we denote by $C_{G}(W)$ the vertices complete to $W$ in $G$; we simply write $C(W)$ when the graph is clear from context.
+A clique is a set of pairwise adjacent vertices.
+A stable set is a set of pairwise nonadjacent vertices.
+A graph whose vertex set can be partitioned into two stable sets is called bipartite.
 A vertex is simplicial if its neighborhood induces a clique.
-An induced path (resp.
+A path (resp.
 hole) is a graph induced by a sequence of vertices $(v_{1},\dots,v_{p})$ whose edge set is $\{v_{i}v_{i+1}:\ i=1,\dots,p-1\}$ (resp.
-$\{v_{i}v_{i+1}:\ i=1,\dots,p-1\}\cup\{v_{1}v_{p}\}$ ).
-Note that sets of vertices inducing cliques with three vertices are considered holes.
-An induced tree is a set of vertices inducing an acyclic graph.
+$\{v_{i}v_{i+1}:\ i=1,\dots,p-1\}\cup\{v_{1}v_{p}\}$ and $p\geq 4$).
+A cycle is a set of edges inducing a connected graph with vertices of degree equal to 2.
+A tree is an acyclic graph.
 A subset of vertices induces a path (resp.
@@
 hole).
+By abuse of definition, we consider that a set of vertices inducing a path is an induced path, similarly, an induced tree is a set of vertices inducing a tree.
 The length of a path or hole is its number of edges.
-A path or hole is called long if its length is at least 4.
 The parity of a path or hole is the parity of its length.
-Given a path induced by the sequence $(w_{1},\dots,w_{k})$, we call $w_{1}$, $w_{k}$ its extremities and the set of vertices $\{w_{2},\dots,w_{k-1}\}$ its interior.
-A tree is a graph with no hole.
+Given a path induced by the sequence $(w_{1},\dots,w_{k})$, we call $w_{1}$, $w_{k}$ its extremities and the set of vertices $\{w_{2},\dots,w_{k-1}\}$ its interior moreover, we say that $P$ ends in $w_{1}$ and $w_{k}$.
 The contraction of an edge $uv$ in $G$ yields a new graph $G/uv$ built from $G$ by deleting $u$, $v$, adding a new vertex $w$, and adding the edges $wz$ for all $z\in N_{G}(u)\cup N_{G}(v)$; parallel edges and loops are removed.
 For $F\subseteq E$, we denote by $G/F$ the graph obtained from $G$ by contracting all the edges in $F$.
-A graph is $k$-chordal if its holes have maximum size $k$; 3-chordal graphs are called chordal.
-Equivalently, chordal graphs are those for which an order $(v_{1},\dots,v_{k})$ of its vertices exists such that $v_{i}$ is simplicial in $G_{i}=G[\{v_{i},\dots,v_{k}\}]$, such an order is called a perfect elimination order.
-Moreover, the maximal cliques of $G$ are among $N_{G_{i}}[v_{i}]$, which shows that they come in a linear number.
-The size of a linear system is its number of rows.
+A graph is $k$-chordal if its holes have maximum size $k$; graphs with no holes are called chordal.
+Equivalently, a graph is chordal if there exists an order $(v_{1},\dots,v_{k})$ of its vertices such that $v_{i}$ is simplicial in $G_{i}=G[\{v_{i},\dots,v_{k}\}]$; such an order is called a perfect elimination order.
+Moreover, the maximal cliques of $G$ are among $\{N_{G_{i}}[v_{i}]:i\in\{1,\dots,k\}\}$, which shows that their number is linear.
 Let $\mathcal{P}=\{x:Ax\leq b\}\subseteq\mathds{R}^{n}$ be a polyhedron, that is, the intersection of finitely many half-spaces.
@@
 Moreover, these facet-defining inequalities are unique up to scalar multiplication.
+The size of a linear system of inequalities is its number of rows.
 A face of $\mathcal{P}$ is a polyhedron contained in $\mathcal{P}$ obtained by setting to equality some inequalities in $Ax\leq b$.
@@
 Each extreme point satisfies $\textrm{dim}~\mathcal{P}$ linearly independent valid inequalities for $\mathcal{P}$ with equality.
-A polyhedron $\mathcal{P}\subseteq\mathds{R}^{n}$ is said to be integer if all its non-empty faces contain an integer point, it is binary if it is bounded and all its extreme points are binary.
-A polytope is a bounded polyhedron, equivalently, it can be defined as the convex hull of a finite number of points $P$ and is denoted by $\textrm{conv}(P)$.
-A cone is a polyhedron described by a linear system with a right hand side equal to 0; it is pointed if it contains no affine space.
-The elements of a cone can be described by conic combinations of finite sets of elements called generators.
-The minimal generating sets of pointed cones are unique up to positive scalar multiplication; we denote the element of such a family by its generators they are in one to one correspondence with its face of dimension 1.
+A polyhedron $\mathcal{P}\subseteq\mathds{R}^{n}$ is said to be integer if all its non-empty faces contain an integer point.
+A polytope is a bounded polyhedron, equivalently, it can be defined as the convex hull of a finite number of points $P$ and is denoted by $\textrm{conv}(P)$, it is binary if all its extreme points are binary.
+A cone is a polyhedron described by a linear system with a right-hand side equal to 0; it is pointed if it contains no affine space.
+Given a set of points $P\subseteq\mathds{R}^{n}$, we denote by $\text{cone}(P)$ the conic hull of $P$’s elements, that is, the minimal cone containing $P$.
+A set of points $P$ is a generating set of the cone $\mathcal{C}$ if $\text{cone}(P)=\mathcal{C}$.
+The minimal generating sets of pointed cones are unique up to positive scalar multiplication; we call the elements of such a set its generators they are in one-to-one correspondence with its faces of dimension 1.
 A polyhedron $\mathcal{P}$ is said to have the integer decomposition property (IDP) if every integer point in the dilation $k\mathcal{P}$ for $k\in\mathds{Z}_{+}$ is obtained as the sum of $k$ elements of $\mathcal{P}$.
 A set of points $X=\{\chi_{1},\dots,\chi_{n}\}$ forms a Hilbert basis if every integer point in the conic hull of $X$ is expressible as an integer nonnegative combination of elements of $X$.
-The Hilbert basis and integer decomposition property are closely related.
-If the extreme points of a polytope $\mathcal{P}$ form a Hilbert basis, then $\mathcal{P}$ has the IDP.
+Hilbert basis and integer decomposition property are closely related.
+If the extreme points of a binary polytope $\mathcal{P}$ form a Hilbert basis, then $\mathcal{P}$ has IDP.
 In this work, the polyhedra we consider are convex hulls of points encoding structured induced subgraphs of an input graph $G=(V,E)$.
-Such an induced subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ (note that $E^{\prime}=E(V^{\prime})$) can be encoded by its natural incidence vector that is $\chi^{V^{\prime}}\in\{0,1\}^{V}$ whose components associated with elements of $V^{\prime}$ are equal to 1 and the others equal to 0, or by its extended incidence vector that is $\xi^{V^{\prime}}=(\chi^{V^{\prime}},\zeta^{E^{\prime}})\in\{0,1\}^{V\cup E}$ where the components of $\zeta^{E^{\prime}}$ associated with elements of $E$ are equal to 1 and the others equal to 0.
-Given a point $(x,y)\in\mathds{R}^{V\cup E}$ we denote by $(x,y)|_{G^{\prime}}$ its restriction to elements of $G^{\prime}$.
-In this paper, given a graph $G$, we consider the following polyhedra:
-  • its induced tree cone $\text{cone}\{\xi^{W}:\text{ $W$ induces a tree of $G$}\}$
-  • its induced tree polytope $\text{conv}\{\xi^{W}:\text{ $W$ induces a tree of $G$}\}$
-  • its induced path cone $\text{cone}\{\xi^{W}:\text{ $W$ induces a path of $G$}\}$
-  • its induced path polytope $\text{conv}\{\xi^{W}:\text{ $W$ induces a path of $G$}\}$
+An induced subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ with $E^{\prime}=E(V^{\prime})$, can be encoded by its incidence vector $\chi^{V^{\prime}}\in\{0,1\}^{V}$ whose components associated with elements of $V^{\prime}$ are equal to 1 and the remaining components equal 0.
+In this article, we consider the extended incidence vector denoted by $\xi^{V^{\prime}}=(\chi^{V^{\prime}},\zeta^{E^{\prime}})\in\{0,1\}^{V}\times\{0,1\}^{E}$ where the components of $\zeta^{E^{\prime}}$ associated with elements of $E$ are equal to 1 and the remaining components equal 0.
+Considering extended descriptions such as proposed here is a common way to obtain compact polyhedral descriptions of combinatorial polytopes [exntededfo].
+In particular, considering extended spaces sometimes allows to obtain compact descriptions, whereas the natural variable may require exponentially many constraints.
+Given a point $(x,y)\in\mathds{R}^{V}\times\mathds{R}^{E}$ we denote by $(x,y)|_{G^{\prime}}$ its restriction to elements of $G^{\prime}$.
+Given a graph $G=(V,E)$, the polytopes we consider in this article are defined as follows:
+  • its induced tree cone $\text{cone}\{\xi^{W}:\text{ $W$ is an induced tree of $G$}\}$
+  • its induced tree polytope $\text{conv}\{\xi^{W}:\text{ $W$ is an induced tree of $G$}\}$
+  • its induced path cone $\text{cone}\{\xi^{W}:\text{ $W$ is an induced path of $G$}\}$
+  • its induced path polytope $\text{conv}\{\xi^{W}:\text{ $W$ is an induced path of $G$}\}$
+Notations
+Here and later, calligraphic letters denote polyhedra.
+Capital letters denote sets of vertices or edges, or graphs (note that $\delta$ is also used for sets of edges).
+Lowercase letters refer to vertices, edges or numbers.
+The symbol $\Gamma$ is reserved for sets of induced trees, while $\Psi$ for sets of induced paths, $\Omega$ for sets of cliques and $\lambda$ refers to vector of integer coefficients.
+Finally, $\Delta$ refers to gaps which we aim to reduce to 0.
 Motivations
-In this paper, we study the problems of finding maximum weight induced trees and paths where both the edges and the vertices have weight, and in particular, filling a gap in polyhedral literature associated with these problems.
+In this article, we study the problems of finding maximum weight induced trees and paths where both the vertices and the edges are weighted.
+In particular, we fill a gap in the polyhedral literature associated with these problems.
 Induced trees
@@
 The problem of computing the maximum size of an induced tree is known to be NP-hard [completePathTree] in general, but it is also the case in bipartite and triangle-free graphs with maximal degree 3 [MATOUSEK2007307].
-Several integer linear programs have been proposed to find maximum weighted such subsets [ILPinducedtree], and extensive computational experiments have been made on general graphs.
+Several integer linear programs have been proposed for the maximum weight induced tree problem [ILPinducedtree] based on different variable spaces, and extensive computational experiments have been made on general graphs.
 Induced paths
@@
 The induced paths of a graph are induced trees, and the type of questions raised in the previous paragraph can be raised considering this structure.
-The authors of [FOX2009494] show that every connected triangle-free graph on $n$ vertices has an induced path of size at least $\sqrt{n}$, and their approach can be generalized to obtain lower bounds for graphs with no clique of given size $k$.
+The authors of [FOX2009494] show that every connected triangle-free graph on $n$ vertices has an induced path of length at least $\sqrt{n}$, and their approach can be generalized to obtain lower bounds for graphs with no clique of size $k$.
 In [esperet], the authors show that every 3-connected planar graph contains an induced path on $\text{log}(|V|)$ vertices.
-In [jfraymond], it is observed that a graph may contain a large path but no large induced path; for this reason, they characterize the sets of edges to be forbidden along a path of size $n$ to force the existence of an induced path of size $log(n)$.
-This result allows them to show that if a graph has a path of length $n$ and no $K_{t}$ as a topological minor, then it contains an induced path of order $(logn)^{\Omega(1/tlog^{2}t)}$.
+In [jfraymond], it is observed that a graph may contain a large path but no large induced path; for this reason, they characterize the sets of edges to be forbidden along a path of length $n$ to force the existence of an induced path of length $log(n)$.
+This result allows them to show that if a graph has a path of length $n$ and no $K_{t}$ as a topological minor, then it contains an induced path of order $log(n)^{\Omega(1/tlog^{2}t)}$.
 Finding a maximum vertex weighted induced path is also NP-hard [completePathTree], but polynomially solvable in $k$-chordal graphs, interval filament, graphs decomposable by clique cutset or by splits whose prime subgraphs are polynomially solvable [GAVRIL2002203], improved in [HANSEN2009135] and on graphs of bounded clique width, asteroidal triple free graphs [inducedpathcliquewifth].
 Besides these combinatorial algorithms, no polyhedral study of the corresponding polytope has been made.
-A related polyhedra has, however, been characterized in [goel2025halfspacerepresentationspathpolytopes]: the polytope whose extreme points are incidence vectors of paths between pairs of leaves in a tree: in trees, induced path and path are the same subgraphs, so the latter paper describes the induced path (between leaves) polytope of trees.
-Finally, extended space integer linear programs for this problem have been proposed [ILPinducedPath, ILPinducedPath2], leading to computational experiments.
+A related polyhedra has however, been considered in [goel2025halfspacerepresentationspathpolytopes]: the polytope whose extreme points in $\{0,1\}^{E}$ encode paths between pairs of leaves in a tree.
+In trees, induced path and path are the same subgraphs, so the latter article describes the induced path (between leaves) polytope of trees with edge variables.
+Finally, extended space integer linear programs have been proposed for the maximum weight induced path [ILPinducedPath, ILPinducedPath2], leading to computational experiments.
 Contributions
-In this paper, we study both the induced tree polytope and the induced path polytope of chordal graphs in two separate sections.
-It turns out that beside being included one in the other, these two polytopes have similar properties.
-Each section starts by respectively showing that the induced path and tree extended incidence vectors of chordal graphs form a Hilbert basis of some cones, which we explicitly describe by linear systems of inequalities.
-We show that intersecting these cones with the correct hyperplane yields the induced tree and path polytope of chordal graphs in extended variable space.
+In this article, we study both the induced tree polytope and the induced path polytopes of chordal graphs in two separate sections.
+It turns out that beside being included one in the other, these two polytopes share similar properties.
+Both following sections follow the same scheme, the first is about induced trees, and the second is about induced path.
+Each section starts by showing that the extended incidence vectors of their corresponding graph structure form a Hilbert basis in chordal graphs.
+As a byproduct, we obtain linear systems of inequalities describing the cones they generate in chordal graphs.
+Finally, we show that intersecting these cones with the correct hyperplane yields the induced tree and path polytope of chordal graphs in extended variable space.
 While the system we provide for the induced tree polytope is easily seen to be compact in chordal graphs, the one we provide for the induced path polytope has an exponential number of inequalities.
@@
 This implies that there exists a compact system describing the induced path polytope of chordal graphs.
-Finally, these results imply that the problems of finding a maximum vertex and edge weight induced tree or path are polynomially solvable in chordal graphs.
+This study allows to solve the problems of finding a maximum vertex and edge weight induced tree or a maximum weight induced path in polynomial time for chordal graphs.
 1 Induced tree polytope of chordal graphs
-In this section, we show that the induced tree’s extended incidence vectors of chordal graphs form a Hilbert basis.
-Then we use this fact to exhibit a linear system describing the induced tree polytope of chordal graphs.
-Given a graph $G$, and a vertex $v$, we denote by $\mathcal{K}^{v}$ the set of maximal cliques containing $v$ and $\mathcal{C}_{T}(G)$ the cone defined by the following system of inequalities:
-$\sum_{u\in K\setminus\{v\}}y_{uv}$ $\leq x_{v}$ $\forall v\in V,\ K\in\mathcal{K}^{v}$ (1)
+In this section, we show that the induced tree’s extended incidence vectors of chordal graphs form a Hilbert basis, and then we use this fact to exhibit a linear system describing the induced tree polytope of chordal graphs.
+1.1 Induced trees of chordal graphs and Hilbert basis
+Given a graph $G$, and a vertex $v$ we denote by $\Omega^{v}$ the set of maximal cliques containing $v$ and by $\mathcal{C}_{T}(G)$ the cone defined by the following system of inequalities:
+$\sum_{u\in K\setminus\{v\}}y_{uv}$ $\leq x_{v}$ $\forall v\in V,\ K\in\Omega^{v}$ (1)
 $-x_{v}$ $\leq 0$ $\forall v\in V$ (2)
 $-y_{e}$ $\leq 0$ $\forall e\in E$ (3)
+For binary points, Inequalities (1) express the fact that no two edges incident to $v$ and contained in a same triangle belong to an induced tree.
 Observe that induced trees’ extended incidence vectors belong to $\mathcal{C}_{T}(G)$.
 Theorem 1.1.
-The induced tree extended incidence vectors of a chordal graph form a Hilbert basis.
-Proof.
-We show that induced trees extended incidence vectors of a chordal graph $G=(V,E)$ form a Hilbert basis of $\mathcal{C}_{T}(G)$.
-As $\mathcal{C}_{T}(G)$ is defined with rational coefficients, each one-dimensional face contains integer points.
-This implies that a Hilbert basis of $\mathcal{C}_{T}(G)$ generates $\mathcal{C}_{T}(G)$.
-We proceed by induction. For initialization, consider the graph $G$ with a single vertex. $\mathcal{C}_{T}(G)$ is the positive half-line trivially generated by the unit vector.
+The induced tree extended incidence vectors of a chordal graph $G$ form a Hilbert basis of $\mathcal{C}_{T}(G)$.
+Proof. We proceed by induction. For initialization, consider the graph $G$ with a single vertex. $\mathcal{C}_{T}(G)$ is the positive half-line trivially generated by the unit vector.
 Induction goes as follows: let $v$ be a simplicial vertex of $G$ and suppose that the integer points of $\mathcal{C}_{T}(G\setminus v)$ are obtained as positive integer combinations of induced trees’ extended incidence vectors.
-Let $(x^{*},y^{*})$ be an integer point of $\mathcal{C}_{T}(G)$, $(x^{*},y^{*})|_{G\setminus v}$ is an integer point of $\mathcal{C}_{T}(G\setminus v)$ by construction.
-By the induction hypothesis, let $(x^{*},y^{*})|_{G\setminus v}$ be generated by the induced trees $\Gamma$ with integer coefficients $\lambda$.
-The proof consists of iteratively modifying $\Gamma$ and $\lambda$ so that they yield a conic combination equal to $(x^{*},y^{*})$.
-At each step, we denote by $(x^{\prime},y^{\prime})$ the point given by the current conic combination associated with $\Gamma$ and $\lambda$.
-For fixed $uv\in\delta(v)$, each operation preserves the value of $(x^{\prime},y^{\prime})|_{H}$ for $H=G\setminus uv$, and preserves nonnegativity and integrality of $\lambda$.
-We define the slackness $\Delta=\sum_{uv\in\delta(v)}\Delta(u,v)$ for $\Delta(u,v)=y_{uv}^{*}-y_{uv}^{\prime}$.
-We consider the following notation $\lambda(\Gamma^{\prime})=\sum_{T\in\Gamma^{\prime}}\lambda_{T}$ for every $\Gamma^{\prime}\subseteq\Gamma$.
-Let us consider the following subsets of $\Gamma$, illustrated in Figure 1 :
+Let $(x^{*},y^{*})$ be an integer point of $\mathcal{C}_{T}(G)$, and $(\lambda,\Gamma)$ an integer nonnegative combination of $G$’s induced trees’ extended incidence vectors satisfying the following properties, where $(x^{\prime},y^{\prime})=\sum_{T\in\Gamma}\lambda_{T}\Gamma_{T}$:
+  1. (i)
+    $(x^{\prime},y^{\prime})\leq(x^{*},y^{*})$
+  2. (ii)
+    $(x^{\prime},y^{\prime})|_{G\setminus v}=(x^{*},y^{*})|_{G\setminus v}$
+  3. (iii)
+    $(\lambda,\Gamma)$ minimize the quantity $\begin{array}[]{l}\\ \Delta(\lambda,\Gamma):=\sum_{e\in\delta(v)}\Delta_{e}(\lambda,\Gamma)\\ \text{where }\Delta_{e}(\lambda,\Gamma):=y_{e}^{*}-y_{e}^{\prime}\ \forall e\in\delta(v)\end{array}$ among the combinations satisfying (i)-(ii).
+$\Delta_{uv}(\lambda,\Gamma)$ corresponds to a Manhattan distance to $(x^{*},y^{*})$.
+An integer nonnegative combination $(\lambda,\Gamma)$ satisfying Conditions (i)-(ii) is given by the induction hypothesis applied on the integer point $(x^{*},y^{*})|_{G\setminus v}$ of $\mathcal{C}_{T}(G\setminus v)$ hence, $(\lambda,\Gamma)$ exists.
+We first show that $\Delta_{uv}(\lambda,\Gamma)=0\ \forall uv\in E$. Fix $uv\in E$ and $H=G\setminus uv$, let us consider the following disjoints subsets of $\Gamma$:
   • $\Gamma^{u}\ \forall u\in N(v)$ containing its elements whose intersection with $N[v]$ equals $\{u\}$.
   • $\Gamma^{u}_{2}\ \forall u\in N(v)$ containing its elements whose intersection with $N(v)$ is $\{u,w\}$ for some $w\in N(v)$.
-  • $\Gamma^{u}_{v}\ \forall u\in N(v)$ containing the elements having $v$ as an extremity and containing $u$
-Note that $u$ is either a leaf or an internal node in the elements of $\Gamma^{u}_{2}$, we provide two different corresponding induced trees in different shades of green, which are the two bottom ones.
-We define the following operation to be applied on $(\Gamma,\lambda)$, which decreases $\Delta$ by an integer value.
-$\mathbf{\Gamma^{u}}$-Reduction: let $T$ be an element of $\Gamma^{u}$.
-Set $\alpha=\min(\lambda_{T},y_{uv}^{*}-y_{uv}^{\prime})$ increase $\lambda_{T\cup\{v\}}$ and decrease $\lambda_{T}$ by $\alpha$.
-By construction, for any $T\in\Gamma^{u}$, $T\cup\{v\}$ induces a tree and that $\mathbf{\Gamma^{u}}$-Reduction preserves integrality of $\lambda$.
-Moreover, $\xi^{T}|_{H}=\xi^{T\cup\{v\}}|_{H}$ implying that $\lambda(\Psi)|_{H}$ is preserved.
-$\mathbf{\Gamma^{u}}$-Reduction applies until $\Gamma^{u}$ is empty or $\Delta(u,v)=0$.
-Suppose that $\Gamma^{u}$ is empty then, $y_{uv}^{*}\leq y_{uv}^{\prime}\ \forall uv\in\delta(v)$. As $(x^{*},y^{*})$ belongs to $\mathcal{C}_{T}(G)$ we have the following:
-$y_{uv}^{*}+\lambda(\Gamma_{2}^{u})=y^{*}_{uv}+y^{*}(\delta(u,N(v)\setminus\{u\}))\leq x_{u}^{*}=x_{u}^{\prime}=\lambda(\Gamma^{u}_{v})+\lambda(\Gamma_{2}^{u})=y_{uv}^{\prime}+\lambda(\Gamma_{2}^{u})$
-The first equality is given by construction, the inequality is given by inequality (1) associated with $u$, and the remaining equalities are obtained by construction.
-This show how $\Delta(u,v)=0$ can be achieved on all edges $uv\in\delta(v)$.
-Finally, the singleton $\{v\}$ can be added to $\Gamma$ and $\lambda_{v}$ set to $x_{v}^{*}-y(\delta(v))^{*}$ to achieve $(x^{\prime},y^{\prime})=(x^{*},y^{*})$ with integer $\lambda$.
+  • $\Gamma^{v}_{u}\ \forall u\in N(v)$ containing the elements having $v$ as an extremity and containing $u$.
+Note that $u$ is either a leaf or an internal node in the elements of $\Gamma^{u}_{2}$.
+We refer to Figure 1 for an illustration of the different types of induced trees considered, it may be used as a visual aid to keep track of the definitions, we provide two different corresponding induced trees in different shades of green; these are the two bottom ones.
+Suppose by contradiction that $\Delta_{uv}(\lambda,\Gamma)$ is nonzero and that $\Gamma^{u}$ is nonempty for some neighbor $u$ of $v$.
+Let $P$ be an element of $\Gamma^{u}$ and set $\alpha=\min(\Delta_{uv}(\lambda,\Gamma),\lambda_{P})$.
+Note that $P\cup\{v\}$ is an induced tree, therefore decreasing $\lambda_{P}$ and increasing $\lambda_{P\cup\{v\}}$ by $\alpha$ yields a nonnegative integer combination of induced paths’ extended incidence vectors of $G$, say $(\lambda^{\prime},\Gamma^{\prime})$.
+As $\xi^{P}|_{G\setminus uv}=\xi^{P\cup\{v\}}|_{G\setminus uv}$, $(\lambda^{\prime},\Gamma^{\prime})$ satisfies Conditions (i)-(ii) and $\Delta_{e}(\lambda,\Gamma)=\Delta_{e}(\lambda^{\prime},\Gamma^{\prime})\ \forall e\in\delta(v)\setminus\{uv\}$.
+Moreover, $\Delta_{uv}(\lambda^{\prime},\Gamma^{\prime})=\Delta_{uv}(\lambda,\Gamma)-\alpha$, Condition (iii) applied to $(\lambda,\Gamma)$ implies that $\alpha=0$.
+Therefore, either $\Gamma^{u}$ is empty or $\Delta_{uv}(\lambda,\Gamma)=0$.
+Suppose that $\Gamma^{u}$ is empty, we now show that $\Delta_{uv}(\lambda,\Gamma)=0$ also holds.
+Given a family $\Gamma^{\prime}$ of induced trees, we denote by $\lambda(\Gamma^{\prime})$ the sum of $\lambda$’s components associated with elements of $\Gamma^{\prime}$.
+As $(x^{*},y^{*})$ belongs to $\mathcal{C}_{T}(G)$ we have the following:
+$y_{uv}^{*}+\lambda(\Gamma_{2}^{u})=y^{*}_{uv}+y^{*}(\delta(u,N(v)\setminus\{u\}))\leq x_{u}^{*}=x_{u}^{\prime}=\lambda(\Gamma^{v}_{u})+\lambda(\Gamma_{2}^{u})=y_{uv}^{\prime}+\lambda(\Gamma_{2}^{u})$
+The first equality is given by the fact that the only elements of $\Gamma$ contributing to $y_{e}^{\prime},$ with $e\in\delta(u,N(v)\setminus\{u\})$ are the ones of $\Gamma^{u}_{2}$ moreover, each element of $\Gamma^{u}_{2}$ contributes to at most one $y_{e}$ for $e\in\delta(u,N(v)\setminus u)$.
+The inequality is given by inequality (1) associated with vertex $u$ and clique $N[v]$ .
+The following equality holds by Condition (ii).
+The two last equalities hold by the facts that the elements of $\Gamma$ contributing to $x_{u}^{\prime}$ are the ones in $\Gamma^{u}\cup\Gamma^{u}_{2}\cup\Gamma^{v}_{u}$, but $\Gamma^{u}$ is supposed to be empty, and that the only elements of $\Gamma$ contributing to $y_{uv}^{\prime}$ are in $\Gamma^{v}_{u}$.
+This shows that $\Delta_{uv}(\lambda,\Gamma)=0$ holds for all edges $uv\in\delta(v)$.
+Finally, the singleton $\{v\}$ can be added to $\Gamma$ and $\lambda_{v}$ set to $x_{v}^{*}-y(\delta(v))^{*}$ to achieve $(x^{\prime},y^{\prime})=(x^{*},y^{*})$ with nonnegative integer coefficients $\lambda$ on induced tree’s extended incidence vectors.
 Note that Theorem 1.1 is not a characterization of the graphs for which the induced tree’s extended incidence vectors form a Hilbert basis.
 In fact, this statement is also true for $C_{4}$.
-However, it is straightforward to see that this property is closed by taking induced subgraphs.
-A natural question would be to characterize the graphs for which this property holds.
+Moreover, it is straightforward to see that this property is closed under taking induced subgraphs.
+A natural question would be to give a forbidden induced subgraph characterization of the graphs for which this property holds.
+A minimal forbidden such graph is given by the complement graph of a hole of length 7.
+1.2 Induced trees of chordal graphs and integral systems
 Theorem 1.1 implies the following.
 Corollary 1.2.
-The induced tree polytope and cone of a chordal graphs have IDP.
+The induced tree polytope of chordal graphs have IDP.
 […]
@@
 $\{(x,y):x(V)-y(E)=1\}$ (4)
-We now prove that $\mathcal{T}(G)$ characterizes induced tree polytopes in chordal graphs.
 Theorem 1.3.
 A graph $G$ is chordal if and only if $\mathcal{T}(G)$ is binary.
-Proof. $\Rightarrow"$ The extreme points of $\mathcal{T}(G)$ are exactly the scalings of $\mathcal{C}_{T}(G)$’s generators that belong to (4). Theorem 1.1 concludes.
-$\Leftarrow"$ By contradiction, suppose that $G$ is not chordal, and let $C$ induce a hole of size at least 4 in $G$, and $v$ belongs to $C$.
-By setting $x_{w}=1\ \forall w\in V\setminus v$, $y_{e}=1\ \forall e\in E(C)$ and $x_{v}=2$ and 0 on all other components, we obtain a point which belongs to $\mathcal{T}(G)$ and is not included in the hypercube implying that $\mathcal{T}(G)$ is itself not contained in the hypercube and hence not binary.
+Proof.
+$\Rightarrow"$ Suppose $G$ to be chordal.
+The extreme points of $\mathcal{T}(G)$ are exactly the scalings of $\mathcal{C}_{T}(G)$’s generators which belong to (4).
+As $\mathcal{C}_{T}(G)$ is defined with rational coefficients, each of its one-dimensional face contains integer points.
+This implies that a Hilbert basis of $\mathcal{C}_{T}(G)$ generates $\mathcal{C}_{T}(G)$.
+Therefore, Theorem 1.1 gives that induced trees extended incidence vectors of a chordal graph $G=(V,E)$ generate $\mathcal{C}_{T}(G)$.
+Extended incidence vectors of induced trees belong to (4) and hence are extreme points of $\mathcal{T}(G)$.
+It remains to show that $\mathcal{T}(G)$ is bounded.
+For each vertex $v$ of $G$, there exists a perfect elimination ordering ending in $v$, say $(v_{1},\dots,v_{n},v)$ (it is a consequence of the fact that minimal separators are cliques [Dirac1961]).
+By definition of perfect elimination orderings, $N(v_{i})\cap\{v_{i+1},\dots,v_{n},v\}$ is a clique and the following inequalities are valid for $\mathcal{C}_{T}(G)$ :
+$y(\delta(v_{i},\{v_{i+1},\dots,v_{n},v\}))\leq x_{v_{i}}\quad\forall i\in\{1,\dots,n\}.$
+Therefore, summing the previous family of inequalities with x(V)y(E)=1x(V)-y(E)=1 yields $x_{v}\leq 1$. Moreover, $y$ variables are sandwiched between 0 and $x$ variables.
+$\Leftarrow"$ By contradiction, suppose that $G$ is not chordal, and let $H$ induce a hole of length at least 4 in $G$, and $v$ belongs to $H$.
+By setting $x_{w}=1\ \forall w\in V\setminus v$, $y_{e}=1\ \forall e\in E(H)$, $x_{v}=2$ and 0 on all other components, we obtain a point $(x,y)$ which belongs to $\mathcal{T}(G)$ and is not included in the hypercube.
+This implies that $\mathcal{T}(G)$ is itself not contained in the hypercube and hence not binary.
 ∎
-The set of inclusion-wise maximal cliques of chordal graphs is polynomial, implying that (1)–(3) has compact size.
-Optimizing a linear function on the vertex and edge set over $\mathcal{T}(G)$ is doable in polynomial time.
+The set of inclusion-wise maximal cliques of chordal graphs being polynomial (1)–(3) has compact size.
+Therefore, optimizing a linear function on the vertex and edge set over $\mathcal{T}(G)$ is doable in polynomial time.
 Corollary 1.4.
@@
 2 Induced path polytope of chordal graphs
-We first recall that chordal graphs are closed by edge contraction as exhibited in [dupontbouillard2024contractions]. We use this result to prove a technical graph theory lemma.
-After that, we show that the extended incidence vectors of induced paths of a chordal graph form a Hilbert basis.
-This allows us to give an exponential description of the induced path polytope of chordal graphs and characterize the inequalities that define facets.
+We first recall that chordal graphs are closed by edge contraction as exhibited in [dupontbouillard2024contractions] and use this result to prove a technical lemma.
+Our proof of the fact that induced paths’ extended incidence vectors of chordal graphs form Hilbert bases follows a similar scheme to the one of the previous section.
+We define $(\lambda,\Psi)$ in a similar way to what we did with $(\lambda,\Gamma)$ in the proof of Theorem 1.1.
+However, the proof of the analogous result for induced path requires a closer study of $(\lambda,\Psi)$.
+Therefore, we first exhibit a set of properties of $(\lambda,\Psi)$ in the form of lemmas before diving into the proof itself.
+In the second section, we deduce an exponential size description of the induced path polytope of chordal graphs and characterize the inequalities that define facets.
 Finally, we show that the non-dominated constraints come in a polynomial number.
@@
 If $G=(V,E)$ is chordal, then $G/F$ is also chordal for any set $F\subseteq E$.
-The proof that chordal graphs’ induced path incidence vectors form a Hilbert basis is constructive; therefore, it is necessary to build induced paths with "good properties," and this is what the following lemma stands for.
+The proof that chordal graphs’ induced path incidence vectors form a Hilbert basis requires to "reorganize" the induced path of a conic combination.
+We mean by reorganize that we build a new induced path having desirable properties from other induced path.
+This is what stands for the following lemma.
 Claim 2.2.
-Given $P_{1},P_{2}$ be two induced paths of a chordal graph having a similar extremity $u$, such that the neighbor $u_{1}$ of $u$ in $P_{1}$ is different from the neighbor $u_{2}$ of $u$ in $P_{2}$.
+Given $P_{1},P_{2}$ be two induced paths of a chordal graph having a similar extremity $u$, and such that the neighbor $u_{1}$ of $u$ in $P_{1}$ is different from the neighbor $u_{2}$ of $u$ in $P_{2}$.
 Either $u_{1}$ is adjacent to $u_{2}$ or $P_{2}\cup P_{1}$ induces a path.
-Proof. By contradiction, suppose that $P_{2}\cup P_{1}$ does not induce a path. Let $w_{1}$ be the other element of $P_{1}$ adjacent to $u_{1}$, by hypothesis, $w_{1}\neq u_{2}$.
-As $P_{1}$ and $P_{2}$ have an extremity in common but their union does not induce a path, either they intersect in some other vertex than $u$ or there exists an edge going from $P_{1}\setminus u$ to $P_{2}\setminus u$.
-In both situations, either $u_{2}$ is adjacent to $u_{1}$ or there exist a cycle $C\subseteq P_{1}\cup P_{2}\textasciidieresis$ containing $\{u,u_{1},w_{1},u_{2}\}$.
-Let us contract all the vertices of $P_{1}\cup P_{2}\setminus\{u_{1},u_{2},u\}$ to a single vertex $w$.
-The subgraph induced by $\{u_{1},u_{2},u,w\}$ contains a cycle of size 4, and $w$ is non-adjacent to $u$ as otherwise $P_{1}$ and $P_{2}$ would not induce paths.
-By chordality of $G$ and Lemma (2.1), $u_{2}$ is adjacent to $u_{1}$.
+Proof.
+By contradiction, suppose that $P_{2}\cup P_{1}$ does not induce a path and that $u_{1}$ is not adjacent to $u_{2}$.
+Let $w_{1}$ be the other element of $P_{1}$ adjacent to $u_{1}$, by hypothesis, $w_{1}\neq u_{2}$.
+As $P_{1}$ and $P_{2}$ have an extremity in common but their union does not induce a path, either their intersection contains $u$ and another vertex, or there exists an edge going from $P_{1}\setminus u$ to $P_{2}\setminus u$.
+In both situations, either $u_{2}$ is adjacent to $u_{1}$ or there exist a cycle with vertex set $C\subseteq P_{1}\cup P_{2}\textasciidieresis$ containing $\{u,u_{1},w_{1},u_{2}\}$.
+Let us contract all the edges in $E(P_{1}\cup P_{2}\setminus\{u_{1},u_{2},u\})$ to a single vertex $w$.
+The subgraph induced by $\{u_{1},u_{2},u,w\}$ contains a cycle of length 4, and $w$ is non-adjacent to $u$ as otherwise $P_{1}$ and $P_{2}$ would not induce paths.
+By chordality of $G$ and Lemma 2.1, $u_{2}$ is adjacent to $u_{1}$, a contradiction.
 To show that a set of vectors forms a Hilbert basis, it can be easier to know the polyhedral description of the cone they generate.
-Therefore, we introduce $\mathcal{C}_{P}(G)$ as the cone defined by the following linear system and show in the next theorem that it is generated by the Hilbert basis formed by the induced paths’ extended incidence vectors.
+Therefore, we introduce $\mathcal{C}_{P}(G)$ as the cone defined by the following linear system.
+$y\Biggl(\delta\Bigl(w,C(K\cup\{w\})\Bigr)\Biggr)+2y\Biggl(\delta\Bigl(w,K\Bigr)\Biggr)$ $\leq 2x_{w}$ $\ \forall w\in V,\ \text{ clique }K\subseteq N(w)$ (5)
+$-x_{w}$ $\leq 0$ $\forall w\in V$ (6)
+$-y_{e}$ $\leq 0$ $\forall e\in E$ (7)
 […]
 Figure 2: Structure of inequalities (5) for a given vertex $w$ and clique $K$ of its neighborhood.
-$y(\delta(w,C(K\cup\{w\})))+2y(\delta(w,K))$ $\leq 2x_{w}$ $\ \forall w\in V,\ \text{ clique }K\subseteq N(w)$ (5)
-$-x_{w}$ $\leq 0$ $\forall w\in V$ (6)
-$-y_{e}$ $\leq 0$ $\forall e\in E$ (7)
-Inequalities (5) express the fact that an induced path contains at most two edges in $\delta(w)$ (for $K=\emptyset$), and the extremities of two such edges cannot induce a triangle (for $K\neq\emptyset$).
+For binary points, Inequalities (5) express the fact that an induced path contains at most two edges in $\delta(w)$ (for $K=\emptyset$), and the extremities of two such edges cannot induce a triangle (for $K\neq\emptyset$).
 We refer to Figure 2 for an illustration where the blue edges are the ones involved with coefficient 2 and the green ones are involved with coefficient 1.
-At most two green edges belong to a same induced path; every blue edge is not contained in any induced path containing some other colored edge.
+At most two green edges belong to a same induced path; given a blue edge $e$, for any other colorful edge $e^{\prime}$, there exists a triangle containing both, therefore no induced path $P$ verifies $\{e,e^{\prime}\}\subseteq E(P)$.
 This shows that extended incidence vectors of induced paths belong to $\mathcal{C}_{P}(G)$.
+Observation 2.3.
+The induced paths’ extended incidence vectors of a chordal graph $G$ belong to $\mathcal{C}_{P}(G)$.
 2.1 Induced paths of chordal graphs and Hilbert basis
-In this section, $G$ denotes a chordal graph, $v$ a simplicial vertex, $u$ a neighbor of $v$ and $H$ the graph obtained by removing $uv$ from $G$.
-We denote by $\Psi$ a family of vertex subsets of $G$ inducing paths, and consider $\lambda$ to be an integer vector indexed by the induced paths of $G$.
-The pair $(\Psi,\lambda)$ acts as the pair $(\Gamma,\lambda)$ used in the previous section.
-We prove that the extended incidence vectors of induced paths of chordal graphs form a Hilbert basis.
-Although the proof follows the same inductive scheme as in the case of induced trees, it is substantially more technical.
-In particular, it relies on a refined decomposition of the family $\Psi$.
-We first introduce this decomposition of $\Psi$, then define five operations acting on $(\Psi,\lambda)$.
-For each operation, we establish a corresponding lemma showing that it preserves the property that all elements of $\Psi$ induce paths of $G$, the nonnegativity of $\lambda$ and the value of $\lambda(\Psi)|_{H}:=\sum_{P\in\Psi}\lambda_{P}\xi^{P}|_{H}$.
-Moreover, we precise its impact on the coefficients $\lambda$.
-Finally, these ingredients are combined to complete the proof.
-Let us consider the following partition of $\Psi$:
+This section is dedicated to the proof of the following conjecture.
+Conjecture 2.4.
+The induced path extended incidence vectors of a chordal graph $G$ form a Hilbert basis of $\mathcal{C}_{P}(G)$.
+We prove Conjecture 2.4 by induction over the vertex set of any chordal graph $G$.
+Note that the initilization is given by the single vertex graph for which $\mathcal{C}_{P}(G)$ is the positive half line trivially generated by a unit vector.
+The induction proof consists in considering an integer point $(x^{*},y^{*})$ in $\mathcal{C}_{P}(G)$, a simplicial vertex $v$ of $G$ (which exists by chordality of $G$) and a nonnegative integer combination of $G$’s induced paths’ extended vectors say, the induced path $\Psi$ with coefficient $\lambda$ satisfying the following properties where $(x^{\prime},y^{\prime})=\sum_{P\in\Psi}\lambda_{P}\xi^{P}$:
+  1. (i)
+    $(x^{\prime},y^{\prime})\leq(x^{*},y^{*})$
+  2. (ii)
+    $(x^{\prime},y^{\prime})|_{G\setminus v}=(x^{*},y^{*})|_{G\setminus v}$
+  3. (iii)
+    $(\lambda,\Psi)$ minimize the quantity $\begin{array}[]{l}\\ \Delta(\lambda,\Psi):=\sum_{e\in\delta(v)}\Delta_{e}(\lambda,\Psi)\\ \text{where }\Delta_{e}(\lambda,\Psi):=y_{e}^{*}-y_{e}^{\prime}\ \forall e\in\delta(v)\end{array}$ among the ones that satisfy Conditions (i)-(ii).
+Note that $(x^{*},y^{*})|_{G\setminus v}$ belong to $\mathcal{C}_{P}(G\setminus v)$, therefore the induction hypothesis implies that $(x^{*},y^{*})|_{G\setminus v}$ is obtained as an integer conic combination of induced paths of $G\setminus v$.
+This combination, say $(\Psi^{\prime},\lambda^{\prime})$, satisfies Conditions (i)-(ii) but is somehow the worst in terms of Condition (iii) as $\Delta(\lambda^{\prime},\Psi^{\prime})=y^{*}(\delta(v))$.
+Nevertheless, this proves that $(\lambda,\Psi)$ exists.
+For the sake of simplicity, we consider that when $\lambda_{P}=0$, $P$ does not belong to $\Psi$.
+We first propose a decomposition of $\Psi$, and in the form of lemmas, we exhibit some properties satisfied by $(\lambda,\Psi)$. Given $u\in N(v)$, let us decompose $\Psi$ into the following subsets:
   • $\Psi^{u}$ is the subset of $\Psi$ whose elements have $u$ as an extremity and do not contain any other element of $N[v]$
-  • $\Psi_{2}^{u}$ is the subset of $\Psi$ having $u$ as an extremity and containing another element of $N(v)$
-  • $\Psi^{u}_{I}$ is the subset of $\Psi$ whose interior contains $u$ and does not contain $v$
-  • $\Psi_{u}^{v}$ is the subset of $\Psi$ containing both $u$ and $v$, note that $v$ is necessarily an extremity of $\Psi_{u}^{v}$’s elements
+  • $\Psi_{2}^{u}$ is the subset of $\Psi$ whose elements have $u$ as an extremity and contain another vertex $w$ in $N(v)$
+  • $\Psi_{I_{2}}^{u}$ is the subset of $\Psi$ whose element’s interior contain $u$ and another element $w$ of $N(v)$.
+  • $\Psi_{I_{1}}^{u}$ is the subset of $\Psi$ whose elements’s interior contain $u$ but no other vertex of $N(v)$
+  • $\Psi_{u}^{v}$ is the subset of $\Psi$ whose elements contain both $u$ and $v$, note that $v$ is necessarily an extremity of $\Psi_{u}^{v}$’s elements
   • $\Psi_{\bot}$ contains the remaining elements.
@@
 Figure 3: Types of induced paths considered in Section 2.1
+We denote by $K_{2}^{u}$ the subset of $N(v)\setminus\{u\}$ contained in at least one path in $\Psi_{2}^{u}$, that is:
+$K_{2}^{u}:=\{w\in N(v)\setminus\{u\}:\exists P\in\Psi_{2}^{u}\text{ s.t. }w\in P\}.$
+As the only induced path of $\Psi$ containing the edge $uv$ belong to $\Psi^{v}_{u}$ the following holds:
+$y_{uv}^{\prime}=\lambda(\Psi_{u}^{v}).$ (8)
 We refer to Figure 3 for an illustration of the different types of path considered, it may be used as a visual aid to keep track of the definitions.
-We denote by $K_{\Psi}$ the elements of $N(v)\setminus\{u\}$ contained in at least one path in $\Psi_{2}^{u}$:
-$K_{\Psi}:=\{w\in N(v)\setminus\{u\}:\exists P\in\Psi_{2}^{u}\text{ s.t. }w\in P\}$
-$K_{\Psi}$ is considered to be changing accordingly to $\Psi$. We further partition $\Psi_{I}^{u}$ into:
-  • $\Psi_{K_{\Psi}}^{u}$ containing the elements of $\Psi$ intersecting $K_{\Psi}$
-  • $\Psi_{\overline{K}_{\Psi}}^{u}$ containing the others.
-We now define some operations to be applied on ($\Psi$$\lambda)$.
-We say that such an operation is valid if it preserves the integrality and nonnegativity of $\lambda$, if the vertex subsets added to $\Psi$ induce paths, and if it preserves the value of $\lambda(\Psi)|_{H}$.
-In each reduction, $\alpha$ is chosen as an integer satisfying $0<\alpha\leq\min\{\lambda_{P}:P\in\Psi,\ \lambda_{P}\text{ is decreased}\}$.
-This ensures that $\lambda$ stays nonnegative.
-$\mathbf{\Psi^{u}}$-Reduction : let $P$ be an element of $\Psi^{u}$. Increase $\lambda_{P\cup\{v\}}$ and decrease $\lambda_{P}$ by $\alpha$.
-$\mathbf{\Psi_{K_{\Psi}}^{u}}$-Reduction : Let $P$ be an element of $\Psi_{K_{\Psi}}^{u}$.
-$P$ contains the two inclusion-wise maximal distinct subsets $P_{1}$ and $P_{2}$ inducing paths ending in $u$.
-Let $P_{2}$ be the one intersecting $N(v)\setminus u$.
-Increase $\lambda_{P_{2}}$, $\lambda_{P_{1}\cup\{v\}}$ and decrease $\lambda_{P}$ by $\alpha$.
-$\mathbf{\Psi_{\overline{K}_{\Psi}}^{u}}$-Reduction : Let $P$ be an element of $\Psi_{\overline{K}_{\Psi}}^{u}$ such that a neighbor $u_{1}$ of $u$ in $P$ is not complete to $K_{\Psi}$.
-Let $P_{1}$ be its inclusion-wise maximal subset inducing a path ending in $u$ and containing $u_{1}$, and let $P^{\prime}$ be the other inclusion-wise maximal subset inducing a path ending in $u$.
-Let $P_{2}$ be an element of $\Psi_{2}^{u}$ whose element $u_{2}$ in $K_{\Psi}$ is non-adjacent to $u_{1}$.
-Decrease $\lambda_{P}$, $\lambda_{P_{2}}$ and increase $\lambda_{P_{1}\cup P_{2}}$, $\lambda_{P^{\prime}\cup\{v\}}$ by $\alpha$.
-$\mathbf{\Psi_{u}^{v}}$-Reduction : Let $P$ be a path of $\Psi_{u}^{v}$ such that the neighbor $u_{1}$ of $u$ distinct from $v$ is not complete to a vertex $u_{2}$ of $K_{\Psi}$ contained in the path $P_{2}\in\Psi_{2}^{u}$.
-Let $P_{1}$ be the inclusion-wise maximal subset of $P$ inducing a path ending in $u$ and containing $u_{1}$.
-Decrease $\lambda_{P}$, $\lambda_{P_{2}}$ and increase $\lambda_{P_{1}\cup P_{2}}$, $\lambda_{\{u,v\}}$ by $\alpha$.
-$\mathbf{\lambda_{\{u,v\}}}$-Reduction : Let $P$ be an element of $\Psi_{\overline{K}_{\Psi}}^{u}$.
-$P$ contains the two distinct inclusion-wise maximal subset $P_{1}$ and $P_{2}$ both inducing paths ending in $u$.
-Note that $P_{1}\cup\{v\}$ induces a path.
-Decrease $\lambda_{u,v}$ and increase $\lambda_{P_{2}}$, $\lambda_{P_{1}\cup\{v\}}$ by $\alpha$.
-Note that in each of the above operations, $\alpha$ is integer, implying that they preserve integrality of $\lambda$.
-We show that under some conditions, the above reductions are valid.
-It remains to verify that the sets of vertices built induce paths and that the value of $\lambda(\Psi)|_{H}$ is preserved.
 Lemma 2.5.
-Suppose that $\Psi^{u}$ is nonempty. $\mathbf{\Psi^{u}}$-Reduction is valid.
-Proof. Given $\in\Psi^{u}, PP$ ends in $u$ and does not intersect $N(v)\setminus\{u\}$, therefore $P\cup\{v\}$ induces a path. Moreover, $\xi^{P}|_{H}=\xi^{P\cup\{v\}}|_{H}$. ∎
+Given $u\in N(v)$, either $\Delta_{uv}(\Psi,\lambda)=0$ or $\Psi^{u}=\emptyset$.
+Proof.
+By contradiction let $P$ be an element of $\Psi^{u}$ and $\alpha=\min\Bigl(\Delta_{uv}(\lambda,\Psi),\lambda_{P}\Bigr)$.
+Decreasing $\lambda_{P}$ and increasing $\lambda_{P\cup\{v\}}$ by $\alpha$ yields a new integer conic combination $(\lambda^{\prime},\Psi^{\prime})$, of induced paths’ extended incidence vectors.
+Moreover, as $\xi^{P}|_{G\setminus uv}=\xi^{P\cup v}|_{G\setminus uv}$, we have that $\Bigl(\sum_{P\in\Psi^{\prime}}\lambda_{P}^{\prime}\xi^{P}\Bigr)\Bigm\lvert_{G\setminus uv}=(x^{\prime},y^{\prime})|_{G\setminus uv}$.
+Therefore $(x^{\prime},y^{\prime})$ satisfies Conditions (i)-(ii), $\Delta_{e}(\lambda^{\prime},\Psi^{\prime})=\Delta_{e}(\lambda,\Psi)\ \forall e\in\delta(v)\setminus uv$ and $\Delta(\lambda^{\prime},\Psi^{\prime})=\Delta(\lambda,\Psi)-\alpha$.
+By Condition (iii) applied to $(\lambda,\Psi)$ we obtain $\alpha=0$ and as $\lambda_{P}>0$ we finally obtain $\Delta_{uv}(\Psi,\lambda)=0$.
+∎
+By Lemma 2.5, if $\Delta_{uv}(\Psi,\lambda)$ is nonzero, we have $\Psi^{u}=\emptyset$. Therefore, the following inequality holds and will be used in the upcoming results:
+$x^{*}_{u}=x_{u}^{\prime}=\lambda(\Psi_{2}^{u})+\lambda(\Psi_{I_{2}}^{u})+\lambda(\Psi_{I_{1}}^{u})+\lambda(\Psi_{u}^{v}).\\$ (9)
 Lemma 2.6.
-Suppose that $\Psi_{K_{\Psi}}^{u}$ is non-empty. $\mathbf{\Psi_{K_{\Psi}}^{u}}$-Reduction is valid and strictly increments $\lambda(\Psi^{u}_{2})$, $\lambda(\Psi^{v}_{u})$.
+Given $u\in N(v)$, either $\Delta_{uv}(\lambda,\Psi)=0$ or $\Psi^{u}_{I_{1}}$ is nonempty.
 Proof.
-By construction, $P\in\Psi_{K_{\Psi}}^{u}$ is decomposed into its two inclusion-wise maximal subsets $P_{1}$ and $P_{2}$ that induce paths ending in $u$.
-At most one of them intersects $N(v)\setminus\{u\}$, which implies that $P_{1}\cup\{v\}$ induces a path.
-Moreover, $\xi^{P}|_{H}=\xi^{P_{1}\cup\{v\}}|_{H}+\xi^{P_{2}}|_{H}.$ Therefore, $\mathbf{\Psi_{K_{\Psi}}^{u}}$-Reduction is valid and increments both $\lambda_{P_{1}\cup\{v\}}$ and $\lambda_{P_{2}}$.
+Suppose that $\Delta_{uv}(\lambda,\Psi)\neq 0$ and that $\Psi^{u}_{I_{1}}=\emptyset$.
+As $(x^{*},y^{*})$ belongs to $\mathcal{C}_{P}(G)$, by inequality (5) associated with $u$ and $K=N[v]\setminus\{u\}$ we get: $x_{u}^{*}\geq y^{*}(\delta(u,K))$ and the right-hand side is equal to $\lambda(\Psi_{2}^{u})+\lambda(\Psi_{I_{2}}^{u})+y_{uv}^{*}$ by definition of $(\lambda,\Psi)$.
+Using equalities (8) and (9) we get : $y_{uv}^{\prime}\geq y_{uv}^{*}$ and hence $\Delta_{uv}(\lambda,\Psi)=0$, a contradiction.
 ∎
 Lemma 2.7.
-Suppose that $\Psi_{\overline{K}_{\Psi}}^{u}$ contains an element $P$ such that a neighbor $u_{1}\in P$ of $u$ is not complete to $K_{\Psi}$, in particular $\Psi_{2}^{u}$ is non empty.
-Then $\mathbf{\Psi_{\overline{K}_{\Psi}}^{u}}$-Reduction is valid.
+Given $u\in N(v)$, the neighbors of $u$ in the elements of $\Psi_{I_{1}}^{u}$ are complete to $K_{2}^{u}\cup\{u\}$.
 Proof.
-Let $P_{1}$ be the inclusion-wise maximal subset of $P$ inducing a path ending in $u$ and containing $u_{1}$, let $P^{\prime}$ be its other subset inducing a path ending in $u$.
-Let $P_{2}$ be an element of $\Psi^{u}_{2}$ such that the neighbor $u_{2}$ of $u$ in $P_{2}$ is non adjacent to $u_{1}$.
-Claim 2.2 implies that $P_{1}\cup P_{2}$ induces a path.
-Moreover, $\xi^{P}|_{H}+\xi^{P_{2}}|_{H}=\xi^{P_{1}\cup P_{2}}|_{H}+\xi^{P^{\prime}\cup\{v\}}|_{H}.$ Therefore $\mathbf{\Psi_{\overline{K}_{\Psi}}^{u}}$-Reduction is valid.
+By contradiction, suppose that a neighbor $u_{1}$ of $u$ in some element $P\in\Psi^{u}_{I_{1}}$ is not adjacent to a vertex $u_{2}\in K_{2}^{u}$ contained in some induced path $P_{2}\in\Psi^{u}_{2}$.
+Let $\alpha=\min\Bigl(\Delta_{uv}(\lambda,\Psi),\lambda_{P},\lambda_{P_{2}}\Bigr)$.
+Let us consider the two inclusion wise maximal induced path contained in $P$ ending in $u$, say $_{1}, PP^{\prime}$, and let $P_{1}$ contain $u_{1}$.
+By Claim 2.2, $P_{1}\cup P_{2}$ induces a path of $G$.
+Decreasing $\lambda_{P}$, $\lambda_{P_{2}}$ and increasing $\lambda_{P_{1}\cup P_{2}}$, $\lambda_{P^{\prime}\cup\{v\}}$ yields a nonnegative integer combination $(\lambda^{\prime},\Psi^{\prime})$ of induced paths’ extended incidence vectors.
+Moreover, the following equality holds $\Bigl(\xi^{P}+\xi^{P_{2}}\Bigr)\Bigm\lvert_{G\setminus uv}=\Bigl(\xi^{P_{1}\cup P_{2}}+\lambda_{P^{\prime}\cup\{v\}}\Bigr)\Bigm\lvert_{G\setminus uv}$.
+Therefore, $(\lambda^{\prime},\Psi^{\prime})$ satisfies Conditions (i)-(ii), $\Delta_{e}(\lambda^{\prime},\Psi^{\prime})=\Delta_{e}(\lambda,\Psi)\ \forall e\in\delta(v)\setminus\{uv\}$ and $\Delta_{uv}(\lambda^{\prime},\Psi^{\prime})=\Delta_{uv}(\lambda,\Psi)-\alpha$.
+We finally get $\alpha=0$ and hence $\Delta_{uv}(\lambda,\Psi)=0$.
 ∎
 Lemma 2.8.
-Suppose that $\Psi_{u}^{v}$ contains an element $P$ such that the neighbor $u_{1}$ of $u$, distinct from $v$, is not complete to a vertex $u_{2}$ of $K_{\Psi}$ contained in the path $P_{2}\in\Psi_{2}^{u}$.
-$\mathbf{\Psi_{u}^{v}}$-Reduction is valid and strictly increases $\lambda_{\{u,v\}}$.
-Proof. Let $P_{1}=P\setminus\{v\}$, by Claim 2.2 $P_{1}\cup P_{2}$ induces a path and $\{u,v\}$ induces a path. Moreover, $\xi^{P}|_{H}+\xi^{P_{2}}|_{H}=\xi^{P_{1}}|_{H}+\xi^{\{u,v\}}|_{H}$. ∎
+Given $u\in N(v)$ either $\Delta_{uv}(\lambda,\Psi)=0$ or $\lambda_{u,v}=0$.
+Proof.
+Suppose by contradiction that both $\lambda_{u,v}$ and $\Delta_{uv}(\lambda,\Psi)$ are nonzero.
+By Lemma 2.6 $\Psi_{I_{1}}^{u}$ is nonempty, let $P$ be an element of $\Psi_{I_{1}}^{u}$.
+$P$ contains the two distinct inclusion-wise maximal subsets $P_{1}$ and $P_{2}$ both inducing paths ending in $u$.
+Let $\alpha=\min(\Delta_{uv}(\lambda,\Psi),\lambda_{P},\lambda_{\{u,v\}})$.
+Note that $P_{1}\cup\{v\}$ induces a path.
+Decreasing $\lambda_{u,v}$, $\lambda_{P}$ and increasing $\lambda_{P_{2}}$, $\lambda_{P_{1}\cup\{v\}}$ by $\alpha$ yields a nonnegative integer combination $(\lambda^{\prime},\Psi^{\prime})$.
+As $\xi^{\{u,v\}}+\xi^{P}=\xi^{\{u,v\}\cup P_{1}}+\xi^{P_{2}}$, $(\lambda^{\prime},\Psi^{\prime})$ satisfies Conditions (i)-(ii) and $\Delta_{uv}(\lambda,\Psi)=\Delta_{uv}(\lambda^{\prime},\Psi^{\prime})$.
+However, $(\Psi^{\prime})^{u}$ is nonempty as it contains $P_{2}$, by Lemma 2.5, $\Delta_{uv}(\lambda,\Psi)=0$, a contradiction.
+∎
 Lemma 2.9.
-Suppose that $\Psi_{\overline{K}_{\Psi}}^{u}$ is non empty. $\mathbf{\lambda_{\{u,v\}}}$-Reduction is valid and strictly decreases $\lambda_{\{u,v\}}$.
+Given $u\in N(v)$, either $\Delta_{uv}(\lambda,\Psi)=0$ or $K_{2}^{u}$ is nonempty.
 Proof.
-Given $P\in\Psi_{\overline{K}_{\Psi}}^{u}$, we consider its two inclusion-wise maximal subsets $P_{1}$ and $P_{2}$ inducing a path ending in $u$.
-$P_{1}\cup\{v\}$ induces a path.
-Moreover, $\xi^{P}|_{H}+\xi^{\{u,v\}}|_{H}=\xi^{P_{1}\cup\{v\}}|_{H}+\xi^{P_{2}}|_{H}$.
+Suppose by contradiction that $\Psi_{2}^{u}$ is empty and that $\Delta_{uv}(\lambda,\Psi)$ is strictly positive.
+By Lemma 2.5, Equality (9) holds.
+Inequality (5) associated with $u$ and the empty clique applied to $(x^{*},y^{*})$ yields: $2x_{u}^{*}\geq y^{*}(\delta(u))$.
+By Lemma 2.8, $\lambda_{u,v}=0$.
+Therefore, each $\lambda_{P}$ for $P\in\Psi_{I_{1}}^{u}\cup\Psi_{I_{2}}\cup\Psi^{v}_{u}$ contributes two times in $y^{*}(\delta(u))$, hence the following holds:
+$2x_{u}^{*}\geq 2\Bigl(\lambda(\Psi_{I_{1}}^{u})+\lambda(\Psi_{I_{2}}^{u})\Bigr)+\lambda(\Psi_{u}^{v})+y_{uv}^{*}.$
+Equalities (8)-(9) yields $y_{uv}^{\prime}\geq y_{uv}^{*}$ implying that $\Delta_{uv}(\lambda,\Psi)=0$, a contradiction. ∎
+Lemma 2.10.
+Given $u\in N(v)$, either $\Delta_{uv}(\lambda,\Psi)=0$ or the neighbors of $u$ in the elements of $\Psi_{u}^{v}$ are complete to $K_{2}^{u}\cup\{u\}$.
+Proof.
+Suppose that $\Delta_{uv}(\lambda,\Psi)$ is nonzero.
+By Lemma 2.8, the elements of $\Psi_{u}^{v}$ have at least 3 vertices.
+Let $P=(v,u,u_{1},\dots)$ be an element of $\Psi^{v}_{u}$.
+By construction, $v$ is complete to $K_{2}^{u}\cup\{u\}$.
+Suppose that $u_{1}$ is not complete to $K_{2}^{u}\cup\{u\}$ and, without loss of generality, not adjacent to the vertex $u_{2}\in P_{2}\cap K_{2}^{u}$ for some induced path $P_{2}\in\Psi_{2}^{u}$.
+Let $P_{1}$ be the inclusion-wise maximal induced path included in $P$ ending in $u$ and containing $u_{1}$.
+By Claim 2.2, $P_{1}\cup P_{2}$ induces a path.
+Let $\alpha=\min(\lambda_{P},\lambda_{P_{2}})$, decreasing $\lambda_{P}$, $\lambda_{P_{2}}$ and increasing $\lambda_{P_{1}\cup P_{2}}$, $\lambda_{\{u,v\}}$ by $\alpha$ yields a nonnegative integer combination $(\lambda^{\prime},\Psi^{\prime})$ of induced paths’ extended incidence vectors.
+As $\xi^{P}+\xi^{P_{2}}=\xi^{P_{1}\cup P_{2}}+\xi^{\{u,v\}}$, $(\lambda^{\prime},\Psi^{\prime})$ satisfies Conditions (i)-(iii), and $\Delta_{uv}(\lambda,\Psi)=\Delta_{uv}(\lambda^{\prime},\Psi^{\prime})$.
+Moreover, we have that $\lambda^{\prime}_{\{u,v\}}$ is nonzero, therefore by Lemma 2.8, $\Delta_{uv}(\lambda,\Psi)=0$.
 ∎
-The proof of the following theorem proceeds by induction over the number of vertices and makes use of the previous definitions and operations.
 Theorem 2.11.
-The extended incidence vectors of induced paths of a chordal graph form a Hilbert basis of $\mathcal{C}_{P}(G)$.
+The extended incidence vectors of induced paths of a chordal graph $G$ form a Hilbert basis of $\mathcal{C}_{P}(G)$.
 Proof.
-The statement is trivial for $|V|=1$.
-Induction goes as follows: As $G$ is chordal, it contains a simplicial vertex $v$.
-The graph $G\setminus v$ is chordal and by hypothesis, the induced paths’ extended incidence vectors of $G\setminus v$ form a Hilbert basis of $\mathcal{C}_{P}(G\setminus v)$.
-Let $(x^{*},y^{*})$ belong to $\mathcal{C}_{P}(G)\cap\mathds{Z}^{V\cup E}$, its restriction $(x^{*},y^{*})|_{G\setminus v}$ belongs to $\mathcal{C}_{P}(G\setminus v)\cap\mathds{Z}^{V\cup E}$.
-From now on, $(\Psi,\lambda)$ denotes an arbitrary family of induced paths of $G$ with integer coefficients satisfying $(x’,y’)|_{G\setminus v}=(x^{*},y^{*})|_{G\setminus v}$, where (x,y)(x’,y’) is the point induced by $(\Psi,\lambda)$.
-All operations defined above transform $(\Psi,\lambda)$ into another such family and preserve: integrality of $\lambda$, the fact that the elements of $\Psi$ induce paths of $G$ and $(x’,y’)|_{H}=(x^{*},y^{*})|_{H}$.
-Note that the latter requirement implies that the operations not only preserve the value of $\lambda(\Psi)$ over elements in $G\setminus v$ but also on edges in $\delta(v)\setminus\{uv\}$.
-The goal of the upcoming sequence of reductions is to achieve $(x^{\prime},y^{\prime})=(x^{*},y^{*})$. For this we consider the slackness;
-$\Delta:=\sum_{uv\in\delta(v)}\Delta(u,v)\text{ where }\Delta(u,v)=y_{uv}^{*}-y_{uv}^{\prime}.$
-The proof proceeds by fixing a neighbor $u$ of $v$ and modifying $(\Psi,\lambda)$ until the equality $\Delta(u,v)=0$ holds.
-Since reductions for a fixed $u$ do not affect $\Psi_{w}^{v}$ for $w\neq u$, and that elements of $\Psi^{u}_{w}$ are the only ones contributing to the value of $y_{vw}^{\prime}$, the procedure can be applied independently for each neighbor of $v$.
-An induced path $P$ is removed from $\Psi$ if $\lambda_{P}$ is set to 0.
-Lemmas 2.32.5 give conditions for some valid reductions strictly decreasing $\Delta(u,v)$ to be applied for respectively $\alpha=\min(\Delta(u,v),\lambda_{P})$, $\alpha=\min(\Delta(u,v),\lambda_{P})$, and $\alpha=\min(\lambda_{P},\lambda_{P_{2}},\Delta(u,v))$.
-Since the corresponding reductions strictly decrease either $\Delta(u,v)$, at some point, none of Lemma 2.32.6 apply for the fixed neighbor $u$.
-This implies that, either $\alpha=0$ and $\Delta(u,v)=0$ which ends the proof or the following holds:
-  • $\Psi^{u}$, $\Psi^{u}_{I_{1}}$ both are empty
-  • the neighbors of $u$ in the elements of $\Psi^{u}_{I_{0}}$ are complete to $K_{\Psi}$.
-By construction, the only elements of $\Psi$ which contribute to the value of $y_{uv}^{\prime}$ belong to $\Psi_{u}^{v}$, therefore the following holds :
-$y_{uv}^{\prime}=\lambda(\Psi_{u}^{v}).$ (8)
-As $\Psi^{u}$ is empty, the following holds:
-$x^{*}_{u}=x_{u}^{\prime}=\lambda(\Psi_{2}^{u})+\lambda(\Psi_{\overline{K}_{\Psi}}^{u})+\lambda(\Psi_{u}^{v})\\$ (9)
-Suppose that $\Psi^{u}_{I_{0}}$ is empty.
-As $(x^{*},y^{*})$ belongs to $\mathcal{C}_{P}(G)$, by inequality (5) associated with $u$ and $K=N[v]\setminus\{u\}$ we get: $x_{u}^{*}\geq y^{*}(\delta(u,K))$ and the right hand side is larger than or equal to $\lambda(\Psi_{2}^{u})+y_{uv}^{*}$.
-Using equalities (8) and (9) we get : $y_{uv}^{\prime}\geq y_{uv}^{*}$ and hence $\Delta(u,v)=0$.
-Assume henceforth that $\Psi_{\overline{K}_{\Psi}}^{u}$ is non-empty.
-Suppose $\Psi_{2}^{u}$ to be empty, Inequality (5) associated with $u$ and $K=\emptyset$ applied to $(x^{*},y^{*})$ yields: $2x_{u}\geq y^{*}(\delta(u))$.
-Note that the right hand side is larger than or equal to $2\lambda(\Psi_{I}^{u})+\lambda(\Psi_{u}^{v})+y_{uv}^{*}$ as each $\lambda_{P}$ for $P\in\Psi_{I}^{u}\cup\Psi^{v}_{u}$ contributes two times in $y^{*}(\delta(u))$ and $\lambda(\Psi^{v}_{u})=y_{uv}^{*}$.
-Equalities (9) and (8) yield $y_{uv}^{\prime}\geq y_{uv}^{*}$ and $\Delta(u,v)=0$.
-We may now assume that $\Psi_{2}^{u}$ is nonempty.
-Therefore, $K$ is non-empty and the neighbors of $u$ in the elements of $\Psi_{\overline{K}_{\Psi}}^{u}$ are complete to $K_{\Psi}$.
-Suppose that the conditions of Lemma 2.6 hold, then applying $\Psi_{u}^{v}$-Reduction with $\alpha=\min(\lambda_{P},\lambda_{P_{2}})$ yields an element of $\Psi_{K_{\Psi}}^{u}$, therefore, the conditions of Lemma 2.4 hold and $\Psi_{K_{\Psi}}^{u}$-Reduction applies.
-This sequence of operations strictly decreases $\Delta(u,v)$.
-Suppose that the conditions of Lemma 2.7 hold, applying $\lambda_{u}^{v}$-Reduction with $\alpha=\min(\lambda_{P},\lambda_{u,v}$) yields an element of $\Psi_{u}^{v}$, and one of $\Psi^{u}$.
-Therefore, the conditions of Lemma 2.3 hold, and $\Psi^{u}$-Reduction applies.
-This sequence of operations also decreases $\Delta(u,v)$.
-As $\Psi^{u}$-Reduction, $\Psi_{K_{\Psi}}^{u}$-Reduction, $\Psi_{\overline{K}_{\Psi}}^{u}$-Reduction, and the two latter sequences of reductions strictly decrease $\Delta(u,v)$, we may assume that none of Lemmas 2.32.7 apply or that $\Delta(u,v)=0$.
-Therefore, the neighbors of $u$ in the elements of $\Psi^{v}_{u}$ are complete to $K_{\Psi}$ and $\lambda_{\{u,v\}}=0$.
-As $(x^{*},y^{*})$ is a point of $\mathcal{C}_{P}(G)$, it satisfies the constraint (5) associated with $K_{\Psi}$ and $u$ that is (where $C(S)$ denotes the set of vertices complete to $S$):
-$2x_{u}^{\prime}=2x_{u}^{*}\geq 2y^{*}(\delta(u,K_{\Psi}))+y^{*}(\delta(u,C(K_{\Psi}\cup\{u\})))$ (10)
-By construction, each element $P$ of $\Psi_{u}^{v}$ contains both $v$ and another neighbor of $u$ and both are complete to $K_{\Psi}$.
-This implies that $\lambda_{P}$ contributes two times in $y^{*}[\delta(u,C(K_{\Psi}\cup\{u\}))]$.
-By construction, $y^{*}(\delta(u,C(K_{\Psi}\cup\{u\})))=y^{*}_{uv}+y^{*}[\delta(u,C(K_{\Psi}\cup\{u\})\setminus\{uv\})]=y^{*}_{uv}+\lambda(\Psi_{u}^{v})+2\lambda(\Psi_{\overline{K}_{\Psi}}^{u})$.
-Using equalities (8) and (9) we get:
-$0\geq y^{*}_{uv}-\lambda(\Psi_{u}^{v})$
-which finally shows that $y_{uv}^{\prime}\geq y_{uv}^{*}$.
-Repeat this procedure until $y_{uv}^{*}=y_{uv}^{\prime}\ \forall uv\in\delta(v)$.
-Setting $\lambda_{\{v\}}=(x^{*}_{v}-y^{*}(\delta(v)))$ for the trivial induced path $\{v\}$ yields the desired integer combination of $(x^{*},y^{*})$ using induced paths’ extended incidence vectors.
-As induced trees of $C_{4}$ coincide with its induced path, Theorem 2.8 is not a characterization of the graph for which the induced path extended incidence vectors have IDP.
-2.2 Extended formulation for the induced path polytope
+Similarly to the proof of Theorem 1.1, the first step is to show that $\Delta_{e}(\lambda,\Psi)$ is equal to 0 for every $e\in\delta(v)$ for $(\lambda,\Psi)$ satisfying Conditions (i)-(iii).
+This is done through the use of an inequality (5) associated with $u$.
+Therefore, we need to exhibit the right clique of $N(u)$.
+We iteratively construct this clique that we denote by $K_{\Psi}$, and note $\Psi_{K_{\Psi}}$ the subset of $\Psi^{u}_{2}\cup\Psi_{I_{2}}^{u}$ whose elements contain a vertex of $K_{\Psi}$.
+Note that $\Psi_{K_{\Psi}}$ is supposed to be changing along to $K_{\Psi}$.
+Start with $K_{\Psi}=K_{2}^{u}$.
+If there exists an element $P$ of $\Psi_{I_{2}}^{u}\setminus\Psi_{K_{\Psi}}$ such that the vertex $w\in\Bigl(P\cap N(u)\Bigr)\setminus N(v)$ is not complete to $K_{\Psi}$, add $P$ to $\Psi_{K_{\Psi}}$ and add the other neighbor of $u$ in $P$ to $K_{\Psi}$.
+This procedure is applied until the neighbors of $u$ in the elements of $\Psi_{I_{2}}^{u}\setminus\Psi_{K_{\Psi}}$ are complete to $K_{\Psi}\cup\{u\}$.
+By Lemmas 2.7-2.10, the neighbors of $u$ in the elements of $\Psi_{I_{1}}^{u}\cup\Psi_{u}^{v}$ are complete to $K_{2}^{u}\cup\{u\}$.
+Moreover, by the previous construction, the elements of $\Psi^{u}_{I_{2}}\setminus\Psi_{K_{\Psi}}$ satisfy this property for the whole clique $K_{\Psi}\cup\{u\}$.
+It remains to show that the neighbors of $u$ in the elements of $\Psi_{I_{1}}^{u}\cup\Psi_{u}^{v}$ are complete to $K_{\Psi}\setminus K_{2}^{u}$.
+We do so by reorganizing the elements of $\Psi_{K_{\Psi}}$.
+Each vertex of $K_{\Psi}\setminus K_{2}^{u}$ is contained in an element of $\Psi_{I_{2}}^{u}\cap\Psi_{K_{\Psi}}$.
+Let $P$ be an element of $\Psi_{I_{1}}^{u}\cup\Psi^{v}_{u}$ and let $P^{\prime}$ be an element of $\Psi_{K_{\Psi}}\setminus\Psi_{2}^{u}$.
+As $P^{\prime}$ has been added to $\Psi_{K_{\Psi}}$ by the previous construction, there exists a sequence of induced path, say $(P_{1},\dots,P_{\ell})$ such that $P_{1}=P^{\prime}$, $P_{\ell}\in\Psi_{2}^{u}$ and all the other elements belong to $\Psi_{I_{2}}^{u}$, such that the vertex $u_{i}\in(N(u)\cap P_{i})\setminus K_{\Psi}$ is nonadjacent to the vertex $v_{i+1}\in K_{\Psi}\cap P_{i+1}\ \forall i\in\{1,\dots,\ell-1\}$.
+For each $P_{i},\ i\in\{1,\dots,\ell-1\}$, we consider the two maximal induced path $U_{i},V_{i}$ included in $P_{i}$ and ending in $u$, with $U_{i}$ containing $u_{i}$ and $V_{i}$ containing $v_{i}$, we set $V_{\ell}=P_{\ell}$.
+Claim 2.2 implies that $U_{i}\cup V_{i+1}\ \forall i\in\{1,\dots,\ell-1\}$ induces a path, see Figure 4 for an illustration where the dashed line represents a nonedge.
+Therefore, for $\alpha=min\{\lambda_{P_{i}}:i\in\{1,\dots,\ell\}\}$ decreasing $\lambda_{P_{i}}\ \forall i\in\{1,\dots,\ell\}$, increasing $\lambda_{V_{i+1}\cup U_{i}}\forall i\in\{1,\dots,\ell\-1\}$ and $\lambda_{V_{1}},$ by $\alpha$ yields a nonnegative combination of induced paths’ extended incidence vectors $(\lambda^{\prime},\Psi^{\prime})$.
+Moreover $\sum_{i=1}^{\ell}\xi^{P_{i}}=\xi^{V_{1}}+\sum_{i=1}^{\ell-1}\xi^{U_{i}\cup V_{i+1}}$, implying that, $(\lambda^{\prime},\Psi^{\prime})$ satisfies Conditions (i)-(iii) and $(\Psi^{\prime})^{u}_{2}$ contains $V_{1}$.
+By Lemmas 2.7 and 2.10, $P$ contains two vertices adjacent to the only vertex in $V_{1}\cap K_{\Psi}$.
+Note that this observation can be made for each element in $\Psi_{K_{\Psi}}\setminus\Psi_{2}^{u}$ and each element in $\Psi^{u}_{I_{1}}\cup\Psi^{v}_{u}$.
+This implies that the neighbors of $u$ in the elements of $\Psi^{u}_{I_{1}}\cup\Psi^{v}_{u}$ have two vertices complete to $K_{\Psi}\cup\{u\}$.
+As $(x^{*},y^{*})$ is a point of $\mathcal{C}_{P}(G)$, it satisfies the constraint (5) associated with $K_{\Psi}$ and $u$ that is:
+$2x_{u}^{\prime}=2x_{u}^{*}\geq y^{*}\Biggl(\delta\Bigl(u,C\left(K_{\Psi}\cup\{u\}\right)\Bigl)\Biggr)+2y^{*}\Biggl(\delta\Bigl(u,K_{\Psi} \Bigr)\Biggr)$ (10)
+By the previous paragraph, each element $P$ of $\Psi_{u}^{v}\cup\Psi_{I_{1}}^{u}\cup\Psi_{I_{2}}^{u}\setminus\Psi_{K_{\Psi}}$ contains $u$ and its neighbors in $P$ are complete to $K_{\Psi}$.
+This implies that $\lambda_{P}$ contributes two times in
+$y^{*}\Biggl(\delta\Bigl(u,C(K_{\Psi}\cup\{u\})\Bigr)\Biggr).$
+Therefore, the following holds:
+$\begin{array}[]{rl}y^{*}\Biggl(\delta\Bigl(u,C(K_{\Psi}\cup\{u\})\Bigr)\Biggr)&=y^{*}_{uv}+y^{*}\Biggl(\delta\Bigl(u,C(K_{\Psi}\cup\{u\})\Bigr)\setminus\{uv\}\Biggr)\\ &=y^{*}_{uv}+\lambda(\Psi_{u}^{v})+2\lambda(\Psi_{I_{1}}^{u})+2\lambda(\Psi_{I_{2}}^{u}\setminus\Psi_{K_{\Psi}}).\end{array}$
+Moreover, each element of $\Psi_{K_{\Psi}}$ intersects $K_{\Psi}$ and contains $u$, implying that it contributes one time in $y^{*}\Biggl(\delta\Bigl(u,K_{\Psi} \Bigr)\Biggr)$.
+Using equality (9) and inequality (10) we get:
+$2\Bigl(\lambda(\Psi_{2}^{u})+\lambda(\Psi_{I_{2}}^{u})+\lambda(\Psi_{I_{1}}^{u})+\lambda(\Psi_{u}^{v})\Bigr)\geq y^{*}_{uv}+\lambda(\Psi_{u}^{v})+2\lambda(\Psi_{I_{1}}^{u})+2\lambda(\Psi_{I_{2}}^{u}\setminus\Psi_{K_{\Psi}})+2\Bigl(\lambda(\Psi_{K_{\Psi}})\Bigr)$
+Moreover, $\Psi_{I_{2}}^{u}=(\Psi_{I_{2}}^{u}\cap\Psi_{K_{\Psi}})\cup(\Psi_{I_{2}}^{u}\setminus\Psi_{K_{\Psi}})$ and $\Psi_{I_{2}}^{u}\cap\Psi_{K_{\Psi}}=\Psi_{K_{\Psi}}\setminus\Psi_{2}^{u}$.
+Therefore, $y_{uv}^{\prime}\geq y_{uv}^{*}$ by equality (8).
+This reasoning can be made for each $uv\in\Delta(v)$.
+Therefore, $\Delta(\lambda,\Psi)=0$.
+Setting $\lambda_{\{v\}}=(x^{*}_{v}-y^{*}(\delta(v)))$ for the trivial induced path $\{v\}$ yields the desired nonnegative integer combination of $(x^{*},y^{*})$ with induced paths’ extended incidence vectors.
+∎
+[…]
+Figure 4: Structure of the paths exhibited in Theorem 2.11
+Remember that $C_{4}$ is a non-chordal graph for which the extended incidence vectors of induced trees form a Hilbert basis.
+Therefore, as induced trees of $C_{4}$ coincide with its induced path, Theorem 2.11 is not a characterization of the graph for which the induced path extended incidence vectors form a Hilbert basis.
+This result also raises the question of a characterization of the graphs for which induced paths’ extended incidence vectors form Hilbert bases.
+2.2 Induced paths of chordal graphs and integral systems
 In this section, we discuss the consequences of Theorem 2.11 and use it to give a compact description of the induced path polytope of chordal graphs.
 Corollary 2.12.
-The induced path cone and polytope of chordal graphs both have IDP.
+The induced path cone and polytope of chordal graphs have IDP.
 Similarly to what we do for the induced tree polytope, we exhibit a description of the induced path polytope of chordal graphs.
@@
 $(\Rightarrow)$ The extreme points of $\mathcal{P}(G)$ are scalings of $\mathcal{C}_{P}(G)$’s generators which belong to (11).
-Theorem 2.8 tells that the set of induced paths’ extended incidence vectors generate $\mathcal{C}_{P}(G)$ moreover, they belong to (11), which concludes the proof.
+Theorem 2.11 tells that the set of induced paths’ extended incidence vectors generate $\mathcal{C}_{P}(G)$ moreover, they belong to (11).
+Finally, $\mathcal{P}(G)$ is included in $\mathcal{T}(G)$ which is bounded by Theorem 1.3.
+Therefore, $\mathcal{P}(G)$ is also bounded.
 $(\Leftarrow)$ The proof is by contradiction. The same non binary point exhibited in the proof of Theorem 1.3 also belongs to $\mathcal{P}(G)$. ∎
@@
 Proof.
-$(\Rightarrow)$ Let us consider the connected components $\overline{G}_{1},\dots,\overline{G}_{k}$ of $\overline{G}[C_{G}(K\cup\{w\})]$ and suppose by contradiction that for $i\leq\ell$, $\overline{G}_{i}$ is bipartite composed of the stable sets induced by $S_{i}^{1},S_{i}^{2}$.
+$(\Rightarrow)$ Let us consider the connected components $\overline{G}_{1},\dots,\overline{G}_{k}$ of $\overline{G}[C_{G}(K\cup\{w\})]$ and suppose by contradiction that for $i\leq\ell$, $\overline{G}_{i}$ is bipartite composed of the stable sets induced by $S_{i}^{1},S_{i}^{2}$ for some $l\in\{1,\dots k\}$.
 By construction, $K_{j}=\bigcup_{i=1}^{\ell}S_{i}^{j}\cup K\cup\{w\},\ j=1,2$ is a clique of $G$ such that $C_{G}(K_{j})=\bigcup_{i=\ell+1}^{k}V(\overline{G}_{i})$.
-By summing $2x_{w}\geq 2y(\delta(w,K_{j}))+y(\delta(w,C(K_{j})))$ for $j=1,2$ we obtain the inequality associated with $K$ and $w$ showing that it is dominated and hence not facet-defining.
-$(\Leftarrow)$ We exhibit |V|+|E|1|V|+|E|-1 linearly independent points satisfying (5) with equality, the fact that the equality constraints submatrix’s rank is at least 1 because of equality 11 concludes.
+By summing the valid inequalities $2x_{w}\geq 2y(\delta(w,K_{j}))+y(\delta(w,C(K_{j})))$ for $j=1,2$ we obtain the inequality associated with $K$ and $w$.
+This shows that it is dominated and hence not facet-defining.
+$(\Leftarrow)$ We exhibit |V|+|E|1|V|+|E|-1 linearly independent points satisfying (5) with equality for a given $w$ and $K$.
+This is sufficient as by equality (11) the dimension of $\mathcal{P}(G)$ is at most |V|+|E|1|V|+|E|-1.
 We consider $K_{w}=K\cup\{w\}$.
@@
   4. 4.
-    $\xi^{\{u,w,v\}}\ \forall uw\in\delta(w)\setminus(\delta(K)\cup\delta(C(K_{w})))$ for some vertex $v\in K$ non adjacent to $u$
-The remaining $|\delta(w,C(K_{w}))|$ points are built the following way: for each connected components $\overline{G}_{c}$ of $\overline{G}[C(K_{w})]$, let us consider $T_{c}\subseteq E(\overline{G}_{c})$ inducing a subgraph containing an odd cycle and hitting every vertex of $G_{c}$, $T_{c}$ exists by hypothesis and may choosen such that $|T_{c}|=|V(\overline{G}_{c})|$.
-This can be done by starting from an odd cycle and then adding the missing edges for this set to span the vertices.
+    $\xi^{\{u,w,v\}}\ \forall uw\in\delta(w)\setminus\Biggl(\delta\Bigl(K\Bigr)\cup\delta\Bigl(C(K_{w})\Bigr)\Biggr)$ for some vertex $v\in K$ non adjacent to $u$
+Points 1 to 4 satisfy inequality (5) for $w$ and $K$ with equality.
+Moreover, they can be ordered so that the corresponding matrix is triangular.
+For the remaining $\left|\delta\Bigl(w,C(K_{w})\Bigr)\right|$ elements, two edges in $\delta\Bigl(C(K_{w})\Bigr)$ are needed to achieve equality; therefore, it is not possible to keep the triangular structure.
+The remaining $\left|\delta\Bigl(w,C(K_{w})\Bigr)\right|$ points are built the following way: for each connected components $\overline{G}_{c}$ of $\overline{G}[C(K_{w})]$, let us consider $T_{c}\subseteq E(\overline{G}_{c})$ inducing a subgraph containing an odd cycle and hitting every vertex of $G_{c}$.
+$T_{c}$ exists and may be chosen such that $|T_{c}|=|V(\overline{G}_{c})|$: start with an odd cycle and then add the missing edges for this set to span the vertices.
 Now, match each edge of $\delta(w,C(K_{w}))$ to a distinct edge of $T=\bigcup T_{c}$ sharing an extremity, note that $|T|=|C(K_{w})|=|\delta(w,C(K_{w}))|$.
@@
 In $G$, $u$ and $v$ are non-adjacent, hence $\xi^{u,v,w}$ is the incidence vector of an induced path for which inequality (5) for $w$ and $K$ is tight.
-This construction is necessary as such a path will contain two edges of $\delta(w,C(K_{w}))$, making it impossible to continue a diagonal equal to 1.
-We refer to Figure 4 for an illustration of the matrix whose rows are the points previously exhibited, and "?" stands for the submatrices associated with incidence vectors of 3-sized paths and whose value does not need to be explicit for the validity of the proof.
-$H$ is similar to the incidence matrix of the non bipartite graph induced by $T$, by Theorem 2.1 of [GROSSMAN1995213] it has full row rank $|C(K_{w})|$.
-Full row rank is obtained by the fact that the matrix is block triangular with a non-zero diagonal over the points (1)–(4), and the fact that the last set of points forms a submatrix of rank $|C(K_{w})|$ which are the only ones involving elements of $\delta(w,C(K_{w}))$.
-This implies that the matrix has rank $|V|+|E|$, therefore the inequality is facet defining.
+We refer to Figure 5 for an illustration of the matrix whose rows are the points previously exhibited and in the same order, therefore, the lines containing the submatrix $H$ correspond to the last set of $\left|\delta\Bigl(w,C(K_{w})\Bigr)\right|$ points.
+The symbol "?" stands for the submatrices associated with incidence vectors of 3-vertex paths and whose value does not need to be explicit for the validity of the proof.
+$H$ is identical to the incidence matrix of the non-bipartite graph induced by $T$, by Theorem 2.1 of [GROSSMAN1995213] it has full row rank $|C(K_{w})|$.
+Full row rank of the whole matrix is obtained by the facts that the matrix is block triangular with a non-zero diagonal over the points (1)–(4), and that $H$ is the maximal submatrix with no 0 row having elements of $\delta\Bigl(w,C(K_{w})\Bigr)$ as column set.
+This implies that the whole matrix has rank $|V|+|E|$, therefore the inequality is facet defining.
 ∎
@@
 We show that the non-dominated inequalities (5) come in a polynomial number and can be enumerated efficiently.
+We introduce a necessary condition on the clique $K$ for inequality 5 to be facet-defining.
+After that, we show that the set of cliques verifying such a property come in a polynomial number in chordal graphs.
 Given a fixed vertex $w$, for any two cliques $K\subseteq N(w)$ and $K^{\prime}\subseteq N(w)$, let $K_{w}=K\cup\{w\}$, $K_{w}^{\prime}=K^{\prime}\cup\{w\}$.
-If $K\subseteq K^{\prime}$, then we have $C(K_{w}^{\prime})\subseteq C(K_{w})$, for this reason, we say that a clique defines an orbit of $w$ if there exists no $K^{\prime}$ strictly containing $K$ such that $K_{w}\cup C(K_{w})\subseteq K_{w}^{\prime}\cup C(K_{w}^{\prime})$.
-Suppose that a clique $K$ is not orbit-defining, let $K^{\prime}$ be the orbit-defining clique containing $K$, inequality (5) associated with $(w,K)$ is dominated with respect to inequality (5) associated with $(w,K^{\prime})$ and non-negativity constraints.
+If $K\subseteq K^{\prime}$, then we have $C(K_{w}^{\prime})\subseteq C(K_{w})$.
+We say that a clique defines an orbit of $w$ if there exists no $K^{\prime}$ strictly containing $K$ such that $K_{w}\cup C(K_{w})\subseteq K_{w}^{\prime}\cup C(K_{w}^{\prime})$.
+Suppose that a clique $K$ is not orbit-defining, let $K^{\prime}$ be the orbit-defining clique containing $K$.
+Inequality (5) associated with $(w,K)$ is dominated with respect to inequality (5) associated with $(w,K^{\prime})$ and non-negativity constraints.
 Theorem 2.16.
-Given a chordal graph $G$, the set of non-dominated inequalities (5) has polynomial size.
+Given a chordal graph $G$, the set of non-dominated inequalities (5) has size $O(|V|^{2})$.
 Proof.
 We build a chordal auxiliary graph $H$ to $G[N(w)]$, whose maximal cliques are in bijection with orbit-defining cliques of $G[N(w)]$.
-Polynomial size of $H$ and linear number of maximal cliques of chordal graphs will conclude.
+Polynomial size of $H$ and linear number of maximal cliques of chordal graphs will concludes.
 The auxiliary graph $H$ is built from $G[N(w)]$ by first adding a true twin to each vertex of $N(w)$, let $W$ be the set of added twins.
 Secondly, add all the edges between elements of $W$.
-A perfect elimination ordering of $H$ is given by any perfect elimination of $G[N(w)]$ and followed by any permutation of the vertices of $W$.
-This proves that $H$ is chordal.
+A perfect elimination ordering of $H$ is given by any perfect elimination of $G[N(w)]$ and followed by any permutation of the vertices of $W$. This proves that $H$ is chordal.
 Let $K_{H}$ be a maximal clique of $H$.
@@
 By construction, $K^{\prime}\cup K_{W^{\prime}}$ is a clique of $H$ strictly containing $K_{H}$, a contradiction.
-Now, let $K$ define an orbit of $G[N(w)]$, we associate to $K$ a unique maximal clique of $H$ as follows.
+Conversely, let $K$ define an orbit of $G[N(w)]$, we associate to $K$ a unique maximal clique of $H$ as follows.
 Let $W_{K}\subseteq W$ be the set of twins of $C(K)\cup K$, $K_{H}=W_{K}\cup K$ is a clique of $H$ by construction.
@@
 By construction we have $K\subsetneq K\cup\{v\}$ and $K\cup C(K)\subseteq K\cup\{v\}\cup C(K\cup\{v\})$, a contradiction to $K$ defining an orbit.
+Since for each vertex $w$, orbit-defining cliques of $G[N(w)]$ are in one-to-one correspondence with the maximal cliques of a chordal graph with at most $2|V|$ vertices, their number is $O(V)$ for fixed $w$.
+Summing over all $w\in V$, the total number of non-dominated inequalities (5) is at most $O(|V|^{2})$.
 ∎
-Since for each vertex $w$, orbit-defining cliques of $G[N(w)]$ are in one-to-one correspondence with the maximal cliques of a chordal graph with at most $2|V|$ vertices, their number is $O(V)$ for fixed $w$.
-Summing over all $w\in V$, the total number of non-dominated inequalities (5) is $O(|V|^{2})$.
 Corollary 2.17.
 There exists an extended formulation for the induced path polytope of chordal graphs with size $O(|V|^{2})$.
+Corollary 2.18.
 The maximum vertex and weighted induced path problem is solvable in polynomial time in chordal graphs.
-It turns out that the target of Theorems 2.12 and 2.13 differ, as there exist orbit-defining cliques in chordal graphs whose associated inequalities are not facet-defining as given by Figure 5.
-Vertices 1, 2, and 3 form an orbit defining clique such that $C(\{1,2,3\})=\{4,5,6,7\}$ however, the complement of the subgraph induced by $\{4,5,6,7\}$ is a hole of size 4 which is bipartite therefore, Theorem 2.12 implies that its associated inequalities are not facet defining.
+The targets of Theorems 2.15 and 2.16 differ: there exist orbit-defining cliques in chordal graphs whose associated inequalities are not facet-defining as given by Figure 6.
+Vertices 1, 2, and 3 form an orbit defining clique such that $C(\{1,2,3\})=\{4,5,6,7\}$ however, the complement of the subgraph induced by $\{4,5,6,7\}$ is a hole of length 4 which is bipartite therefore, Theorem 2.15 implies that besides defining an orbit, its associated inequalities are not facet defining.
 11223344556677
@@
 Conclusion
-In this paper, we show that the induced tree and path extended incidence vectors of chordal graphs form a Hilbert basis.
-We use this fact to give a compact linear system of inequalities whose binaricity characterizes chordality.
+In this article, we show that the induced tree and path extended incidence vectors of chordal graphs respectively, form two Hilbert bases.
+We use this fact to give two compact linear systems of inequalities whose binaricity characterizes chordality.
 It turns out that there exist non-chordal graphs for which induced tree and path extended incidence vectors form a Hilbert basis, which raises the question of characterizing such graphs.
-In general, showing that a set of vectors $\mathcal{V}$ forms a Hilbert basis is used to show that a linear system whose rows are elements of $\mathcal{V}$ is TDI.
-In this paper, we are not considering the programs whose rows are extended induced path or induced tree of chordal graphs.
-Finding some problems that are naturally formulated as integer linear programs whose linear constraint set has such a structure would lead to TDI-ness results.
+These results imply that both corresponding linear optimization problems are polynomially solvable.
+In general, showing that some subsets of vectors of a ground set $\mathcal{V}$ forms Hilbert bases is used to show that the linear system whose rows are elements of $\mathcal{V}$ is TDI.
+In this work, we do not consider such systems and observe that extended incidence vectors of induced trees and path form Hilbert bases in chordal graphs.
+It is still a question whether these new Hilbert basis can be of any use to find new TDI systems.
 Statements and Declarations